### v1

# ANID: How Far Are We? Evaluating the Discrepancies Between AI-synthesized Images and Natural Images through Multimodal Guidance

Abstract
In the rapidly evolving field of Artificial Intelligence Generated Content (AIGC), one of the key challenges is distinguishing AI-synthesized images from natural images. Despite the remarkable capabilities of advanced AI generative models in producing visually compelling images, significant discrepancies remain when these images are compared to natural ones. To systematically investigate and quantify these discrepancies, we introduce an AI-Natural Image Discrepancy Evaluation benchmark aimed at addressing the critical question: how far are AI-generated images (AIGIs) from truly realistic images? We have constructed a large-scale multimodal dataset, the Distinguishing Natural and AI-generated Images (DNAI) dataset, which includes over 440,000 AIGI samples generated by 8 representative models using both unimodal and multimodal prompts, such as Text-to-Image (T2I), Image-to-Image (I2I), and Text vs. Image-to-Image (TI2I). Our fine-grained assessment framework provides a comprehensive evaluation of the DNAI dataset across five key dimensions: naive visual feature quality, semantic alignment in multimodal generation, aesthetic appeal, downstream task applicability, and coordinated human validation. Extensive evaluation results highlight significant discrepancies across these dimensions, underscoring the necessity of aligning quantitative metrics with human judgment to achieve a holistic understanding of AI-generated image quality. Code is available at https://github.com/ryliu68/ANID.
Introduction
With the rapid advancement of deep learning techniques, the proliferation of Artificial Intelligence Generated Content (AIGC) has garnered significant attention across various domains, such as e-commerce, gaming, medicine, animation, and autonomous driving (Li et al. 2024; Qian et al. 2024). AI-generated Images (AIGI) have been one of the mainstream forms of AIGC, and a variety of AI image generative models have been proposed to make the generated synthetic images as realistic as natural images, ranging from the earlier generative adversarial networks (GANs) (Tao et al. 2023, 2022), advanced diffusion models (DMs) (Xu et al. 2023b; Wei et al. 2023) to the large multimodal generative model like DALL·E (Ramesh et al. 2022).
Despite the proliferation of AI image generative models, AI-generated images are not qualified for real-world applications, and the discrepancies between generated synthetic images and realistic natural images still exist (Li et al. 2023a). Therefore, many AI-generated image quality assessment and evaluation methods have been proposed (Wang et al. 2023; Xu et al. 2023a; Hu et al. 2023). For example, Wang et al. (Wang et al. 2023) established an AIGC Image Quality Assessment Database called AIGCIQA2023 from 6 text-to-image generative models, and devised a subjective evaluation framework to assess human visual preferences for each AI-generated image from three aspects, e.g., quality, authenticity, and text-to-image correspondence. PKU-I2IQA (Yuan et al. 2023) constructed a human perception-based image-to-image database named PKU-I2IQA and conducted a subject analysis based on both no-reference and full-reference methods. Different from prior studies focusing on establishing the AIGI dataset and conducting a subjective evaluation with human feedback, another group of studies (e.g. Pick-a-Pic (Kirstain et al. 2023), HPS v2 (Wu et al. 2023), TIFA (Hu et al. 2023), and ImageReward (Xu et al. 2023a)) focused on training a unified score model as the automatic evaluation metrics for AIGC image quality assessment (AIGIQA) measures, in contrast with traditional image quality assessment for natural images measures like Inception Score (Salimans et al. 2016) and FID (Heusel et al. 2017).
However, some issues and problems have not yet been fully answered and solved. First, prior studies only considered unimodal promopted generation, e.g. Text-to-Image (T2I) or Image-to-Image (I2I), ignoring multimodal prompted content like Text vs.Image-to-Image (TI2I). The data size of the established T2I or I2I dataset is relatively small (around thousands of images), and might be insufficient for comprehensive analysis and evaluation. Second, prior studies typically focused on the perceptual quality of AIGC images with traditional image quality assessment metrics, the text-to-image correspondence, and subjective human preference. No studies have fully answered the important questions about what discrepancies exist between generated synthetic images and realistic natural images, and they lacked a systematic and compressive assessment and evaluation to investigate and interpret the discrepancies between AI-generated synthetic images and realistic natural images. The investigation and understanding of the discrepancies are key ways to realize the practical application of AIGC images in real-world scenarios and make breakthroughs in AIGC.
To address the above challenges, we proposed an AI-Natural Image Discrepancy Evaluation Benchmark to investigate and interpret what discrepancies still remained between AI-generated images and natural images, finally answering the important question “How far are AI generative models with respect to the visual forms of images?”. Especially, we have two important contributions with reference to the key problems. Large Multimodal Evaluation Dataset construction: we have curated a large multimodal evaluation dataset called DNAI (Distinguishing Natural and AI-generated) dataset, with about 440,000 AIGI generated from 8 representative generative models based on both unimodal and mutimodal guidance including Text-to-Image (T2I), Image-to-Image (I2I), and Text vs. Image-to-Image (TI2I). The data size of our DNAI dataset scales to 100x that of prior datasets. Fine-grained Evaluation Framework, we propose a fine-grained evaluation framework to conduct the systematic and comprehensive assessment and evaluation of DNAI, covering 5 diverse and important aspects, including naive visual feature quality, semantic alignment among multimodal generation, aesthetic appeal, downstream applicability, and the coordinated human validation.
Through our DNAI dataset and fine-grained evaluation framework, we conduct extensive benchmark analysis and evaluation and conclude key insights to answer the discrepancy questions:
  • Significant discrepancies in key Areas: AI-generated images exhibit substantial discrepancies from natural images in terms of quantitative measures from all aspects, with about 10% to 30%.
  • Multimodal Alignment: Different prompted generation might have different semantic alignment scores. Generated images prompted with texts including both T2I and TI2I demonstrated remarkable semantic alignments.
  • Downstream Task Applicability: Significant differences exist in the usability of AI-generated images vs. natural ones in downstream tasks, highlighting the need for further improvements of generative models to ensure practical applicability in real-world scenarios.
  • Human Evaluation vs. Quantitative Metrics: Human evaluation results reveal larger discrepancies compared with quantitative metrics. It validates the necessity of incorporating human evaluation for coordination.
Related Work
Natural Image Evaluation:
Over the past decades, numerous image quality assessment methods have been developed to evaluate traditional natural images (N. et al. 2015). These methods have focused on various visual features and properties, such as perceptual appearance (Zhang et al. 2018), naturalness (Ma et al. 2018), and aesthetics (Esfandarani and Milanfar 2018). For instance, BRISQUE evaluates natural image quality by analyzing spatial domain features, while PIQE assesses perceptual quality through block-based image segmentation, both serving as effective no-reference metrics. Other notable measures include FID and Inception Scores, which assess the quality and diversity of generated images, and SSIM (Wang et al. 2004) and PSNR, which quantify structural similarity and pixel-level accuracy. Beyond quantitative measures, human evaluations have also been leveraged to estimate natural image quality. For example, NIMA (Esfandarani and Milanfar 2018), proposed by Talebi and Milanfar, uses a deep CNN trained on human-rated images to predict aesthetic quality. Similarly, Wong et al. introduced the AVA dataset, which facilitates aesthetic assessment based on human preferences (Murray, Marchesotti, and Perronnin 2012).
AI-generated Image Evaluation:
Recently, AI-generated content (AIGC) has seen significant advancements with the rise of generative models. Several methods and benchmarks have been introduced to evaluate AI-generated images (Zhang et al. 2023; Hu et al. 2023). These methods primarily focus on three aspects: perceptual quality, text-image correspondence, and aesthetics (Xu et al. 2023a; Wang et al. 2023), often training unified models to assess the overall quality of AI-generated images. For example, ImageReward (Xu et al. 2023a) presents a framework for aligning image generation with human preferences, AGIQA-3k (Li et al. 2023a) offers a comprehensive benchmark for assessing AI-generated image quality, and QBench (Wu et al. 2024) evaluates content quality across multiple dimensions.
Despite advancements, little research has delved into the fine-grained differences between AI-generated and natural images, and existing datasets are often too small for comprehensive evaluation. To address these gaps, we propose an AI-Natural Image Difference Evaluation Benchmark, featuring a large Distinguishing AI-Natural Image Dataset and a systematic, fine-grained evaluation framework to thoroughly explore these differences.
Figure 1: Overview of AI-Natural Image Discrepancy Evaluation Benchmark (ANID): (a) A large Distinguishing AI-Natural Image (DANI) dataset is curated, guided by three types of prompts. (b) A fine-grained evaluation framework is proposed to measure and evaluate the potential discrepancies between AI-generated images and natural images from five aspects covering naive image quality, semantic alignment, aesthetic appeal, downstream applicability, and coordinated human assessment.
AI-Natural Image Discrepancy Evaluation Benchmark
To tackle the challenges of current AI-generated image evaluation, we construct an AI-Natural Image Discrepancy evaluation benchmark (ANID) to evaluate the potential discrepancies between AI-generated images and natural images. Figure 1 shows the overview of our ANID benchmark.
Our ANID benchmark contributes to two core components: (1) A Distinguishing AI-Natural Image (DANI) Dataset is constructed, in which we collect AI-generated images from diverse generative models for real natural images in MS COCO dataset (Lin et al. 2014) based on three types of guidance prompts e.g. text-only, image-only and text-image prompts; (2) We devise a systematic and comprehensive evaluation framework to measure and evaluate the potential differences between AI-generated images and real natural images from five aspects including naive image quality, semantic alignment, aesthetic appeal, downstream applicability, and the coordinated human assessment. In the following parts, we introduce the DANI Dataset and the evaluation framework in detail.
Distinguishing AI-Natural Image Dataset
We construct the Distinguished Natural and AI-generated Image Dataset based on the classical natural image dataset MS COCO dataset. We select 5,000 images covering diverse captions from the MS COCO dataset and collect a total of 25,000 text-image pairs by pairing each image with five different pieces of text descriptions. The 25,000 text-image pairs are seen as the referenced natural image set.
For each text-image sample in the referenced natural image set, we collect AI-generated images from 8 representative generative models, guided by three types of prompts, including text-only prompts, image-only prompts, and text-image prompts. The 8 representative generative models are from diverse generative models ranging from the earlier generative adversarial networks (GANs) (e.g. GALIP (Tao et al. 2023), DF-GAN (Tao et al. 2022)) to the recent benchmarking diffusion models (DMs) like Stable Diffusion v1.4, v1.5, v2.1, XL (Rombach et al. 2022), Versatile-Diffusion (Xu et al. 2023b) and OpenAI DALL·E 2 (Ramesh et al. 2022). The referenced natural images and the generated images comprise our distinguishing AI-Natural image dataset. More details about data construction and collection are presented in Section Appendix A.
Table 1 shows the statistics and properties of our DANI Dataset compared with the existing evaluation dataset. Our DANI Dataset has a total of 440,000 AI-generated images from 8 representative generated models. For details, please refer to Appendix Sec. A. Compared with state-of-the-art evaluation datasets, our DANI Dataset has convincing contributions:
  • DANI Dataset is the most extensive dataset for AI-generated image evaluation, and the data size scales to 100x.
  • DANI Dataset is guided by both unimodal prompts and multi-modal prompts (text-image). Prior datasets barely considered text prompts.
  • Diverse generative models are exploited, ranging from the earlier GAN models, the recent popular diffusion models, and the influential large commercial generative model DALL·E 2 issued by OpenAI.
Our large and comprehensive dataset works as the foundation of the evaluation benchmark and enables thorough evaluation and analysis of the differences between AI-generated images and realistic natural images.
Dataset Model Text Image Text vs. Image AIGI
AGIQA-1K 2 1,080 - - 1,080
AGIQA-3K 6 300 - - 2,982
AIGCIQA2023 6 100 - - 2,400
PKU-I2IQA 2 200 200 - 1,600
DNAI (Ours) 8 25,000 5,000 25,000 440,000
Table 1: Comparison of DANI dataset and Counterparts.
Fine-Grained Evaluation Framework
On top of the DANI dataset, we construct a fine-grained evaluation framework to systematically measure and interpret the differences between AI-generated images and realistic natural images from five different aspects, including naive image quality, semantic alignment, aesthetic appeal, downstream applicability, and coordinated human assessment.
1. Naive Image Quality:
We leverage traditional image quality assessment methods for natural images to measure the AI-Natural image differences. In order to perform fine-grained analysis and exploration, we estimate the AI-Natural image differences from three levels by considering the inherent properties of image-style content, ranging from the low-level visual features, frame-level visual features to the holistic content distribution.
  • First, we use the pixel-level similarity measures like SSIM, LPIPS, and DISTS to compute the visual similarities between the paired AI-generated image and natural image and use the inverse similarity values to quantify the AI-Natural image differences from the low-level visual view.
  • Second, we use frame-level visual features to qualify the perceptional quality of images and compare the global-level visual quality differences between AI-generated images and natural images from the high-level visual view. The exploited frame-level visual features include PIQE (N. et al. 2015), IL-NIQE (Zhang, Zhang, and Bovik 2015), MUSIQ (Ke et al. 2021), DBCNN (Zhang et al. 2020), LIQE (Ma et al. 2018), Inception Score (Salimans et al. 2016), CLIPIQA (Radford et al. 2021a), TReS (Gu et al. 2015), HyperIQA (Su et al. 2020), UNIQUE (Zhang et al. 2021), BRISQUE , NIQE (Mittal, Soundararajan, and Bovik 2013), NRQM (Ma et al. 2018), etc. These metrics provide insights into global-level visual quality differences.
  • Third, we investigate the visual content distribution to capture the holistic differences between AI-generated images and natural images by computing the FID and Inception Score of the whole AI-generated image dataset and the referenced natural image set.
Based on the different levels of image quality measures, our framework can evaluate the fine-grained AI-Natural image differences, covering both low-level and high-level. More details about the utilized traditional image quality assessment measures are described in Appendix Sec. C.
2. Semantic Alignment:
Generative models require the referenced prompts as semantic guidance to generate images. Therefore, semantic alignment can be an important quality indicator of the AIGIs. We use the widely used CLIP model (Radford et al. 2021b) to measure semantic alignment and compare the different CLIP Scores from natural text, image, and the paired AI-generated image-text prompt. The studied AI-generated images were originally generated from the 8 Representative generative models based on three types of prompts as described in Appendix Sec. B; the semantic alignment evaluation can provide a fine-grained analysis of the potential differences in semantic alignment with respect to both unimodal and multimodal guidance.
3. Aesthetic Appeal:
Aesthetic Appeal evaluation is to estimate the visual appeal and artistic quality of images, reflecting the visual attractiveness and artistic quality of images. We utilize the classical aesthetic measures NIMA (Esfandarani and Milanfar 2018) and LAION-AES (Schuhmann et al. 2022) as the quantitative metrics and compare the aesthetic scores for the paired AI-generated and natural images.
4. Downstream Applicability:
This aspect is devised to investigate the practical utility of AI-generated images in downstream tasks and evaluate whether AI-generated images can have different practical utilities with realistic natural images in downstream application tasks. We mainly focus on two classical downstream tasks, e.g., image recognition and object segmentation. For the image recognition task, we evaluate the Classification Mismatch Rate (MR) rates between AI-generated images and natural images when using a pre-trained image recognizer like ResNet-152 (He et al. 2016) model. For the object segmentation task, we use the Intersection over Union (IoU) metric (Everingham et al. 2010) as quantitative measures and investigate the segmentation results for AI-generated images with the segmentation model U2NET (Qin et al. 2020).
5. Human Assessment:
We involve human assessments to coordinate the above evaluation aspects. We develop the human assessment interface to demonstrate AI-generated images with the referenced natural images and text descriptions and collect human ratings (on a scale from 1 to 5) alongside the following three aspects: naive image quality, semantic alignment, and Aesthetic Appeal as human assessments. Human participants are provided with example images before the evaluation to understand high and low scores without being given specific numerical values, ensuring unbiased and informed ratings. Detailed information about human assessment procedures is provided in Appendix Sec. D.
Through the integration of diverse and comprehensive evaluation dimensions, our fine-grained evaluation framework offers a systematic assessment solution to investigate and interpret the differences that still remain between AI-generated images and natural images.
Figure 2: We evaluate eight image-generating models from 6 aspects. The numerical values in the radar chart represent the mean difference rate of each model.
Evaluation
With respect to the 5 different aspects, we conducted benchmark experiments and performed experimental evaluation alongside the following research questions.
RQ1: What are the fine-grained discrepancies between AIGIs and natural images?
  • RQ1.a: What discrepancies between I2I-guided AIGIs and their natural counterparts can be interpreted via the pixel-level quality measures?
  • RQ1.b: What discrepancies between multimodal-guided AIGIs and their natural counterparts can be revealed via the frame-level metrics?
  • RQ1.c: How do the structural visual content distributions differ between AIGIs and natural images?
RQ2: How significant are the discrepancies in semantic alignment between AIGIs and natural images in different types of guided prompts?
RQ3: What are the discrepancies in aesthetic appeal between AIGIs and natural images?
RQ4: How do AIGIs differ from natural images in downstream task applicability?
RQ5: Are human assessment results consistent with quantitative measures? What are the discrepancies revealed from human evaluation?
Experimental Setting
We use the described quantitative evaluation metrics, and details of all metrics are presented in Appendix Sec. B.
Full-reference metrics:
For full-reference metrics like pixel-level image quality, we directly report the calculated value cause it already can present the discrepancy.
No-reference metrics
We calculate Difference Rate (DR) for each quantitative metric in non-reference scenarios, by subtracting the value of each AI-generated image from its corresponding natural reference image and then averaging these difference rates, defined as follows:
Difference $\text{{Difference Rate}}=\frac{1}{n}\sum_{i=1}^{n}\frac{\left|m(\mathbf{X}_{i})-m(\mathbf{N}_{i})\right|}{m(\mathbf{N}_{i})}$ (1)
where $n$ is the number of images, $m$ represents the metric, $\mathbf{X}$ denotes the generated images, and $\mathbf{N}$ refers to the referenced natural images.
RQ1: Naive Quality Results
To evaluate the image quality of AI-generated images versus natural images, we conducted a comprehensive analysis using a variety of metrics that assess visual feature quality, naturalness, and similarity of images. Our findings reveal substantial discrepancies between AIGIs and natural images across these dimensions.
RQ1.a Pixel-level
We evaluated the visual similarity between AI-generated and natural images using structural metrics like SSIM, LPIPS, DISTS, and PSNR. The results, as shown in Table 2, indicate that AI-generated images exhibit significantly lower similarity to natural counterparts. AI-generated images show a 20% to 50% reduction in SSIM, highlighting a major loss in structural fidelity. LPIPS and DISTS further reveal notable perceptual dissimilarities, with LPIPS scores ranging from 0.19 to 0.60. Additionally, PSNR values, between 12.42 and 23.25, suggest that AI-generated images have higher noise levels, reducing their overall visual similarity.
SD_V14 SD_V15 Versatile SD_V21 SD_XL DALL·E 2
SSIM ↑ 0.35 0.35 0.46 0.44 0.70 0.38
PSNR ↑ 13.81 13.80 16.52 16.00 23.25 12.42
VIF ↑ 0.02 0.02 0.05 0.04 0.17 0.03
VSI ↑ 0.84 0.84 0.89 0.87 0.96 0.83
FSIM ↑ 0.61 0.61 0.71 0.67 0.86 0.61
LPIPS ↓ 0.60 0.60 0.39 0.52 0.19 0.61
DISTS ↓ 0.27 0.27 0.17 0.23 0.10 0.24
MAD ↓ 211.61 211.73 196.86 202.29 143.66 214.86
Table 2: Full-reference Evaluation Results (I2I).
RQ1.b Frame-level
We assessed the frame-level quality of AI-generated and natural images using a suite of metrics, including MUSIQ, DBCNN, HyperIQA, LIQE, BRISQUE, NRQM and NIQE. Figure 2 uses radar figures to visualize the results from the six metrics, and indicates that there are significant quality differences between AI-generated and natural images. These deviations typically range from 10% to 20%, irrespective of the guidance method used, with some instances showing quality differences as high as 30%. This underscores the considerable gaps between AIGIs and natural images in terms of global image quality. Moreover, the results of NRQM and NIQE reveal that AI-generated images significantly differ from natural images in terms of naturalness. Specifically, AI-generated images show notable deviations from natural images in visual realism and adherence to natural scene statistics. The NRQM difference rates range from 5% to 20%, while the NIQE difference rates range from 10% to 50%, highlighting the challenges AI-generated images face in replicating the inherent naturalness of real-world scenes. More detailed results with more metrics are presented in Appendix Sec. C.
Figure 3: The FID and Inception Score of AI-generated images from different source models and guidance.
RQ1.c Visual Content Distribution
We present the FID and Inception Score results, presented in Figure 3, suggesting that AI-generated images often exhibit relatively higher FID scores and lower Inception Scores. This finding indicates that current generative models, particularly GAN-based models and earlier versions of Stable Diffusion, struggle to consistently produce high-quality images. Notably, these models tend to generate better images when a reference image is provided (TI2I), as it has more detailed guidance.
In conclusion, our evaluation of image visual feature quality, naturalness, and similarity reveals significant disparities between AI-generated and natural images. These differences highlight the current limitations of AIGC technologies in matching the visual quality and realism of natural images.
RQ2: Alignment Results
In our evaluation of alignment results, we assessed how well AI-generated images correspond to their respective reference prompts across different guidance types: Text-to-Image (T2I), Image-to-Image (I2I), and Text-and-Image-to-Image (TI2I). And we find that AI-generated images often struggle to maintain high alignment with their prompts, especially the image-only reference provided. The results, shown in Table 3, indicate that AI-generated images typically achieve lower CLIP Score compared to natural images, with alignment discrepancies ranging from 4.02% to 36.69%. We also find that even if the AIGIs can get a higher CLIP score, especially with text and image-guided, they are looking stranger to humans as Table 3 illustrated. The CLIP score, which measures the alignment between text descriptions and generated images by embedding both into a shared space and calculating cosine similarity, revealed a noticeable gap in how accurately AIGC images reflect the input prompts. This suggests that current AIGC models often struggle to maintain high fidelity to the provided references, leading to variations that may impact the intended alignment. Our findings underscore the challenges that these models face in consistently generating content to accurately reflect the specified input conditions, highlighting both their strengths and limitations in real-world applications.
GALIP DFGAN SD_V14 SD_V15 Versatile SD_V21 SD_XL DALL·E 2
T2I 10.72 18.51 9.78 9.69 9.71 9.93 10.20 9.46
I2I - - 36.46 36.69 4.92 26.96 3.98 5.41
TI2I - - 9.21 9.20 8.03 8.48 4.02 -
Table 3: The difference rate of CLIP score (%).
RQ3: Aesthetic Appeal
We assessed the aesthetic appeal of AI-generated and natural images using NIMA and LAION-AES metrics and found that AI-generated images generally fall short of natural images in terms of aesthetic quality. Figure 4 presents the results, which reveal that AI-generated images receive lower scores for aesthetic appeal, with discrepancies ranging from 3.34% to 23.85%. These findings indicate that while AI models can produce visually appealing images, they often lack the nuanced artistic quality and emotional impact that are characteristic of natural images.
Figure 4: The aesthetic difference rate between AI-generated and naturally generated images.
RQ4: Applicability Results
When evaluating the applicability of AI-generated images, we focused on two crucial downstream tasks: image recognition and semantic segmentation. Our findings indicate significant discrepancies between the performance of AIGC images and natural images in these tasks.
Image Recognition:
For image recognition, we measured the classification mismatch rate, which quantifies the differences in predicted labels between AIGC images and their natural counterparts. The results in Table 4 showed a substantial gap, with the mismatch rate ranging from 29.89% to 94.44%. This indicates that AI-generated images often fail to achieve the same level of accuracy as natural images, highlighting the limitations of current AIGC models in producing reliable content for image recognition tasks. The high mismatch rate underscores the need for further advancements in generative models to improve their applicability in real-world recognition tasks.
Overall, our results indicate that AI-generated images currently exhibit significant limitations in applicability for both image recognition and semantic segmentation tasks. The high classification mismatch rates and variable mean IoU values highlight the gap between AIGC and natural images in practical applications. These findings emphasize the importance of ongoing research and development to enhance the reliability and accuracy of AIGC models, ensuring they can be effectively used in real-world scenarios.
GALIP DFGAN SD_V14 SD_V15 Versatile SD_V21 SD_XL DALL·E 2
T2I 72.02 87.43 94.44 94.29 43.72 84.42 30.41 69.74
I2I - - 68.62 68.33 70.73 68.01 67.73 53.42
TI2I - - 62.84 62.35 62.28 58.56 29.89 -
Table 4: The results of Classification Mismatch Rate (MR).
GALIP DFGAN SD_V14 SD_V15 Versatile SD_V21 SD_XL DALL·E 2
T2I 0.23 0.23 0.25 0.25 0.26 0.25 0.26 0.26
I2I - - 0.38 0.38 0.69 0.50 0.82 0.42
TI2I - - 0.43 0.44 0.30 0.53 0.82 -
Table 5: The results of Segmentation IoU.
Semantic Segmentation:
In semantic segmentation, we evaluated the performance using the mean Intersection over Union (IoU) metric and show the results in Table 5. The mean IoU values for AIGC images varied between 0.23 and 0.82, indicating a wide range of segmentation accuracy. While some generated images approached the performance of natural images, a significant portion of AIGIs fell short, demonstrating poor segmentation results. This variability suggests that while AIGC models can sometimes produce high-quality segmentation, they often struggle to consistently match the accuracy of natural images.
Visual Question Answering:
To further evaluate the performance of AI-generated images and natural images in downstream tasks, we conducted tests on the VQA model BLIP-2 (Li et al. 2023b). The evaluation involved three different types of questions: (1) identifying objects in the image (Object), (2) determining whether the image aligns with a given text by answering ”yes” or ”no” (Align), and (3) predicting the similarity probability between the image and the text (Prob.). Table 6 summarizes the discrepancies (%) between natural and AI-generated images across various models. For Object questions, the highest discrepancies occurred in the I2I setting, with SD_V14 and SD_V15 reaching 77.27% and 75.51%, respectively, while T2I discrepancies ranged from 36.08% (DALL·E 2) to 68.96% (DFGAN). Align questions showed lower discrepancies, with T2I values ranging from 24.72% (Versatile) to 34.65% (DFGAN) and I2I values from 15.90% (Versatile) to 29.24% (SD_V15). For similarity probability (Prob.), T2I discrepancies varied between 36.61% (DALL·E) and 62.01% (DFGAN), while I2I results ranged from 31.03% (DALL·E) to 62.80% (SD_V14), highlighting model-dependent variations.
These results indicate that AI-generated images, especially those produced under certain generative settings, face significant challenges in downstream VQA tasks compared to natural images. The findings also reveal performance variations across generative models, underscoring the need for improved alignment and semantic understanding in AI-generated images.
GALIP DFGAN SD_V14 SD_V15 Versatile SD_V21 SD_XL DALL·E 2
Object T2I 47.84 68.96 39.70 38.97 43.46 38.50 38.42 36.08
I2I - - 77.27 75.51 28.78 71.59 28.37 26.03
TI2I - - 41.18 38.06 37.64 39.27 26.69 -
Align T2I 29.81 34.65 27.52 28.03 24.72 31.99 29.78 29.79
I2I - - 28.48 29.24 15.90 27.07 16.33 25.50
TI2I - - 27.52 25.79 19.26 27.31 15.85 -
Prob. T2I 54.61 62.01 54.74 52.69 49.67 54.84 58.68 36.61
I2I - - 62.80 58.39 36.16 56.14 42.88 31.03
TI2I - - 52.75 48.71 43.69 46.30 43.24 -
Table 6: The results of VQA task w.r.t Object, Alignment, and Similarity probability.
SD_V21 SD_XL Versatile DALL·E 2
T2I I2I TI2I T2I I2I TI2I T2I I2I TI2I T2I I2I
Quality 3.54 3.68 2.88 3.44 2.40 4.39 4.11 3.49 3.31 4.37 3.32
Alignment 3.45 3.43 2.48 3.16 2.44 4.47 4.22 3.59 3.40 4.50 2.96
Aesthetic 3.38 3.55 2.70 3.21 2.35 4.29 3.96 3.32 3.19 4.28 3.23
Table 7: Human evaluation results.
Figure 5: The difference rate provided by human & quantitative statistics, where ‘H’ represents human and ‘Q’ represents quantitative.
Human Validation
To rigorously assess the differences between AI-Natural images and to evaluate the effectiveness of existing metrics in measuring AIGC image quality, we conducted human evaluations focusing on three key aspects: Image Quality, Alignment, and Aesthetic Appeal. We excluded results for Applicability since this dimension inherently involves human-generated ground truth labels. Participants rated images on a scale from 1 (poor) to 5 (excellent). Across all three dimensions, AI-generated images generally received lower scores, often between 2 and 3, compared to natural images.
Image Quality: Participants assessed the overall quality of the images, considering factors such as clarity, detail, and the presence of artifacts. The results revealed a noticeable gap between AIGC images and their natural counterparts, with the generated images often lacking the same level of detail and clarity.
Alignment: This aspect measured how well the generated images matched the given prompts, revealing a notable discrepancy. Generated images frequently deviated from the descriptions, indicating issues with model adherence to specified conditions.
Aesthetic: The aesthetic appeal of the images was rated based on visual attractiveness, composition, and overall artistic quality. While some generated images displayed impressive artistic qualities, the overall consistency and aesthetic appeal were generally considered inferior to that of natural images.
Overall, our human evaluation results are shown in Table 7 and Figure 5 reveal significant discrepancies between AI-generated images and natural images across all three assessed aspects: quality, alignment, and aesthetic appeal. AIGC images consistently received lower scores, emphasizing the need for further advancements to narrow these gaps. These findings underscore the importance of continued research and development in AI image generation to bring AIGC images closer to the standards of natural images.
Moreover, the results from human validation indicate that traditional quantitative metrics often fail to align with human preferences, suggesting that they are not fully compatible with assessing AIGC images. This highlights the necessity for developing new measures tailored specifically to the unique characteristics of AI-generated content.
More Evaluations
To analyze the differences between AI-generated and natural images across various categories, we evaluated the performance of generated data from eight categories (’person’, ’animal’, ’indoor’, ’outdoor’, ’vehicle’, ’food’, ’sports’, ’accessory’) across three key aspects: (1) Naive Quality, including pixel-level similarity (SSIM), frame-level quality (CLIPIQA), and content distribution (FID); (2) Alignment, measured by CLIP Score; and (3) Aesthetic Quality, assessed by LAION-AES. Partial results are presented in Figures 6 and 7, while the full experimental findings are provided in Appendix Sec. E (Figures 13 and 14 ).
From Figure 6, we observe substantial variability in the differential rates for categories such as “food” and “sports”. Compared to “sports” images, the “food” demonstrate significantly higher discrepancies in image quality (CLIPIQA), revealing AI models struggle to generate natural food images. In Figure 7, the FID and SSIM metrics highlight distinct trends across categories. The “accessory” category exhibits the highest FID values, indicating poorer content distribution, while the “animal” category consistently achieves higher SSIM scores, reflecting better pixel-level similarity. These results suggest that certain categories, such as “accessory” and “indoor”, remain particularly challenging for AIGC models to reproduce accurately, especially in maintaining semantic consistency and achieving balanced content generation.
Overall, these findings demonstrate that AI-generated images still fall short of matching natural images in terms of naturalness and image quality. Additionally, the pronounced performance discrepancies across categories underscore an imbalance in the training data used for AIGC models. Addressing these discrepancies will require more balanced datasets and targeted improvements in model architecture and training strategies.
Figure 6: Differential rates of AI-generated images in the food and sports categories across naive image quality, alignment, and aesthetic metrics.
Figure 7: FID and SSIM metrics for AI-generated images across different source models and categories. FID evaluates distribution-level image quality, while SSIM measures pixel-level similarity to assess overall image quality.
Conclusion and Discussion
In this paper, we develop a systematic and comprehensive framework to evaluate the discrepancies between AI-Natural images using the formulated DNAI dataset. Our findings reveal significant disparities in naive image quality, semantic alignment, aesthetics, and downstream task applicability. While AIGC models have achieved impressive innovations, AIGIs still lag behind natural images in these critical areas. Besides, human evaluations also reveal lower scores for AIGC in terms of visual quality, semantic alignment, and aesthetics, emphasizing the gaps between AIGIs and natural images. Finally, the poor utility of AIGIs in downstream tasks indicates practical application challenges for AIGC.
This study also has limitations, including reliance on specific quantitative metrics and models available at the time. Future studies should incorporate more generative models, larger datasets, and sophisticated metrics to capture further discrepancies. Addressing these limitations and improving models will help reduce the perceptual gaps between AIGC images and natural images, facilitating better integration into real-world applications.
Appendix
Appendix Overview
The appendix provides supplementary details and additional experimental results that were not included in the main paper due to space limitations. It is organized as follows:
  • Section A: In-depth Description of the DNAI Dataset.
  • Section B: Detailed Overview of the Evaluated Generative Models.
  • Section C: Comprehensive Summary of Evaluation Metrics.
  • Section D: Extended Information on Human Perceptual Evaluations.
  • Section E: Discussion on the Safety Mechanisms in Existing Diffusion Models.
A Details of the DNAI Dataset
Overview
The DNAI dataset is a carefully curated collection of both natural and AI-generated images, designed to support comprehensive evaluations of image generation models. The natural images are sourced from the COCO validation set. AI-generated images were produced using eight different models. These images were generated under three distinct guidance modes: Text-to-Image (T2I), Image-to-Image (I2I), and Text vs. Image-to-Image (TI2I). Representative samples from the DNAI dataset are shown in Figure 8. For a detailed breakdown of the DNAI dataset composition, please refer to Table 8.
Figure 8: Examples of AI-generated content (AIGC) created using five state-of-the-art generative models: SD_v1.4, SD_v1.5, Versatile Diffusion, SD_v2.1, and SD_XL. Each case is illustrated with three rows corresponding to Text-to-Image (T2I), Image-to-Image (I2I), and Text-and-Image-to-Image (TI2I) generation modes. For the complete set of AIGC images, please refer to our DNAI dataset, which will be available publicly soon.
Types DFGAN GALIP SD_V1.4 SD_V1.5 Versartile SD_V2.1 SD_XL DALL·E 2
T2I 25,000 25,000 25,000 25,000 25,000 25,000 25,000 5,000
I2I - - 25,000 25,000 25,000 25,000 25,000 5,000
TI2I - - 25,000 25,000 25,000 25,000 25,000 -
Table 8: The details of the DNAI dataset.
Natural Images
The DNAI dataset features a carefully curated selection of natural images sourced from the COCO validation set. This subset comprises 5,000 images, each paired with at least five unique captions, totaling 25,014 captions. These rich and varied textual descriptions enhance the dataset’s utility across a wide range of image-to-text and text-to-image tasks, making it an invaluable resource for evaluating and training image generation and understanding models. In this study, we use these images and their accompanying (five) captions to guide the generative models in creating new images.
Generated Images
The DNAI dataset includes AI-generated images created using several advanced generative models across three distinct guidance modes: Text-to-Image (T2I), Image-to-Image (I2I), and Text-and-Image-to-Image (TI2I). These modes collectively contributed a significant volume of images to the dataset. The models utilized in this process include DF-GAN, GALIP, various versions of Stable Diffusion (v1.4, v1.5, v2.1, XL), Versatile Diffusion, and DALL·E 2. Each model, except DALL·E 2, generated 25,000 images per guidance mode. DALL·E 2 produced 5,000 images per mode, resulting in a total of 440,000 generated images. It is important to note that DF-GAN and GALIP were used only for T2I, while DALL·E 2 was employed for both T2I and I2I modes. For detailed information, please refer to Table 8.
Image Resolution
The resolution of the generated images varies depending on the model used. Images generated by GALIP are 224x224 pixels, while those produced by DF-GAN are 256x256 pixels. The other models, including various versions of Stable Diffusion and DALL·E 2, generate images at a resolution of 512x512 pixels.
Dataset Composition
The DNAI dataset offers a robust foundation for evaluating the performance and capabilities of various AIGC models under different conditions and guidance modes. By encompassing a broad spectrum of images and resolutions, the dataset ensures comprehensive coverage of the potential use cases and challenges associated with AI-generated content. The combination of natural and generated images, alongside diverse textual descriptions, facilitates a thorough assessment of image generation models, supporting the development of more accurate and reliable AI systems.
B Evaluated Generative Models
GALIP (Generative Adversarial Latent Image Processing): (Tao et al. 2023) is a generative model that produces high-quality images by leveraging adversarial techniques. It utilizes latent image processing to enhance both the quality and fidelity of generated content. GALIP’s architecture involves a generator and discriminator, where the generator creates images from latent vectors, and the discriminator evaluates their authenticity. Through iterative training, GALIP achieves highly detailed and realistic image synthesis.
DFGAN (Deep Fusion Generative Adversarial Network): (Tao et al. 2022) focuses on generating images from text descriptions using deep fusion techniques. The model integrates textual information into the image generation process through multiple layers of fusion, ensuring alignment with input descriptions. DFGAN operates in two stages: the first generates a coarse image based on the text, and the second refines the image, enhancing detail and coherence.
Stable Diffusion v1.4 (SD_V14): (Rombach et al. 2022) is part of the Stable Diffusion family, designed for image synthesis through iterative diffusion processes. Starting with pure noise, the model refines it into coherent images through a series of steps, guided by a learned model that predicts noise distribution. SD_V14 is recognized for producing high-resolution images with fine textures and complex structures.
Stable Diffusion v1.5 (SD_V15): (Rombach et al. 2022) builds on SD_V14, introducing enhancements in architecture and training techniques to improve image quality, consistency, and the handling of complex scenes. While continuing to use the diffusion process, SD_V15 incorporates better optimization strategies and larger datasets, yielding superior results.
Versatile Diffusion: (Xu et al. 2023b) is a diffusion-based model designed for a variety of image generation tasks, including style transfer, image inpainting, and super-resolution. Its versatility lies in its ability to adapt the diffusion process to the specific requirements of each task, making it suitable for diverse applications.
Stable Diffusion v2.1 (SD_V21): (Rombach et al. 2022) is an advanced iteration of the Stable Diffusion series, offering significant improvements in image fidelity and generation speed. With optimizations in the diffusion process, better noise handling, and enhanced training algorithms, SD_V21 produces more realistic and high-quality images, making it highly effective for a range of creative and practical uses.
Stable Diffusion XL (SD_XL): (Rombach et al. 2022) represents the most advanced model in the Stable Diffusion series, capable of generating extremely high-resolution images with intricate details. SD_XL uses an extended number of diffusion steps and a larger network architecture to manage increased complexity, making it ideal for applications requiring ultra-high resolution and precision, such as detailed artworks and large-scale prints.
DALL·E 2: (Ramesh et al. 2022) is a generative model developed by OpenAI that excels in creating images from textual descriptions. As an advancement over the original DALL·E, it offers improved capabilities in generating high-quality, diverse, and coherent images. DALL·E 2 combines CLIP (Contrastive Language-Image Pre-training) with a transformer-based generative model to translate complex and abstract textual inputs into visual content, making it renowned for its creative and realistic outputs.
These generative models represent significant advancements in AI-driven image synthesis, each with unique strengths tailored to specific applications. From text-to-image generation and high-resolution synthesis to versatile image processing, these models push the boundaries of AI-generated content.
Figure 9: We evaluate eight image-generating models from 13 aspects. The numerical values in the radar chart represent the mean difference rate of each model.
C Evaluation Metrics
Overview
In this study, we selected a comprehensive set of metrics to evaluate the image quality, alignment, and aesthetics of both natural and AI-generated images. These metrics, summarized in Table 9, are carefully chosen to highlight the differences across multiple dimensions, providing a robust framework for assessing the quality and realism of AIGC in comparison to natural images.
To evaluate image quality, we divided the assessment into three sub-aspects: pixel-level similarity, frame-level visual metrics, and visual content distribution.
  • Pixel-level: Metrics such as PSNR, LPIPS, DISTS, VIF, VSI, FSIM, and MAD assess structural, perceptual, and textural fidelity. These metrics measure how closely the generated images match natural ones in terms of structural integrity and perceptual similarity.
  • Frame-level: Metrics including PIQE, IL-NIQE, MUSIQ, DBCNN, CNNIQA, CLIPIQA, BRISQUE, TReS, HyperIQA, LIQE, and UNIQUE provide no-reference assessments of image quality. They evaluate attributes such as distortion levels, natural scene statistics, and learned features to predict the overall quality of the images. Additionally, naturalness metrics like NIQE and NRQM help gauge how closely AI-generated images resemble natural scenes based on statistical properties.
  • Visual Content Distribution: Metrics such as FID and Inception Score examine overall content distribution differences between AI-generated and natural images across the entire dataset. These metrics capture broad differences in content quality.
For alignment, we utilize the CLIP Score (Radford et al. 2021b) to measure the correspondence between text descriptions and the generated images. This metric evaluates how well the generated content aligns with the given prompts by embedding both text and image into a shared space and calculating their similarity.
In terms of aesthetic evaluation, we employ NIMA and LAION-AES metrics to assess the visual appeal and artistic quality of the images. These metrics predict human perception of image aesthetics based on deep learning models and various visual features.
By integrating these diverse metrics, our framework offers a detailed and multi-faceted evaluation of AI-generated images, capturing the nuances and differences between natural and synthetic content. The comprehensive nature of these metrics ensures a robust and reliable assessment, guiding improvements in AIGC technologies.
The chosen metrics, as detailed in Table 9, collectively provide a comprehensive toolkit for evaluating image quality, alignment, and aesthetics, each focusing on different aspects of visual perception and fidelity.
Naive Image Quality
We evaluated the naive image quality of AI-generated images versus natural images using a range of metrics. These metrics assess both pixel-level and frame-level qualities, as well as visual content distribution, to capture the holistic differences between AI-generated and natural images.
Pixel-Level Quality Metrics:
  • SSIM (Wang et al. 2004) (Structural Similarity Index) compares the structural information between a reference and a test image, considering luminance, contrast, and structure for a comprehensive measure of similarity.
  • PSNR (Peak Signal-to-Noise Ratio) measures the ratio between the maximum possible power of an image and the power of corrupting noise, with higher values indicating better image quality.
  • VIF (Sheikh and Bovik 2006) (Visual Information Fidelity) quantifies the amount of visual information preserved in the test image relative to the reference, based on natural scene statistics and the human visual system.
  • VSI (Zhang, Shen, and Li 2014) (Visual Saliency-based Index) evaluates perceptual similarity by focusing on how salient (prominent) features in the images compare.
  • FSIM (Zhang et al. 2011) (Feature Similarity Index) measures similarity using phase congruency and gradient magnitude, emphasizing feature similarity, which is crucial for human perception.
  • LPIPS (Zhang et al. 2018) (Learned Perceptual Image Patch Similarity) uses deep network features to assess perceptual similarity, capturing human visual perception more effectively by comparing feature activations in a deep neural network.
  • DISTS (Ding et al. 2022) (Deep Image Structure and Texture Similarity) combines structural and textural similarity metrics using deep learning features to evaluate overall image similarity.
  • MAD (Larson and Chandler 2010) (Mean Absolute Deviation) provides a straightforward assessment of pixel-level differences, with lower values indicating greater similarity between images.
Frame-Level Quality Metrics:
  • PIQE (N. et al. 2015) (Perception-based Image Quality Evaluator) provides a no-reference quality assessment by analyzing image blocks and estimating distortion levels.
  • IL-NIQE (Zhang, Zhang, and Bovik 2015) (Integrated Local Natural Image Quality Evaluator) evaluates image quality based on natural scene statistics and local features, offering a no-reference assessment.
  • MUSIQ (Ke et al. 2021) (Multi-Scale Image Quality) assesses image quality at multiple scales, considering both local and global features for a comprehensive no-reference evaluation.
  • DBCNN (Zhang et al. 2020) (Deep Bilinear Convolutional Neural Network) predicts image quality using deep learning, based on bilinear pooling of convolutional features.
  • LIQE (Ma et al. 2018) (Learning-based Image Quality Evaluator) combines traditional quality metrics with learned features for robust no-reference quality assessment.
  • CNNIQA (Kang et al. 2014) (Convolutional Neural Network Image Quality Assessment) leverages deep convolutional networks for no-reference image quality assessment using learned features.
  • CLIPIQA (Radford et al. 2021a) (CLIP Image Quality Assessment) uses CLIP model embeddings to assess image quality by comparing the alignment between image and text descriptions.
  • TReS (Radford et al. 2021a) (Textural and Edge-based Similarity) evaluates image quality by focusing on textural and edge-based features, providing a no-reference assessment.
  • HyperIQA (Su et al. 2020) (Hyper Network-based Image Quality Assessment) uses a hypernetwork to predict image quality, leveraging deep learning for no-reference quality assessment.
  • UNIQUE (Zhang et al. 2021) (Universal Quality Index with Deep Features) evaluates unique aspects of image quality using deep learning features, offering a no-reference assessment.
  • BRISQUE (Mittal, Moorthy, and Bovik 2012) (Blind/Referenceless Image Spatial Quality Evaluator) assesses image quality based on natural scene statistics, providing a no-reference measure.
  • NIQE (Mittal, Soundararajan, and Bovik 2013) (Natural Image Quality Evaluator) evaluates how natural an image appears by comparing its features with those derived from a dataset of natural images, without needing a reference image.
  • NRQM (Ma et al. 2018) (Naturalness Image Quality Metric) assesses how closely an image aligns with the statistical properties of natural scenes based on natural scene statistics and perceptual models.
Visual Content Distribution Metrics:
  • FID (Heusel et al. 2017) (Fréchet Inception Distance) measures the similarity between two sets of images by comparing the means and covariances of their feature representations, providing a comprehensive assessment of both quality and diversity.
  • Inception Score (Salimans et al. 2016) evaluates the quality of generated images by considering the confidence of the classifier and the diversity of the generated samples, with higher scores indicating better quality and diversity.
We present the additional results of the Naive Image Quality assessment in Figure 9.
Alignment
To assess alignment, we employ the CLIP Score (Contrastive Language-Image Pre-training Score), which measures the correspondence between text descriptions and generated images. This metric evaluates how well the visual content aligns with the textual prompts by embedding both into a shared space and calculating their cosine similarity.
Aesthetic
  • NIMA (Esfandarani and Milanfar 2018) (Neural Image Assessment) is a deep learning-based model that predicts the aesthetic quality of images. It generates a score that reflects human perception of visual appeal.
  • LAION-AES (Schuhmann et al. 2022) (LAION Aesthetic Score) evaluates images’ aesthetic quality using various visual features, providing a numerical value to indicate their aesthetic appeal.
Together, these metrics offer a robust toolkit for evaluating images, focusing on alignment with prompts and aesthetic quality, thereby covering key aspects of visual perception and fidelity.
Aspects Sub_aspects Netrucs Reference
Naive Quality Pixel-level SSIM TRUE
PSNR TRUE
LPIPS TRUE
DISTS TRUE
VIF TRUE
VSI TRUE
FSIM TRUE
MAD TRUE
Frame-level PIQE FALSE
IL-NIQE FALSE
MUSIQ FALSE
DBCNN FALSE
CNNIQA FALSE
CLIPIQA FALSE
BRISQUE FALSE
TReS FALSE
HyperIQA FALSE
LIQE FALSE
UNIQUE FALSE
NIQE FALSE
NRQM FALSE
Content Distribution FID TRUE
Inception Score TRUE
Alignment CLIP Score FALSE
Aesthtic NIMA FALSE
LAION-AES FALSE
Table 9: The detail of evaluation metrics.
D Human Perceptual Evaluation
Interface for Human Evaluation
The human evaluation of our DNAI dataset was conducted using a custom-designed interface (Figure 11). This interface allowed participants to evaluate AI-generated images in comparison to reference content across three key aspects: Quality, Alignment, and Aesthetics. The reference content included prompts used to generate the images, categorized into three types: 1) text, 2) image, and 3) text and image. The AI-generated images evaluated were produced by four different models: SD_V2.1, SD_XL, Versatile Diffusion, and DALL·E 2.
Evaluation Process
A total of 58 volunteers participated in the evaluation process. Each volunteer assessed the images based on the provided reference content. To ensure the validity of the data collected, we included duplicate images in the evaluation system. By comparing the scores given to these identical images, we could determine the consistency of each participant’s evaluations. Only data from participants who showed consistent scoring were considered valid, resulting in 47 effective samples.
Participant Demographics
The volunteers who participated in the evaluation ranged in age from 19 to 46 years old, with a gender ratio of 6:4 (male to female). All participants held at least an undergraduate degree, ensuring a well-educated sample group. Alongside scoring data, we collected demographic details such as age and gender. Upon completing their evaluations, participants were also invited to provide feedback on the evaluation system and their experience with assessing the differences between natural and AI-generated images.
Evaluation Metrics
Participants were instructed to evaluate each image based on three aspects:
  1. 1.
    Quality: Overall visual fidelity and clarity of the image.
  2. 2.
    Alignment: Accuracy in matching the image to the provided prompt.
  3. 3.
    Aesthetic: Visual appeal and artistic quality of the image.
To facilitate the scoring process, we provided reference examples within the evaluation system (Figure 12). These examples illustrated what constituted good or poor performance in each aspect without giving specific scores, thereby preventing any potential bias in participant evaluations.
Summary
The comprehensive human evaluation process ensured that the collected data was reliable and reflected genuine human perception. The demographic diversity and the rigorous validation of participant responses provided a robust foundation for assessing the performance of various AIGC models in generating high-quality, aligned, and aesthetically pleasing images. This approach not only highlighted the strengths and weaknesses of the evaluated models but also offered valuable insights into the perceptual differences between AI-generated and natural images.
Figure 10: Examples for unsafe references.
Figure 11: Interface for Human Evaluation. Reference content: the prompt guidance used to generate images, including three types: 1) text, 2) image, and 3) text and image. AIGC images: AI-generated images that come from four different models. Participants can rely on the reference content to evaluate each image on three aspects, i.e., Quality, Alignment, and Aesthetic.
Figure 12: Before the evaluation process, the participant will be given some examples to learn which image should be good or bad in a specific aspect.
E More Evaluations
This section provides a detailed evaluation of AI-generated images across different categories, focusing on naive image quality, alignment, aesthetics, pixel-level similarity, and content distribution. These analyses aim to uncover discrepancies between AI-generated and natural images and identify category-specific challenges.
From Figure 13, we observe notable variability in the differential rates across categories such as ”indoor” and ”sports” among different generative models. Images in the ”indoor” category exhibit significantly higher discrepancies in alignment (CLIP Score) and aesthetics (LAION-AES) compared to ”sports”, indicating that AI-generated sports images are more aligned with natural images. In contrast, generating natural-looking indoor images remains a challenge due to their complex structures and diverse visual features.
Figure 14 presents trends in SSIM and FID metrics across categories and models, where SSIM evaluates pixel-level similarity, and FID measures distribution-level content alignment. The ”accessory” category consistently shows higher FID values, reflecting poorer content distribution, while the ”animal” category achieves higher SSIM scores, demonstrating better pixel-level similarity. These findings suggest that ”accessory” and ”indoor” images pose greater challenges for generative models in maintaining semantic consistency and achieving balanced content generation compared to other categories.
Overall, these results indicate that AI-generated images still lag behind natural images in achieving comparable levels of naturalness and image quality. Moreover, the substantial performance discrepancies across categories highlight an imbalance in the training data used for generative models. This imbalance likely limits their ability to generalize effectively across diverse content types, underscoring the need for more balanced datasets and targeted advancements in model design.
Figure 13: The differential rate of data from different categories models across naive image quality, alignment, and aesthetic.
Figure 14: The FID and SSIM of AI-generated images from different source models and categories. The FID and SSIM are used to evaluate image quality at the pixel level and content distribution level.
F Discussion
Safety Mechanism Activation in SD 1.4, 1.5 DALL·E 2
T2I I2I TI2I T2I(%) I2I(%) TI2I(%)
SD v1.4 104 263 141 4.16 10.51 5.64
SD v1.5 133 270 145 5.32 10.79 5.80
DALL·E 2 20 - - 4 - -
Table 10: Unsafe Prompt.
In the rapidly evolving field of AI-generated content (AIGC), the balance between model robustness and safety mechanisms remains a critical area of study. A noteworthy issue has emerged with popular image synthesis models, such as Stable Diffusion v1.4, v1.5, and DALL·E 2, which occasionally fail to generate images in response to seemingly benign prompts. Some examples of these prompts are shown in Figure 10. This unexpected behavior suggests an oversensitivity in the models’ safety filters, which are intended to prevent the creation of inappropriate or sensitive content.
This case study delves into instances where typical prompts, which ostensibly do not contain objectionable content, nonetheless trigger these safety mechanisms. Through a series of experiments, we systematically presented a variety of benign prompts to both model versions and documented the conditions under which it is hard to output the AIGC images (Table 10). Our findings suggest that certain benign text, benign image, or combinations of text and image are misinterpreted by the models’ safety algorithms as being potentially harmful or sensitive.
The implications of these findings are twofold. First, they highlight a critical need for refining the sensitivity of safety algorithms to reduce false positives, thereby ensuring that the AIGC models do not unduly limit creative expression or practical application. Second, they serve as a stark reminder of the challenges inherent in balancing content safety with the functional robustness of generative models. This case study aims to contribute to the ongoing discussion on optimizing safety mechanisms in AIGC models, striving to enhance their utility and accessibility without compromising essential safeguards.

### v2

@@
-ANID: How Far Are We? Evaluating the Discrepancies Between AI-synthesized Images and Natural Images through Multimodal Guidance
+D-Judge: How Far Are We? Accessing the Discrepancies Between AI-synthesized Images and Natural Images through Multimodal Guidance
 Abstract
-In the rapidly evolving field of Artificial Intelligence Generated Content (AIGC), one of the key challenges is distinguishing AI-synthesized images from natural images.
-Despite the remarkable capabilities of advanced AI generative models in producing visually compelling images, significant discrepancies remain when these images are compared to natural ones.
-To systematically investigate and quantify these discrepancies, we introduce an AI-Natural Image Discrepancy Evaluation benchmark aimed at addressing the critical question: how far are AI-generated images (AIGIs) from truly realistic images?
-We have constructed a large-scale multimodal dataset, the Distinguishing Natural and AI-generated Images (DNAI) dataset, which includes over 440,000 AIGI samples generated by 8 representative models using both unimodal and multimodal prompts, such as Text-to-Image (T2I), Image-to-Image (I2I), and Text vs.
-Image-to-Image (TI2I).
-Our fine-grained assessment framework provides a comprehensive evaluation of the DNAI dataset across five key dimensions: naive visual feature quality, semantic alignment in multimodal generation, aesthetic appeal, downstream task applicability, and coordinated human validation.
-Extensive evaluation results highlight significant discrepancies across these dimensions, underscoring the necessity of aligning quantitative metrics with human judgment to achieve a holistic understanding of AI-generated image quality.
-Code is available at https://github.com/ryliu68/ANID.
-Introduction
-With the rapid advancement of deep learning techniques, the proliferation of Artificial Intelligence Generated Content (AIGC) has garnered significant attention across various domains, such as e-commerce, gaming, medicine, animation, and autonomous driving (Li et al.
-2024; Qian et al.
-2024).
-AI-generated Images (AIGI) have been one of the mainstream forms of AIGC, and a variety of AI image generative models have been proposed to make the generated synthetic images as realistic as natural images, ranging from the earlier generative adversarial networks (GANs) (Tao et al.
-2023, 2022), advanced diffusion models (DMs) (Xu et al.
-2023b; Wei et al.
-2023) to the large multimodal generative model like DALL·E (Ramesh et al.
-2022).
-Despite the proliferation of AI image generative models, AI-generated images are not qualified for real-world applications, and the discrepancies between generated synthetic images and realistic natural images still exist (Li et al.
-2023a).
-Therefore, many AI-generated image quality assessment and evaluation methods have been proposed (Wang et al.
-2023; Xu et al.
-2023a; Hu et al.
-2023).
-For example, Wang et al.
-(Wang et al.
-2023) established an AIGC Image Quality Assessment Database called AIGCIQA2023 from 6 text-to-image generative models, and devised a subjective evaluation framework to assess human visual preferences for each AI-generated image from three aspects, e.g., quality, authenticity, and text-to-image correspondence.
-PKU-I2IQA (Yuan et al.
-2023) constructed a human perception-based image-to-image database named PKU-I2IQA and conducted a subject analysis based on both no-reference and full-reference methods.
-Different from prior studies focusing on establishing the AIGI dataset and conducting a subjective evaluation with human feedback, another group of studies (e.g.
-Pick-a-Pic (Kirstain et al.
-2023), HPS v2 (Wu et al.
-2023), TIFA (Hu et al.
-2023), and ImageReward (Xu et al.
-2023a)) focused on training a unified score model as the automatic evaluation metrics for AIGC image quality assessment (AIGIQA) measures, in contrast with traditional image quality assessment for natural images measures like Inception Score (Salimans et al.
-2016) and FID (Heusel et al.
-2017).
-However, some issues and problems have not yet been fully answered and solved.
-First, prior studies only considered unimodal promopted generation, e.g.
-Text-to-Image (T2I) or Image-to-Image (I2I), ignoring multimodal prompted content like Text vs.Image-to-Image (TI2I).
-The data size of the established T2I or I2I dataset is relatively small (around thousands of images), and might be insufficient for comprehensive analysis and evaluation.
-Second, prior studies typically focused on the perceptual quality of AIGC images with traditional image quality assessment metrics, the text-to-image correspondence, and subjective human preference.
-No studies have fully answered the important questions about what discrepancies exist between generated synthetic images and realistic natural images, and they lacked a systematic and compressive assessment and evaluation to investigate and interpret the discrepancies between AI-generated synthetic images and realistic natural images.
-The investigation and understanding of the discrepancies are key ways to realize the practical application of AIGC images in real-world scenarios and make breakthroughs in AIGC.
-To address the above challenges, we proposed an AI-Natural Image Discrepancy Evaluation Benchmark to investigate and interpret what discrepancies still remained between AI-generated images and natural images, finally answering the important question “How far are AI generative models with respect to the visual forms of images?”.
+In Artificial Intelligence Generated Content (AIGC), distinguishing AI-synthesized images from natural ones remains a key challenge.
+Despite advancements in generative models, significant discrepancies persist.
+To systematically investigate and quantify these discrepancies, we introduce an AI-Natural Image Discrepancy accessing benchmark (D-Judge) aimed at addressing the critical question: how far are AI-generated images (AIGIs) from truly realistic images?
+We construct D-ANI, a dataset with 5,000 natural images and over 440,000 AIGIs generated by nine models using Text-to-Image (T2I), Image-to-Image (I2I), and Text and Image-to-Image (TI2I) prompts.
+Our framework evaluates the discrepancy across five dimensions: naive image quality, semantic alignment, aesthetic appeal, downstream applicability, and human validation.
+Results reveal notable gaps, emphasizing the importance of aligning metrics with human judgment.
+Source code and datasets are available at https://shorturl.at/l83W2.
+Artificial Intelligence Generated Content, AI-synthesized Images, Multimodal, Discrepancy, Evaluation Framework
+1 Introduction
+With the rapid advancement of deep learning techniques, the proliferation of Artificial Intelligence Generated Content (AIGC) has garnered significant attention across various domains, such as e-commerce, gaming, medicine, animation, and autonomous driving (Li et al., 2024; Qian et al., 2024).
+AI-generated Images (AIGI) have been one of the mainstream forms of AIGC, and a variety of AI image generative models have been proposed to make the generated synthetic images as realistic as natural images, ranging from the earlier generative adversarial networks (GANs) (Tao et al., 2023, 2022), advanced diffusion models (DMs) (Xu et al., 2023b; Wei et al., 2023) to the large multimodal generative model like DALL·E (Ramesh et al., 2022).
+Figure 1: Toy example demonstrating the key aspects of D-Judge.
+Despite the rapid development of AI image generative models, AI-generated images still fall short of real-world application standards due to discrepancies with realistic natural images (Li et al., 2023a).
+To address this, various AI-generated image quality assessment methods have emerged (Wang et al., 2023; Xu et al., 2023a; Hu et al., 2023).
+For instance, Wang et al.
+(Wang et al., 2023) introduced the AIGCIQA2023 database and a subjective evaluation framework based on quality, authenticity, and text-image correspondence.
+Similarly, PKU-I2IQA (Yuan et al., 2023) developed a perception-based database for image-to-image generation using no-reference and full-reference methods.
+Other studies, such as Pick-a-Pic (Kirstain et al., 2023), HPS v2 (Wu et al., 2023), TIFA (Hu et al., 2023), and ImageReward (Xu et al., 2023a), focused on training unified score models for automatic AIGC image quality assessment, moving beyond traditional metrics like Inception Score (Salimans et al., 2016) and FID (Heusel et al., 2017).
+Despite advancements in AIGC research, several critical issues remain unresolved.
+First, there is a lack of systematic and comprehensive studies addressing the discrepancies between AI-generated images and realistic natural images, which is crucial for enabling the practical application of AIGC images in real-world scenarios and achieving meaningful breakthroughs in the field.
+Most prior studies primarily focus on perceptual quality using traditional image quality metrics, text-to-image correspondence, or subjective human preferences, leaving a gap in understanding the broader implications of these discrepancies.
+Second, previous research predominantly considers unimodal prompts, such as T2I or I2I, while neglecting multimodal prompts like TI2I, and the datasets used are often small, typically comprising only a few thousand images, which limits their capacity for comprehensive analysis and evaluation.
+To address the above challenges, we proposed an AI-Natural Image Discrepancy Evaluation Benchmark (D-Judge) to investigate and interpret what discrepancies still remained between AI-generated images and natural images, finally answering the important question “How far are AI generative models with respect to the visual forms of images?”.
 Especially, we have two important contributions with reference to the key problems.
-Large Multimodal Evaluation Dataset construction: we have curated a large multimodal evaluation dataset called DNAI (Distinguishing Natural and AI-generated) dataset, with about 440,000 AIGI generated from 8 representative generative models based on both unimodal and mutimodal guidance including Text-to-Image (T2I), Image-to-Image (I2I), and Text vs.
-Image-to-Image (TI2I).
+Large Multimodal Evaluation Dataset construction: we have curated a large multimodal evaluation dataset called D-ANI (Distinguishing Natural and AI-generated) dataset, with about 440,000 AIGI generated from 9 representative generative models based on both unimodal and mutimodal guidance including Text-to-Image (T2I), Image-to-Image (I2I), and Text and Image-to-Image (TI2I).
 The data size of our DNAI dataset scales to 100x that of prior datasets.
 Fine-grained Evaluation Framework, we propose a fine-grained evaluation framework to conduct the systematic and comprehensive assessment and evaluation of DNAI, covering 5 diverse and important aspects, including naive visual feature quality, semantic alignment among multimodal generation, aesthetic appeal, downstream applicability, and the coordinated human validation.
-Through our DNAI dataset and fine-grained evaluation framework, we conduct extensive benchmark analysis and evaluation and conclude key insights to answer the discrepancy questions:
-  • Significant discrepancies in key Areas: AI-generated images exhibit substantial discrepancies from natural images in terms of quantitative measures from all aspects, with about 10% to 30%.
-  • Multimodal Alignment: Different prompted generation might have different semantic alignment scores.
-Generated images prompted with texts including both T2I and TI2I demonstrated remarkable semantic alignments.
-  • Downstream Task Applicability: Significant differences exist in the usability of AI-generated images vs.
-natural ones in downstream tasks, highlighting the need for further improvements of generative models to ensure practical applicability in real-world scenarios.
+The toy example is illustrated in Figure 1.
+For example, given a text prompt, an image is generated using a generative model and paired with a natural reference image retrieved from the natural database.
+D-Judge then evaluates the discrepancy rate between the two images across the aforementioned dimensions.
+Through our D-ANI dataset and fine-grained evaluation framework D-Judge, we conduct extensive benchmark analysis and evaluation and conclude key insights to answer the discrepancy questions:
+  • Significant discrepancies in key Areas: AIGIs exhibit substantial discrepancies from natural images in terms of quantitative measures from all aspects, up to 40%, 33.57%, 27.66%, and 96.95% in frame-level image quality, semantic alignment, aesthetic appeal, and downstream applicability, respectively.
+  • Multimodal Alignment: Different prompted generations might have different semantic alignment discrepancies compared to natural ones.
+Generated images prompted with texts including both T2I and TI2I demonstrated less semantic discrepancy than I2I.
+  • Downstream Task Applicability: Significant differences in usability between AI-generated and natural images are observed in downstream tasks, particularly in fine-grained object recognition and VQA, with discrepancy rates reaching as high as 94.29% and 96.95%, respectively.
+This underscores the need for further advancements in generative models to enhance their practical applicability in real-world scenarios.
   • Human Evaluation vs.
-Quantitative Metrics: Human evaluation results reveal larger discrepancies compared with quantitative metrics.
+Quantitative Metrics: Human evaluation results reveal larger discrepancies compared with quantitative ones.
 It validates the necessity of incorporating human evaluation for coordination.
-Related Work
+2 Related Work
 Natural Image Evaluation:
-Over the past decades, numerous image quality assessment methods have been developed to evaluate traditional natural images (N.
-et al.
-2015).
-These methods have focused on various visual features and properties, such as perceptual appearance (Zhang et al.
-2018), naturalness (Ma et al.
-2018), and aesthetics (Esfandarani and Milanfar 2018).
+Over the past decades, numerous image quality assessment methods have been developed to evaluate natural images (N.
+et al., 2015).
+These methods have focused on various visual features and properties, such as perceptual appearance (Zhang et al., 2018), naturalness (Ma et al., 2018), and aesthetics (Esfandarani & Milanfar, 2018).
 For instance, BRISQUE evaluates natural image quality by analyzing spatial domain features, while PIQE assesses perceptual quality through block-based image segmentation, both serving as effective no-reference metrics.
-Other notable measures include FID and Inception Scores, which assess the quality and diversity of generated images, and SSIM (Wang et al.
-2004) and PSNR, which quantify structural similarity and pixel-level accuracy.
+Other notable measures include FID and Inception Scores, which assess the quality and diversity of generated images, and SSIM (Wang et al., 2004) and PSNR, which quantify structural similarity and pixel-level accuracy.
 Beyond quantitative measures, human evaluations have also been leveraged to estimate natural image quality.
-For example, NIMA (Esfandarani and Milanfar 2018), proposed by Talebi and Milanfar, uses a deep CNN trained on human-rated images to predict aesthetic quality.
+For example, NIMA (Esfandarani & Milanfar, 2018), proposed by Talebi and Milanfar, uses a deep CNN trained on human-rated images to predict aesthetic quality.
 Similarly, Wong et al.
-introduced the AVA dataset, which facilitates aesthetic assessment based on human preferences (Murray, Marchesotti, and Perronnin 2012).
+introduced the AVA dataset, which facilitates aesthetic assessment based on human preferences (Murray et al., 2012).
 AI-generated Image Evaluation:
 Recently, AI-generated content (AIGC) has seen significant advancements with the rise of generative models.
-Several methods and benchmarks have been introduced to evaluate AI-generated images (Zhang et al.
-2023; Hu et al.
-2023).
-These methods primarily focus on three aspects: perceptual quality, text-image correspondence, and aesthetics (Xu et al.
-2023a; Wang et al.
-2023), often training unified models to assess the overall quality of AI-generated images.
-For example, ImageReward (Xu et al.
-2023a) presents a framework for aligning image generation with human preferences, AGIQA-3k (Li et al.
-2023a) offers a comprehensive benchmark for assessing AI-generated image quality, and QBench (Wu et al.
-2024) evaluates content quality across multiple dimensions.
+Several methods and benchmarks have been introduced to evaluate AI-generated images (Zhang et al., 2023; Hu et al., 2023).
+These methods primarily focus on three aspects: perceptual quality, text-image correspondence, and aesthetics (Xu et al., 2023a; Wang et al., 2023), often training unified models to assess the overall quality of AI-generated images.
+For example, ImageReward (Xu et al., 2023a) presents a framework for aligning image generation with human preferences, AGIQA-3k (Li et al., 2023a) offers a comprehensive benchmark for assessing AI-generated image quality, and QBench (Wu et al., 2024a) evaluates content quality across multiple dimensions.
 Despite advancements, little research has delved into the fine-grained differences between AI-generated and natural images, and existing datasets are often too small for comprehensive evaluation.
-To address these gaps, we propose an AI-Natural Image Difference Evaluation Benchmark, featuring a large Distinguishing AI-Natural Image Dataset and a systematic, fine-grained evaluation framework to thoroughly explore these differences.
-Figure 1: Overview of AI-Natural Image Discrepancy Evaluation Benchmark (ANID): (a) A large Distinguishing AI-Natural Image (DANI) dataset is curated, guided by three types of prompts.
-(b) A fine-grained evaluation framework is proposed to measure and evaluate the potential discrepancies between AI-generated images and natural images from five aspects covering naive image quality, semantic alignment, aesthetic appeal, downstream applicability, and coordinated human assessment.
-AI-Natural Image Discrepancy Evaluation Benchmark
-To tackle the challenges of current AI-generated image evaluation, we construct an AI-Natural Image Discrepancy evaluation benchmark (ANID) to evaluate the potential discrepancies between AI-generated images and natural images.
-Figure 1 shows the overview of our ANID benchmark.
-Our ANID benchmark contributes to two core components: (1) A Distinguishing AI-Natural Image (DANI) Dataset is constructed, in which we collect AI-generated images from diverse generative models for real natural images in MS COCO dataset (Lin et al.
-2014) based on three types of guidance prompts e.g.
-text-only, image-only and text-image prompts; (2) We devise a systematic and comprehensive evaluation framework to measure and evaluate the potential differences between AI-generated images and real natural images from five aspects including naive image quality, semantic alignment, aesthetic appeal, downstream applicability, and the coordinated human assessment.
-In the following parts, we introduce the DANI Dataset and the evaluation framework in detail.
-Distinguishing AI-Natural Image Dataset
+To address these gaps, we propose D-Judge as AI-Natural image difference Assessment Benchmark, featuring a large Distinguishing AI-Natural Image Dataset and a systematic, fine-grained evaluation framework to thoroughly explore these differences.
+3 D-Judge
+Figure 2: The D-ANI dataset of D-Judge, guided by three types of prompts. Each generated image is paired with both the corresponding text and image from the naturally collected dataset (e.g., COCO).
+Figure 3: The framework of D-Judge evaluates the Discrepancy Rate ($DR$) between AI-generated images and natural reference images across five dimensions: ➊ Naive Quality, ➋ Alignment, ➌ Aesthetic, and ➍ Applicability, complemented by ➎ human validation.
+The images with green borders represent reference natural images, while those with yellow borders are AI-generated counterparts.
+This comprehensive evaluation integrates pixel-level, frame-level, and distribution-level quality checks, alongside advanced tools like OpenAI CLIP, OpenAI Laion Aesthetic Evaluator, Qalign, ViT, U2Net, and BLIP2 to robustly assess the gap between the two image types.
+Where $CS$ means CLIP Score for the text and given images, and $AES$ means Aesthetic score.
+To tackle the challenges of current AI-generated image evaluation, we construct an AI-Natural Image Discrepancy evaluation benchmark (D-Judge) to assess the potential discrepancies between AI-generated images and natural images and answer how far AIGIs are from Natural Images.
+Firstly, a Distinguishing AI-Natural Image (D-ANI) Dataset is constructed, in which we collect AI-generated images from diverse generative models for real natural images in MS COCO dataset (Lin et al., 2014) based on three types of guidance prompts, e.g., text-only, image-only and text-image prompts, as shown in Figure 2; Then, we devise a systematic and comprehensive evaluation framework as Figure 3 illustrated to measure and evaluate the potential differences between AI-generated images and real natural images from five aspects including naive image quality, semantic alignment, aesthetic appeal, downstream applicability, and the coordinated human assessment.
+In the following parts, we introduce the D-ANI Dataset and the evaluation framework in detail.
+3.1 Distinguishing AI-Natural Image Dataset
 We construct the Distinguished Natural and AI-generated Image Dataset based on the classical natural image dataset MS COCO dataset.
@@
 The 25,000 text-image pairs are seen as the referenced natural image set.
-For each text-image sample in the referenced natural image set, we collect AI-generated images from 8 representative generative models, guided by three types of prompts, including text-only prompts, image-only prompts, and text-image prompts.
-The 8 representative generative models are from diverse generative models ranging from the earlier generative adversarial networks (GANs) (e.g.
-GALIP (Tao et al.
-2023), DF-GAN (Tao et al.
-2022)) to the recent benchmarking diffusion models (DMs) like Stable Diffusion v1.4, v1.5, v2.1, XL (Rombach et al.
-2022), Versatile-Diffusion (Xu et al.
-2023b) and OpenAI DALL·E 2 (Ramesh et al.
-2022).
+For each text-image sample in the referenced natural image set, we collect AI-generated images from 9 representative generative models, guided by three types of prompts, including text-only prompts, image-only prompts, and text-image prompts.
+The 9 representative generative models are from diverse generative models ranging from the earlier generative adversarial networks (GANs) (e.g.
+GALIP (Tao et al., 2023), DF-GAN (Tao et al., 2022)) to the recent diffusion models (DMs) like Stable Diffusion v1.4 (SD14), v1.5 (SD15), v2.1 (SD21), XL (SDxl) (Rombach et al., 2022), Versatile-Diffusion (VD) (Xu et al., 2023b), OpenAI DALL·E 2 (D·E 2) (Ramesh et al., 2022) and DALL·E 3 (D·E 3) (OpenAI, 2024).
 The referenced natural images and the generated images comprise our distinguishing AI-Natural image dataset.
-More details about data construction and collection are presented in Section Appendix A.
-Table 1 shows the statistics and properties of our DANI Dataset compared with the existing evaluation dataset.
-Our DANI Dataset has a total of 440,000 AI-generated images from 8 representative generated models.
+More details about data construction and collection are presented in Appendix Sec.
+B.
+Table 1 shows the statistics and properties of our D-ANI Dataset compared with the existing evaluation dataset.
+Our D-ANI Dataset has a total of 440,000 AI-generated images from 9 representative generated models.
 For details, please refer to Appendix Sec.
-A.
-Compared with state-of-the-art evaluation datasets, our DANI Dataset has convincing contributions:
-  • DANI Dataset is the most extensive dataset for AI-generated image evaluation, and the data size scales to 100x.
-  • DANI Dataset is guided by both unimodal prompts and multi-modal prompts (text-image). Prior datasets barely considered text prompts.
-  • Diverse generative models are exploited, ranging from the earlier GAN models, the recent popular diffusion models, and the influential large commercial generative model DALL·E 2 issued by OpenAI.
-Our large and comprehensive dataset works as the foundation of the evaluation benchmark and enables thorough evaluation and analysis of the differences between AI-generated images and realistic natural images.
+B.
+Compared with state-of-the-art evaluation datasets, our D-ANI Dataset has convincing contributions:
+  • D-ANI Dataset is the most extensive dataset for AI-generated image evaluation, and the data size scales to 100x.
+  • D-ANI Dataset is guided by both unimodal prompts and multi-modal prompts (text-image). Prior datasets barely considered text prompts.
+  • Comprehensive and Diverse generative models are exploited, ranging from the earlier GAN models, the recent popular diffusion models, and the influential large commercial generative model DALL·E 2 and DALL·E 3 issued by OpenAI.
+Our large and comprehensive dataset works as the foundation of the assessment benchmark and enables thorough evaluation and analysis of the differences between AI-generated images and realistic natural images.
+Table 1: Comparison of D-ANI dataset and Counterparts.
 Dataset Model Text Image Text vs. Image AIGI
@@
 PKU-I2IQA 2 200 200 - 1,600
-DNAI (Ours) 8 25,000 5,000 25,000 440,000
-Table 1: Comparison of DANI dataset and Counterparts.
-Fine-Grained Evaluation Framework
-On top of the DANI dataset, we construct a fine-grained evaluation framework to systematically measure and interpret the differences between AI-generated images and realistic natural images from five different aspects, including naive image quality, semantic alignment, aesthetic appeal, downstream applicability, and coordinated human assessment.
-1. Naive Image Quality:
+D-ANI (Ours) 9 25,000 5,000 25,000 445,000
+3.2 D-Judge Framework
+On top of the D-ANI dataset, we construct the fine-grained D-Judge to systematically measure and interpret the differences between AI-generated images and realistic natural images from five key different aspects as following aspects.
+1. Naive Image Quality Discrepancy:
 We leverage traditional image quality assessment methods for natural images to measure the AI-Natural image differences.
-In order to perform fine-grained analysis and exploration, we estimate the AI-Natural image differences from three levels by considering the inherent properties of image-style content, ranging from the low-level visual features, frame-level visual features to the holistic content distribution.
-  • First, we use the pixel-level similarity measures like SSIM, LPIPS, and DISTS to compute the visual similarities between the paired AI-generated image and natural image and use the inverse similarity values to quantify the AI-Natural image differences from the low-level visual view.
-  • Second, we use frame-level visual features to qualify the perceptional quality of images and compare the global-level visual quality differences between AI-generated images and natural images from the high-level visual view.
+In order to perform fine-grained analysis and exploration, we estimate the AI-Natural image differences from three levels by considering the inherent properties of image-style content, ranging from the pixel-level visual features, frame-level visual features to the holistic content distribution.
+  • First, we use the full-referenced pixel-level similarity measures like SSIM, LPIPS, and DISTS to compute the visual similarities between the paired AI-generated image and natural image and use these calculated values to quantify the AI-Natural image differences from the low-level visual view.
+  • Second, we use non-referenced frame-level visual features to qualify the perceptional quality of images and compute the frame-level visual quality difference rates between AI-generated images and natural images from the high-level visual view.
 The exploited frame-level visual features include PIQE (N.
-et al.
-2015), IL-NIQE (Zhang, Zhang, and Bovik 2015), MUSIQ (Ke et al.
-2021), DBCNN (Zhang et al.
-2020), LIQE (Ma et al.
-2018), Inception Score (Salimans et al.
-2016), CLIPIQA (Radford et al.
-2021a), TReS (Gu et al.
-2015), HyperIQA (Su et al.
-2020), UNIQUE (Zhang et al.
-2021), BRISQUE , NIQE (Mittal, Soundararajan, and Bovik 2013), NRQM (Ma et al.
-2018), etc.
-These metrics provide insights into global-level visual quality differences.
-  • Third, we investigate the visual content distribution to capture the holistic differences between AI-generated images and natural images by computing the FID and Inception Score of the whole AI-generated image dataset and the referenced natural image set.
-Based on the different levels of image quality measures, our framework can evaluate the fine-grained AI-Natural image differences, covering both low-level and high-level.
+et al., 2015), IL-NIQE (Zhang et al., 2015), MUSIQ (Ke et al., 2021), DBCNN (Zhang et al., 2020), LIQE (Ma et al., 2018), Inception Score (Salimans et al., 2016), CLIPIQA (Radford et al., 2021b), TReS (Gu et al., 2015), HyperIQA (Su et al., 2020), UNIQUE (Zhang et al., 2021), BRISQUE , NIQE (Mittal et al., 2013), NRQM (Ma et al., 2018), etc.
+The Discrepancy Rate ($DR$) of these metrics provides insights into frame-level visual quality differences.
+  • Third, we investigate the full-reference visual content distribution to capture the holistic differences between AI-generated images and natural images by computing the FID, CLIP-FID, and Inception Score of the AI-generated image set and the referenced natural image set.
+The results of these set-level measures can reflect the difference between AIGIs and natural ones from an integral perspective.
+Based on the different levels of image quality measures, our framework can assess and measure the fine-grained AI-Natural images differences with Full-reference and Non-reference angles, covering both low-level and high-level.
 More details about the utilized traditional image quality assessment measures are described in Appendix Sec.
 C.
-2. Semantic Alignment:
+2. Semantic Alignment Discrepancy:
 Generative models require the referenced prompts as semantic guidance to generate images.
 Therefore, semantic alignment can be an important quality indicator of the AIGIs.
-We use the widely used CLIP model (Radford et al.
-2021b) to measure semantic alignment and compare the different CLIP Scores from natural text, image, and the paired AI-generated image-text prompt.
-The studied AI-generated images were originally generated from the 8 Representative generative models based on three types of prompts as described in Appendix Sec.
-B; the semantic alignment evaluation can provide a fine-grained analysis of the potential differences in semantic alignment with respect to both unimodal and multimodal guidance.
-3. Aesthetic Appeal:
-Aesthetic Appeal evaluation is to estimate the visual appeal and artistic quality of images, reflecting the visual attractiveness and artistic quality of images.
-We utilize the classical aesthetic measures NIMA (Esfandarani and Milanfar 2018) and LAION-AES (Schuhmann et al.
-2022) as the quantitative metrics and compare the aesthetic scores for the paired AI-generated and natural images.
-4. Downstream Applicability:
+We use the widely used CLIP model (Radford et al., 2021a) to calculate the difference rates of Clip Scores, $DR_{CS}$ to represent alignment discrepancies between AIGIs and natural images, and provide a fine-grained analysis with respect to both unimodal and multimodal guidance.
+3. Aesthetic Appeal Discrepancy:
+Aesthetic Appeal is to estimate the visual appeal and artistic quality of images, reflecting the visual attractiveness and artistic quality of images.
+We utilize the classical aesthetic measures NIMA (Esfandarani & Milanfar, 2018), LAION-AES (Schuhmann et al., 2022a), and QAlign (Wu et al., 2024b) as the quantitative metrics and compare the difference rates of aesthetic scores $DR_{aes}$ for the paired AI-generated and natural images.
+4. Downstream Applicability Discrepancy:
 This aspect is devised to investigate the practical utility of AI-generated images in downstream tasks and evaluate whether AI-generated images can have different practical utilities with realistic natural images in downstream application tasks.
-We mainly focus on two classical downstream tasks, e.g., image recognition and object segmentation.
-For the image recognition task, we evaluate the Classification Mismatch Rate (MR) rates between AI-generated images and natural images when using a pre-trained image recognizer like ResNet-152 (He et al.
-2016) model.
-For the object segmentation task, we use the Intersection over Union (IoU) metric (Everingham et al.
-2010) as quantitative measures and investigate the segmentation results for AI-generated images with the segmentation model U2NET (Qin et al.
-2020).
-5. Human Assessment:
+We mainly focus on three common downstream tasks, e.g., image recognition, object segmentation and Multimodal Visual Question Answering (VQA).
+For the image recognition task, we evaluate the Difference recognition rates ($DR_{re}$) between AI-generated images and natural images when using a pre-trained image recognizer like ResNet-152 (He et al., 2016) model.
+For the object segmentation, we use the Intersection over Union (IoU) (Everingham et al., 2010) as quantitative measures and investigate the segmentation discrepancy for AI-generated images with U2Net (Qin et al., 2020).
+For VQA, we design three types of questions: ‘object’ (identifying objects in the image), ‘alignment’ (determining if the image aligns with a textual description, returning ‘yes’ or ‘no’), and ‘similarity probability’ (calculating the alignment probability between text and image).
+Using BLIP2 (Li et al., 2023b), we evaluate the Difference Rate ($DR$) between AI-generated images and reference natural images for each question type.
+5. Human Validation:
 We involve human assessments to coordinate the above evaluation aspects.
@@
 Detailed information about human assessment procedures is provided in Appendix Sec.
-D.
-Through the integration of diverse and comprehensive evaluation dimensions, our fine-grained evaluation framework offers a systematic assessment solution to investigate and interpret the differences that still remain between AI-generated images and natural images.
-Figure 2: We evaluate eight image-generating models from 6 aspects. The numerical values in the radar chart represent the mean difference rate of each model.
-Evaluation
+D.3.
+Through the integration of diverse and comprehensive evaluated dimensions, our D-Judge offers a systematic assessment solution to investigate and interpret the differences that still remain between AI-generated images and natural images.
+Figure 4: We evaluate nine image-generating models from 6 aspects. The numerical values in the radar chart represent the mean difference rate of each model.
+4 Evaluation
 With respect to the 5 different aspects, we conducted benchmark experiments and performed experimental evaluation alongside the following research questions.
-RQ1: What are the fine-grained discrepancies between AIGIs and natural images?
+RQ1: What are the fine-grained naive image quality discrepancies between AIGIs and natural images?
   • RQ1.a: What discrepancies between I2I-guided AIGIs and their natural counterparts can be interpreted via the pixel-level quality measures?
@@
   • RQ1.c: How do the structural visual content distributions differ between AIGIs and natural images?
-RQ2: How significant are the discrepancies in semantic alignment between AIGIs and natural images in different types of guided prompts?
+RQ2: How significant are the discrepancies in the semantic alignment of the given text with natural images and AIGIs in different types of guidance?
 RQ3: What are the discrepancies in aesthetic appeal between AIGIs and natural images?
@@
 RQ5: Are human assessment results consistent with quantitative measures? What are the discrepancies revealed from human evaluation?
-Experimental Setting
-We use the described quantitative evaluation metrics, and details of all metrics are presented in Appendix Sec. B.
+4.1 Experimental Setting
+We use the described quantitative evaluation metrics, which can be divided into Full-reference and No-reference, to present the $DR$ for involved assessment dimensions (more details of all metrics are presented in Appendix Sec.
+C.
 Full-reference metrics:
-For full-reference metrics like pixel-level image quality, we directly report the calculated value cause it already can present the discrepancy.
+For full-reference metrics, which are used for pixel-level image quality, visual content distribution, and downstream object segmentation task, i.e., IoU, we directly report the calculated value to reflect the difference cause it already can present the discrepancy.
 No-reference metrics
 We calculate Difference Rate (DR) for each quantitative metric in non-reference scenarios, by subtracting the value of each AI-generated image from its corresponding natural reference image and then averaging these difference rates, defined as follows:
-Difference $\text{{Difference Rate}}=\frac{1}{n}\sum_{i=1}^{n}\frac{\left|m(\mathbf{X}_{i})-m(\mathbf{N}_{i})\right|}{m(\mathbf{N}_{i})}$ (1)
+$\text{{DR}}=\frac{1}{n}\sum_{i=1}^{n}\frac{\lvert((M(\mathbf{X}_{i})-m_{\text{min}})-(M(\mathbf{N}_{i})-m_{\text{min}}))\rvert}{m_{\text{max}}-m_{\text{min}}}$ (1)
 where $n$ is the number of images, $M$ represents the metric, $\mathbf{X}$ denotes the generated images, and $\mathbf{N}$ refers to the referenced natural images.
-RQ1: Naive Quality Results
-To evaluate the image quality of AI-generated images versus natural images, we conducted a comprehensive analysis using a variety of metrics that assess visual feature quality, naturalness, and similarity of images.
+Additionally, $m_{\text{min}}$ and $m_{\text{max}}$ represent the minimum and maximum values across all images, respectively, especially when the value range of the corresponding metric is undefined.
+4.2 RQ1: Naive Quality Results
+To evaluate the image quality of AI-generated images versus natural images, we conducted a comprehensive analysis using a variety of metrics that assess the discrepancies between AIGIs and natural images across pixel-level similarity, frame-level quality and visual content distribution.
 Our findings reveal substantial discrepancies between AIGIs and natural images across these dimensions.
-RQ1.a Pixel-level
-We evaluated the visual similarity between AI-generated and natural images using structural metrics like SSIM, LPIPS, DISTS, and PSNR.
-The results, as shown in Table 2, indicate that AI-generated images exhibit significantly lower similarity to natural counterparts.
-AI-generated images show a 20% to 50% reduction in SSIM, highlighting a major loss in structural fidelity.
-LPIPS and DISTS further reveal notable perceptual dissimilarities, with LPIPS scores ranging from 0.19 to 0.60.
-Additionally, PSNR values, between 12.42 and 23.25, suggest that AI-generated images have higher noise levels, reducing their overall visual similarity.
-SD_V14 SD_V15 Versatile SD_V21 SD_XL DALL·E 2
-SSIM ↑ 0.35 0.35 0.46 0.44 0.70 0.38
-PSNR ↑ 13.81 13.80 16.52 16.00 23.25 12.42
-VIF ↑ 0.02 0.02 0.05 0.04 0.17 0.03
+4.2.1 RQ1.a Pixel-level
+We evaluated the visual similarity between AI-generated and natural images using structural metrics like SSIM, LPIPS, DISTS, PSNR, etc.
+The mean similarity values of these metrics, as shown in Table 2, indicate that AI-generated images exhibit significantly lower similarity to natural counterparts.
+AI-generated images show a 30% to 65% reduction in SSIM, highlighting a major loss in structural fidelity.
+LPIPS and DISTS further reveal notable perceptual dissimilarities, with LPIPS scores ranging from 0.19 to 0.61.
+Additionally, PSNR values between 12.29 and 23.11 suggest that AI-generated images have higher noise levels, reducing their overall visual similarity.
+Table 2: Full-reference Evaluation Results (I2I).
+SD14 SD15 VD SD21 SDxl D·E 2
+SSIM ↑ 0.35 0.35 0.46 0.43 0.70 0.37
+PSNR ↑ 13.72 13.71 16.41 15.86 23.11 12.29
+VIF ↑ 0.02 0.02 0.05 0.04 0.18 0.03
 VSI ↑ 0.84 0.84 0.89 0.87 0.96 0.83
-FSIM ↑ 0.61 0.61 0.71 0.67 0.86 0.61
+FSIM ↑ 0.61 0.61 0.71 0.66 0.86 0.60
 LPIPS ↓ 0.60 0.60 0.39 0.52 0.19 0.61
 DISTS ↓ 0.27 0.27 0.17 0.23 0.10 0.24
-MAD ↓ 211.61 211.73 196.86 202.29 143.66 214.86
-Table 2: Full-reference Evaluation Results (I2I).
-RQ1.b Frame-level
-We assessed the frame-level quality of AI-generated and natural images using a suite of metrics, including MUSIQ, DBCNN, HyperIQA, LIQE, BRISQUE, NRQM and NIQE.
-Figure 2 uses radar figures to visualize the results from the six metrics, and indicates that there are significant quality differences between AI-generated and natural images.
-These deviations typically range from 10% to 20%, irrespective of the guidance method used, with some instances showing quality differences as high as 30%.
+MAD ↓ 212.91 213.09 197.92 203.89 144.92 217.12
+4.2.2 RQ1.b Frame-level
+We assessed the frame-level discrepancies of AI-generated and natural images using a suite of metrics, including DBCNN, HyperIQA, LIQE, PIQE, NRQM and QAlign.
+Figure 4 uses radar figures to visualize the mean difference rates from the six metrics, and indicates that there are significant quality differences between AI-generated and natural images at the frame level.
+These deviations are typically around 10%, irrespective of the guidance prompt used, with some instances showing quality differences as high as 40%.
 This underscores the considerable gaps between AIGIs and natural images in terms of global image quality.
-Moreover, the results of NRQM and NIQE reveal that AI-generated images significantly differ from natural images in terms of naturalness.
+Moreover, the results of NRQM reveal that AI-generated images significantly differ from natural images in terms of naturalness.
 Specifically, AI-generated images show notable deviations from natural images in visual realism and adherence to natural scene statistics.
-The NRQM difference rates range from 5% to 20%, while the NIQE difference rates range from 10% to 50%, highlighting the challenges AI-generated images face in replicating the inherent naturalness of real-world scenes.
+The NRQM difference rates range from 5% to 32%, with images generated by D·E 3 exhibiting particularly high rates, underscoring the challenges AI-generated images encounter in replicating the inherent naturalness of real-world scenes.
 More detailed results with more metrics are presented in Appendix Sec.
-C.
-Figure 3: The FID and Inception Score of AI-generated images from different source models and guidance.
-RQ1.c Visual Content Distribution
-We present the FID and Inception Score results, presented in Figure 3, suggesting that AI-generated images often exhibit relatively higher FID scores and lower Inception Scores.
-This finding indicates that current generative models, particularly GAN-based models and earlier versions of Stable Diffusion, struggle to consistently produce high-quality images.
+D.1 Figure 12.
+Figure 5: The FID (left, ↓) and Inception Score (right, ↑) of AI-generated images from different source models and guidance.
+4.2.3 RQ1.c Visual Content Distribution
+We present the FID and Inception Score results, presented in Figure 5, suggesting that AI-generated images often exhibit relatively higher FID scores (up to 50) and lower Inception Scores (low to 14).
+This finding indicates that current generative models, particularly GAN-based models and earlier versions of Stable Diffusion, struggle to produce high-quality images as naturally collected images consistently.
 Notably, these models tend to generate better images when a reference image is provided (TI2I), as it has more detailed guidance.
-In conclusion, our evaluation of image visual feature quality, naturalness, and similarity reveals significant disparities between AI-generated and natural images.
+In conclusion, our evaluation of image quality across pixel-level, frame-level, and content distribution-level reveals significant disparities between AI-generated and natural images.
 These differences highlight the current limitations of AIGC technologies in matching the visual quality and realism of natural images.
-RQ2: Alignment Results
-In our evaluation of alignment results, we assessed how well AI-generated images correspond to their respective reference prompts across different guidance types: Text-to-Image (T2I), Image-to-Image (I2I), and Text-and-Image-to-Image (TI2I).
-And we find that AI-generated images often struggle to maintain high alignment with their prompts, especially the image-only reference provided.
-The results, shown in Table 3, indicate that AI-generated images typically achieve lower CLIP Score compared to natural images, with alignment discrepancies ranging from 4.02% to 36.69%.
-We also find that even if the AIGIs can get a higher CLIP score, especially with text and image-guided, they are looking stranger to humans as Table 3 illustrated.
-The CLIP score, which measures the alignment between text descriptions and generated images by embedding both into a shared space and calculating cosine similarity, revealed a noticeable gap in how accurately AIGC images reflect the input prompts.
+4.3 RQ2: Alignment Results
+We calculate the difference rate of CLIP Scores ($DR_{CS}$) to assess how well the AIGIs align with natural ones in terms of semantic alignment across different guidance types: Text-to-Image (T2I), Image-to-Image (I2I), and Text-and-Image-to-Image (TI2I).
+And we find that AIGIs often struggle to maintain semantic alignment as natural ones, especially the image-only reference provided.
+The results, shown in Table 3, indicate that the $DR_{CS}$ range from 4.02% to 36.69%.
+We also find that even if the AIGIs can get a higher CS sometimes, especially with text and image-guided, they are looking stranger to humans.
 This suggests that current AIGC models often struggle to maintain high fidelity to the provided references, leading to variations that may impact the intended alignment.
 Our findings underscore the challenges that these models face in consistently generating content to accurately reflect the specified input conditions, highlighting both their strengths and limitations in real-world applications.
-GALIP DFGAN SD_V14 SD_V15 Versatile SD_V21 SD_XL DALL·E 2
-T2I 10.72 18.51 9.78 9.69 9.71 9.93 10.20 9.46
-I2I - - 36.46 36.69 4.92 26.96 3.98 5.41
-TI2I - - 9.21 9.20 8.03 8.48 4.02 -
-Table 3: The difference rate of CLIP score (%).
-RQ3: Aesthetic Appeal
+Table 3: The results of $DR_{CS}$ (%).
+GALIP DFGAN SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
+T2I 12.16 16.11 10.67 10.53 9.71 10.90 11.59 10.93 10.19
+I2I - - 33.43 33.57 4.89 24.04 3.56 5.53 -
+TI2I - - 9.90 9.73 8.22 9.20 3.57 - -
+4.4 RQ3: Aesthetic Appeal
 We assessed the aesthetic appeal of AI-generated and natural images using NIMA and LAION-AES metrics and found that AI-generated images generally fall short of natural images in terms of aesthetic quality.
-Figure 4 presents the results, which reveal that AI-generated images receive lower scores for aesthetic appeal, with discrepancies ranging from 3.34% to 23.85%.
-These findings indicate that while AI models can produce visually appealing images, they often lack the nuanced artistic quality and emotional impact that are characteristic of natural images.
-Figure 4: The aesthetic difference rate between AI-generated and naturally generated images.
-RQ4: Applicability Results
-When evaluating the applicability of AI-generated images, we focused on two crucial downstream tasks: image recognition and semantic segmentation.
+Figure 6 presents the aesthetic difference rate ($DR_{aes}$) results, which reveal that AI-generated images have noteworthy discrepancies compared to natural ones, with $DR_{aes}$ ranging from 3.51% to 27.66%.
+These findings indicate that while AI models can produce visually appealing images, AIGIs’ often lack the nuanced artistic quality and emotional impact that are characteristic of natural images.
+Figure 6: The $DR_{aes}$ between AIGIs and natural ones.
+4.5 RQ4: Applicability Results
+When evaluating the applicability of AI-generated images, we focused on three crucial downstream tasks: image recognition, semantic segmentation and VQA.
 Our findings indicate significant discrepancies between the performance of AIGC images and natural images in these tasks.
 Image Recognition:
-For image recognition, we measured the classification mismatch rate, which quantifies the differences in predicted labels between AIGC images and their natural counterparts.
-The results in Table 4 showed a substantial gap, with the mismatch rate ranging from 29.89% to 94.44%.
-This indicates that AI-generated images often fail to achieve the same level of accuracy as natural images, highlighting the limitations of current AIGC models in producing reliable content for image recognition tasks.
-The high mismatch rate underscores the need for further advancements in generative models to improve their applicability in real-world recognition tasks.
-Overall, our results indicate that AI-generated images currently exhibit significant limitations in applicability for both image recognition and semantic segmentation tasks.
-The high classification mismatch rates and variable mean IoU values highlight the gap between AIGC and natural images in practical applications.
-These findings emphasize the importance of ongoing research and development to enhance the reliability and accuracy of AIGC models, ensuring they can be effectively used in real-world scenarios.
-GALIP DFGAN SD_V14 SD_V15 Versatile SD_V21 SD_XL DALL·E 2
-T2I 72.02 87.43 94.44 94.29 43.72 84.42 30.41 69.74
-I2I - - 68.62 68.33 70.73 68.01 67.73 53.42
-TI2I - - 62.84 62.35 62.28 58.56 29.89 -
-Table 4: The results of Classification Mismatch Rate (MR).
-GALIP DFGAN SD_V14 SD_V15 Versatile SD_V21 SD_XL DALL·E 2
-T2I 0.23 0.23 0.25 0.25 0.26 0.25 0.26 0.26
-I2I - - 0.38 0.38 0.69 0.50 0.82 0.42
-TI2I - - 0.43 0.44 0.30 0.53 0.82 -
-Table 5: The results of Segmentation IoU.
+We measured the classification Difference Rate (DR) (%) to identify the differences in predicted labels between AIGC images and their natural counterparts.
+The results in Table 8 showed a substantial gap, with the DR ranging from 27.46% to 94.29%, indicating that the generated images can’t accurately present the object as expected.
 Semantic Segmentation:
-In semantic segmentation, we evaluated the performance using the mean Intersection over Union (IoU) metric and show the results in Table 5.
-The mean IoU values for AIGC images varied between 0.23 and 0.82, indicating a wide range of segmentation accuracy.
-While some generated images approached the performance of natural images, a significant portion of AIGIs fell short, demonstrating poor segmentation results.
-This variability suggests that while AIGC models can sometimes produce high-quality segmentation, they often struggle to consistently match the accuracy of natural images.
+We evaluated the performance using the Intersection over Union (IoU) metric and show the discrepancy results in Table 9.
+The $IoU$ values (higher is better) for AIGC images varied between 0.33 and 0.82, indicating a wide range of segmentation between AIGC images and their natural counterparts.
 Visual Question Answering:
-To further evaluate the performance of AI-generated images and natural images in downstream tasks, we conducted tests on the VQA model BLIP-2 (Li et al.
-2023b).
-The evaluation involved three different types of questions: (1) identifying objects in the image (Object), (2) determining whether the image aligns with a given text by answering ”yes” or ”no” (Align), and (3) predicting the similarity probability between the image and the text (Prob.).
-Table 6 summarizes the discrepancies (%) between natural and AI-generated images across various models.
-For Object questions, the highest discrepancies occurred in the I2I setting, with SD_V14 and SD_V15 reaching 77.27% and 75.51%, respectively, while T2I discrepancies ranged from 36.08% (DALL·E 2) to 68.96% (DFGAN).
-Align questions showed lower discrepancies, with T2I values ranging from 24.72% (Versatile) to 34.65% (DFGAN) and I2I values from 15.90% (Versatile) to 29.24% (SD_V15).
-For similarity probability (Prob.), T2I discrepancies varied between 36.61% (DALL·E) and 62.01% (DFGAN), while I2I results ranged from 31.03% (DALL·E) to 62.80% (SD_V14), highlighting model-dependent variations.
-These results indicate that AI-generated images, especially those produced under certain generative settings, face significant challenges in downstream VQA tasks compared to natural images.
-The findings also reveal performance variations across generative models, underscoring the need for improved alignment and semantic understanding in AI-generated images.
-GALIP DFGAN SD_V14 SD_V15 Versatile SD_V21 SD_XL DALL·E 2
-Object T2I 47.84 68.96 39.70 38.97 43.46 38.50 38.42 36.08
-I2I - - 77.27 75.51 28.78 71.59 28.37 26.03
-TI2I - - 41.18 38.06 37.64 39.27 26.69 -
-Align T2I 29.81 34.65 27.52 28.03 24.72 31.99 29.78 29.79
-I2I - - 28.48 29.24 15.90 27.07 16.33 25.50
-TI2I - - 27.52 25.79 19.26 27.31 15.85 -
-Prob. T2I 54.61 62.01 54.74 52.69 49.67 54.84 58.68 36.61
-I2I - - 62.80 58.39 36.16 56.14 42.88 31.03
-TI2I - - 52.75 48.71 43.69 46.30 43.24 -
-Table 6: The results of VQA task w.r.t Object, Alignment, and Similarity probability.
-SD_V21 SD_XL Versatile DALL·E 2
+Further, we apply these AI-generated image to the VQA task by BLIP-2 (Li et al., 2023b) involved three different types of questions: Object, Alignment (Align), and similarity probability (Prob.).
+Table 4 summarizes the discrepancies (%) between natural and AI-generated images across various questions.
+For Object, the highest discrepancies occurred in the I2I setting, with SD14 and SD15 reaching 96.41% and 96.95%, respectively, while T2I discrepancies ranged from 64.07% (DALL·E 3) to 88.43% (DFGAN).
+Align questions showed lower discrepancies, with T2I values ranging from 8.55% (DFGAN) to 21.18% (DALL·E 3) and I2I values from 7.87% (VD) to 10.07% (SD15).
+For similarity probability (Prob.), T2I discrepancies varied between 37.31% (DALL·E 2) and 62.69% (DFGAN), while I2I results ranged from 32.81% (DALL·E) to 61.56% (SD15).
+Overall, our findings highlight significant limitations of AI-generated images in tasks such as image recognition, semantic segmentation, and visual question answering (VQA).
+While AI-generated images perform comparably to natural images in simpler tasks, such as binary text-image alignment (e.g., responding with “yes” or “no”), they fall short in more complex scenarios, including accurately estimating text-image alignment probabilities, identifying specific objects, and segmenting object areas.
+These shortcomings limit their broader applicability in diverse downstream tasks compared to natural images.
+Our results underscore the need for continued research and development to improve the reliability and accuracy of AIGC models, enabling their effective deployment in real-world applications.
+Table 4: The Difference Rate (DR) of VQA tasks.
+GALIP DFGAN SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
+Object T2I 73.34 88.43 68.72 68.63 74.21 68.34 66.93 57.85 64.07
+I2I - - 96.41 96.95 50.43 91.14 51.65 46.29 -
+TI2I - - 67.22 66.65 66.29 67.22 49.20 - -
+Align T2I 10.57 8.55 14.02 15.79 12.32 13.02 13.00 12.99 21.18
+I2I - - 9.71 9.53 7.87 10.07 8.01 8.93 -
+TI2I - - 13.31 12.70 10.03 10.99 8.46 - -
+Prob. T2I 53.13 62.69 53.07 54.21 50.57 54.95 57.09 37.31 49.22
+I2I - - 61.05 61.56 39.57 58.67 43.64 32.81 -
+TI2I - - 50.96 50.34 45.93 49.97 41.48 - -
+4.6 Human Validation
+To rigorously assess the differences between AI-Natural images and to evaluate the effectiveness of existing metrics in measuring AIGC image quality, we conducted human evaluations focusing on three key aspects: Image Quality, Alignment, and Aesthetic Appeal.
+  • Image Quality: Participants assessed the overall quality of the images, considering factors such as clarity, detail, and the presence of artifacts.
+  • Alignment: This aspect measured how well the generated images matched the given prompts.
+  • Aesthetic: The aesthetic appeal of the images was rated based on visual attractiveness, composition, and overall artistic quality.
+We excluded results for Applicability since this dimension inherently involves human-generated ground truth labels.
+Participants rated images on a scale from 1 (poor) to 5 (excellent).
+Our human evaluation results are shown in Table 5 and Figure 7 reveal significant discrepancies between AI-generated images and natural images across all three assessed aspects: quality, alignment, and aesthetic appeal.
+AIGC images consistently received lower scores across all three dimensions; AI-generated images generally received lower scores, often between 2 and 3, compared to natural images.
+This emphasizes the need for further advancements to narrow these gaps.
+These findings underscore the importance of continued research and development in AI image generation to bring AIGC images closer to the standards of natural images.
+Moreover, the result in Figure 7 from human validation indicates that traditional quantitative metrics often fail to align with human preferences, suggesting that they are not fully compatible with assessing AIGC images.
+This highlights the necessity for developing new measures tailored specifically to the unique characteristics of AI-generated content.
+Table 5: Human score results across three aspects.
+SD21 SDxl VD D·E 2
 T2I I2I TI2I T2I I2I TI2I T2I I2I TI2I T2I I2I
@@
 Aesthetic 3.38 3.55 2.70 3.21 2.35 4.29 3.96 3.32 3.19 4.28 3.23
-Table 7: Human evaluation results.
 Figure 7: The difference rate provided by human & quantitative statistics, where ‘H’ represents human and ‘Q’ represents quantitative.
-Human Validation
-To rigorously assess the differences between AI-Natural images and to evaluate the effectiveness of existing metrics in measuring AIGC image quality, we conducted human evaluations focusing on three key aspects: Image Quality, Alignment, and Aesthetic Appeal.
-We excluded results for Applicability since this dimension inherently involves human-generated ground truth labels.
-Participants rated images on a scale from 1 (poor) to 5 (excellent).
-Across all three dimensions, AI-generated images generally received lower scores, often between 2 and 3, compared to natural images.
-Image Quality: Participants assessed the overall quality of the images, considering factors such as clarity, detail, and the presence of artifacts.
-The results revealed a noticeable gap between AIGC images and their natural counterparts, with the generated images often lacking the same level of detail and clarity.
-Alignment: This aspect measured how well the generated images matched the given prompts, revealing a notable discrepancy.
-Generated images frequently deviated from the descriptions, indicating issues with model adherence to specified conditions.
-Aesthetic: The aesthetic appeal of the images was rated based on visual attractiveness, composition, and overall artistic quality.
-While some generated images displayed impressive artistic qualities, the overall consistency and aesthetic appeal were generally considered inferior to that of natural images.
-Overall, our human evaluation results are shown in Table 7 and Figure 5 reveal significant discrepancies between AI-generated images and natural images across all three assessed aspects: quality, alignment, and aesthetic appeal.
-AIGC images consistently received lower scores, emphasizing the need for further advancements to narrow these gaps.
-These findings underscore the importance of continued research and development in AI image generation to bring AIGC images closer to the standards of natural images.
-Moreover, the results from human validation indicate that traditional quantitative metrics often fail to align with human preferences, suggesting that they are not fully compatible with assessing AIGC images.
-This highlights the necessity for developing new measures tailored specifically to the unique characteristics of AI-generated content.
-More Evaluations
-To analyze the differences between AI-generated and natural images across various categories, we evaluated the performance of generated data from eight categories (’person’, ’animal’, ’indoor’, ’outdoor’, ’vehicle’, ’food’, ’sports’, ’accessory’) across three key aspects: (1) Naive Quality, including pixel-level similarity (SSIM), frame-level quality (CLIPIQA), and content distribution (FID); (2) Alignment, measured by CLIP Score; and (3) Aesthetic Quality, assessed by LAION-AES.
-Partial results are presented in Figures 6 and 7, while the full experimental findings are provided in Appendix Sec.
-E (Figures 13 and 14 ).
-From Figure 6, we observe substantial variability in the differential rates for categories such as “food” and “sports”.
-Compared to “sports” images, the “food” demonstrate significantly higher discrepancies in image quality (CLIPIQA), revealing AI models struggle to generate natural food images.
-In Figure 7, the FID and SSIM metrics highlight distinct trends across categories.
-The “accessory” category exhibits the highest FID values, indicating poorer content distribution, while the “animal” category consistently achieves higher SSIM scores, reflecting better pixel-level similarity.
-These results suggest that certain categories, such as “accessory” and “indoor”, remain particularly challenging for AIGC models to reproduce accurately, especially in maintaining semantic consistency and achieving balanced content generation.
-Overall, these findings demonstrate that AI-generated images still fall short of matching natural images in terms of naturalness and image quality.
-Additionally, the pronounced performance discrepancies across categories underscore an imbalance in the training data used for AIGC models.
-Addressing these discrepancies will require more balanced datasets and targeted improvements in model architecture and training strategies.
-Figure 6: Differential rates of AI-generated images in the food and sports categories across naive image quality, alignment, and aesthetic metrics.
-Figure 7: FID and SSIM metrics for AI-generated images across different source models and categories.
-FID evaluates distribution-level image quality, while SSIM measures pixel-level similarity to assess overall image quality.
-Conclusion and Discussion
-In this paper, we develop a systematic and comprehensive framework to evaluate the discrepancies between AI-Natural images using the formulated DNAI dataset.
-Our findings reveal significant disparities in naive image quality, semantic alignment, aesthetics, and downstream task applicability.
+4.7 Discrepancies Across Categories
+To analyze the differences between AI-generated and natural images across various categories, we evaluated the discrepancy of AIGI from eight categories (‘person’, ‘animal’, ‘indoor’, ‘outdoor’, ‘vehicle’, ‘food’, ‘sports’, ‘accessory’) across three key aspects: ➊ Naive Quality, including pixel-level similarity (SSIM), frame-level quality (CLIPIQA), and content distribution (FID, CLIP-FID, and Inception Score); ➋ Alignment, measured by CLIP Score; and ➌ Aesthetic Quality, assessed by LAION-AES.
+The results in Figure 8 reveal substantial variability across different categories, even for the AIGI from the same model (please refer to Appendix Sec.
+D.2, Figures 16 and 17 for integrated details).
+This could be attributed to the significant imbalance in the distribution of training data across these categories (Schuhmann et al., 2022b; Aghajanyan et al., 2023).
+Figure 8: The difference rate on food and sport subcategories across quality, alignment, and aesthetic.
+5 Conclusion and Discussion
+In this paper, we develop a systematic and comprehensive D-Judge to evaluate the discrepancies between AI-Natural images using the constructed D-ANI dataset.
+Our findings reveal significant disparities in naive image quality, semantic alignment, aesthetics, and downstream task applicability, especially in the low-level and content distribution-level image quality and fine-grained downstream tasks, like object recognition and object VAQ.
 While AIGC models have achieved impressive innovations, AIGIs still lag behind natural images in these critical areas.
@@
 Addressing these limitations and improving models will help reduce the perceptual gaps between AIGC images and natural images, facilitating better integration into real-world applications.
-Appendix
-Appendix Overview
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Appendix A Appendix Overview
 The appendix provides supplementary details and additional experimental results that were not included in the main paper due to space limitations. It is organized as follows:
-  • Section A: In-depth Description of the DNAI Dataset.
-  • Section B: Detailed Overview of the Evaluated Generative Models.
+  • Section B: In-depth Description of the D-ANI Dataset.
   • Section C: Comprehensive Summary of Evaluation Metrics.
-  • Section D: Extended Information on Human Perceptual Evaluations.
+  • Section D: Additional Evaluation Results.
   • Section E: Discussion on the Safety Mechanisms in Existing Diffusion Models.
-A Details of the DNAI Dataset
-Overview
-The DNAI dataset is a carefully curated collection of both natural and AI-generated images, designed to support comprehensive evaluations of image generation models.
+Appendix B Details of the D-ANI Dataset
+B.1 Overview
+The D-ANI dataset is a carefully curated collection of both natural and AI-generated images, designed to support comprehensive evaluations of image generation models.
 The natural images are sourced from the COCO validation set.
@@
 Image-to-Image (TI2I).
-Representative samples from the DNAI dataset are shown in Figure 8.
-For a detailed breakdown of the DNAI dataset composition, please refer to Table 8.
-Figure 8: Examples of AI-generated content (AIGC) created using five state-of-the-art generative models: SD_v1.4, SD_v1.5, Versatile Diffusion, SD_v2.1, and SD_XL.
-Each case is illustrated with three rows corresponding to Text-to-Image (T2I), Image-to-Image (I2I), and Text-and-Image-to-Image (TI2I) generation modes.
-For the complete set of AIGC images, please refer to our DNAI dataset, which will be available publicly soon.
-Types DFGAN GALIP SD_V1.4 SD_V1.5 Versartile SD_V2.1 SD_XL DALL·E 2
-T2I 25,000 25,000 25,000 25,000 25,000 25,000 25,000 5,000
-I2I - - 25,000 25,000 25,000 25,000 25,000 5,000
-TI2I - - 25,000 25,000 25,000 25,000 25,000 -
-Table 8: The details of the DNAI dataset.
-Natural Images
-The DNAI dataset features a carefully curated selection of natural images sourced from the COCO validation set.
+Representative samples from the D-ANI dataset are shown in Figures 9,  10 and  11.
+For a detailed breakdown of the D-ANI dataset composition, please refer to Table 6.
+Figure 9: Examples of AI-generated content (AIGC) created using Text-to-Image (T2I) guidance.
+Figure 10: Examples of AI-generated content (AIGC) created using Image-to-Image (I2I) guidance.
+Figure 11: Examples of AI-generated content (AIGC) created using Text-and-Image-to-Image (TI2I) guidance.
+B.2 Natural Images
+The D-ANI dataset features a carefully curated selection of natural images sourced from the COCO validation set.
 This subset comprises 5,000 images, each paired with at least five unique captions, totaling 25,014 captions.
@@
 In this study, we use these images and their accompanying (five) captions to guide the generative models in creating new images.
-Generated Images
-The DNAI dataset includes AI-generated images created using several advanced generative models across three distinct guidance modes: Text-to-Image (T2I), Image-to-Image (I2I), and Text-and-Image-to-Image (TI2I).
+B.3 Generated Images
+The D-ANI dataset includes AI-generated images created using several advanced generative models across three distinct guidance modes: Text-to-Image (T2I), Image-to-Image (I2I), and Text-and-Image-to-Image (TI2I).
 These modes collectively contributed a significant volume of images to the dataset.
-The models utilized in this process include DF-GAN, GALIP, various versions of Stable Diffusion (v1.4, v1.5, v2.1, XL), Versatile Diffusion, and DALL·E 2.
-Each model, except DALL·E 2, generated 25,000 images per guidance mode.
-DALL·E 2 produced 5,000 images per mode, resulting in a total of 440,000 generated images.
-It is important to note that DF-GAN and GALIP were used only for T2I, while DALL·E 2 was employed for both T2I and I2I modes.
+The models utilized in this process include DF-GAN, GALIP, various versions of Stable Diffusion (v1.4, v1.5, v2.1, XL), Versatile Diffusion, DALL·E 2 and DALL·E 3.
+Each model, except DALL·E 2 and DALL·E 3, generated 25,000 images per guidance mode.
+DALL·E 2 and DALL·E 3 produced 5,000 images for per mode, resulting in a total of 440,000 generated images.
+It is important to note that DF-GAN, GALIP and DALL·E 3 were used only for T2I, while DALL·E 2 was employed for both T2I and I2I modes.
 For detailed information, please refer to Table 6.
-Image Resolution
-The resolution of the generated images varies depending on the model used.
-Images generated by GALIP are 224x224 pixels, while those produced by DF-GAN are 256x256 pixels.
-The other models, including various versions of Stable Diffusion and DALL·E 2, generate images at a resolution of 512x512 pixels.
-Dataset Composition
-The DNAI dataset offers a robust foundation for evaluating the performance and capabilities of various AIGC models under different conditions and guidance modes.
-By encompassing a broad spectrum of images and resolutions, the dataset ensures comprehensive coverage of the potential use cases and challenges associated with AI-generated content.
-The combination of natural and generated images, alongside diverse textual descriptions, facilitates a thorough assessment of image generation models, supporting the development of more accurate and reliable AI systems.
-B Evaluated Generative Models
-GALIP (Generative Adversarial Latent Image Processing): (Tao et al.
-2023) is a generative model that produces high-quality images by leveraging adversarial techniques.
+The involved generative models are as follows:
+  • GALIP (Generative Adversarial Latent Image Processing): (Tao et al., 2023) is a generative model that produces high-quality images by leveraging adversarial techniques.
 It utilizes latent image processing to enhance both the quality and fidelity of generated content.
@@
 Through iterative training, GALIP achieves highly detailed and realistic image synthesis.
-DFGAN (Deep Fusion Generative Adversarial Network): (Tao et al.
-2022) focuses on generating images from text descriptions using deep fusion techniques.
+  • DFGAN (Deep Fusion Generative Adversarial Network): (Tao et al., 2022) focuses on generating images from text descriptions using deep fusion techniques.
 The model integrates textual information into the image generation process through multiple layers of fusion, ensuring alignment with input descriptions.
 DFGAN operates in two stages: the first generates a coarse image based on the text, and the second refines the image, enhancing detail and coherence.
-Stable Diffusion v1.4 (SD_V14): (Rombach et al.
-2022) is part of the Stable Diffusion family, designed for image synthesis through iterative diffusion processes.
+  • Stable Diffusion v1.4 (SD14): (Rombach et al., 2022) is part of the Stable Diffusion family, designed for image synthesis through iterative diffusion processes.
 Starting with pure noise, the model refines it into coherent images through a series of steps, guided by a learned model that predicts noise distribution.
-SD_V14 is recognized for producing high-resolution images with fine textures and complex structures.
-Stable Diffusion v1.5 (SD_V15): (Rombach et al.
-2022) builds on SD_V14, introducing enhancements in architecture and training techniques to improve image quality, consistency, and the handling of complex scenes.
-While continuing to use the diffusion process, SD_V15 incorporates better optimization strategies and larger datasets, yielding superior results.
-Versatile Diffusion: (Xu et al.
-2023b) is a diffusion-based model designed for a variety of image generation tasks, including style transfer, image inpainting, and super-resolution.
+SD14 is recognized for producing high-resolution images with fine textures and complex structures.
+  • Stable Diffusion v1.5 (SD15): (Rombach et al., 2022) builds on SD14, introducing enhancements in architecture and training techniques to improve image quality, consistency, and the handling of complex scenes.
+While continuing to use the diffusion process, SD15 incorporates better optimization strategies and larger datasets, yielding superior results.
+  • Versatile Diffusion: (Xu et al., 2023b) is a diffusion-based model designed for a variety of image generation tasks, including style transfer, image inpainting, and super-resolution.
 Its versatility lies in its ability to adapt the diffusion process to the specific requirements of each task, making it suitable for diverse applications.
-Stable Diffusion v2.1 (SD_V21): (Rombach et al.
-2022) is an advanced iteration of the Stable Diffusion series, offering significant improvements in image fidelity and generation speed.
-With optimizations in the diffusion process, better noise handling, and enhanced training algorithms, SD_V21 produces more realistic and high-quality images, making it highly effective for a range of creative and practical uses.
-Stable Diffusion XL (SD_XL): (Rombach et al.
-2022) represents the most advanced model in the Stable Diffusion series, capable of generating extremely high-resolution images with intricate details.
-SD_XL uses an extended number of diffusion steps and a larger network architecture to manage increased complexity, making it ideal for applications requiring ultra-high resolution and precision, such as detailed artworks and large-scale prints.
-DALL·E 2: (Ramesh et al.
-2022) is a generative model developed by OpenAI that excels in creating images from textual descriptions.
-As an advancement over the original DALL·E, it offers improved capabilities in generating high-quality, diverse, and coherent images.
-DALL·E 2 combines CLIP (Contrastive Language-Image Pre-training) with a transformer-based generative model to translate complex and abstract textual inputs into visual content, making it renowned for its creative and realistic outputs.
+  • Stable Diffusion v2.1 (SD21): (Rombach et al., 2022) is an advanced iteration of the Stable Diffusion series, offering significant improvements in image fidelity and generation speed.
+With optimizations in the diffusion process, better noise handling, and enhanced training algorithms, SD21 produces more realistic and high-quality images, making it highly effective for a range of creative and practical uses.
+  • Stable Diffusion XL (SDxl): (Rombach et al., 2022) represents the most advanced model in the Stable Diffusion series, capable of generating extremely high-resolution images with intricate details.
+SDxl uses an extended number of diffusion steps and a larger network architecture to manage increased complexity, making it ideal for applications requiring ultra-high resolution and precision, such as detailed artworks and large-scale prints.
+  • DALL·E 2 (D·E 2): (Ramesh et al., 2022) is a generative model by OpenAI designed to create images from textual descriptions.
+It produces high-quality, diverse, and coherent outputs by combining CLIP (Contrastive Language-Image Pre-training) with a transformer-based generative architecture.
+DALL·E 2 excels at translating complex and abstract textual prompts into visually captivating content, establishing itself as a benchmark model in text-to-image generation for its creativity and realism.
+  • DALL·E 3 (D·E 3): (OpenAI, 2024) is the latest evolution in OpenAI’s text-to-image generative models, offering significant advancements over DALL·E 2.
+It excels in interpreting nuanced and detailed prompts, producing outputs that are more coherent, visually appealing, and aligned with the given descriptions.
+DALL·E 3 integrates seamlessly with GPT-4, leveraging its advanced language understanding to ensure a deeper alignment between textual input and generated images.
+This integration enhances both the quality and controllability of image generation, making DALL·E 3 a powerful tool for a wide range of creative and professional applications.
 These generative models represent significant advancements in AI-driven image synthesis, each with unique strengths tailored to specific applications.
 From text-to-image generation and high-resolution synthesis to versatile image processing, these models push the boundaries of AI-generated content.
-Figure 9: We evaluate eight image-generating models from 13 aspects. The numerical values in the radar chart represent the mean difference rate of each model.
-C Evaluation Metrics
-Overview
+B.4 Image Resolution
+The resolution of the generated images varies across models.
+GALIP produces images at 224x224 pixels, while DF-GAN generates slightly higher-resolution images at 256x256 pixels.
+Models such as various versions of Stable Diffusion and DALL·E 2 generate images at a resolution of 512x512 pixels.
+Notably, DALL·E 3 delivers the highest resolution among these models, producing images at 1024x1024 pixels.
+B.5 Composition
+The D-ANI dataset offers a robust foundation for evaluating the performance and capabilities of various AIGC models under different conditions and guidance modes.
+By encompassing a broad spectrum of images and resolutions, the dataset ensures comprehensive coverage of the potential use cases and challenges associated with AI-generated content.
+The combination of natural and generated images, alongside diverse textual descriptions, facilitates a thorough assessment of image generation models, supporting the development of more accurate and reliable AI systems.
+Table 6: The details of the D-ANI dataset.
+Types DFGAN GALIP SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
+T2I 25,000 25,000 25,000 25,000 25,000 25,000 25,000 5,000 5,000
+I2I - - 25,000 25,000 25,000 25,000 25,000 5,000 -
+TI2I - - 25,000 25,000 25,000 25,000 25,000 - -
+Appendix C Evaluation Metrics
+C.1 Overview
 In this study, we selected a comprehensive set of metrics to evaluate the image quality, alignment, and aesthetics of both natural and AI-generated images.
@@
 To evaluate image quality, we divided the assessment into three sub-aspects: pixel-level similarity, frame-level visual metrics, and visual content distribution.
-  • Pixel-level: Metrics such as PSNR, LPIPS, DISTS, VIF, VSI, FSIM, and MAD assess structural, perceptual, and textural fidelity.
+  • Pixel-level: Metrics such as SSIM, PSNR, LPIPS, DISTS, VIF, VSI, FSIM, and MAD assess structural, perceptual, and textural fidelity.
 These metrics measure how closely the generated images match natural ones in terms of structural integrity and perceptual similarity.
-  • Frame-level: Metrics including PIQE, IL-NIQE, MUSIQ, DBCNN, CNNIQA, CLIPIQA, BRISQUE, TReS, HyperIQA, LIQE, and UNIQUE provide no-reference assessments of image quality.
+  • Frame-level: Metrics including PIQE, IL-NIQE, MUSIQ, DBCNN, CNNIQA, CLIPIQA, BRISQUE, TReS, HyperIQA, LIQE, UNIQUE and QAlign provide no-reference assessments of image quality.
 They evaluate attributes such as distortion levels, natural scene statistics, and learned features to predict the overall quality of the images.
 Additionally, naturalness metrics like NIQE and NRQM help gauge how closely AI-generated images resemble natural scenes based on statistical properties.
-  • Visual Content Distribution: Metrics such as FID and Inception Score examine overall content distribution differences between AI-generated and natural images across the entire dataset.
+  • Visual Content Distribution: Metrics such as FID, CLIP-FID and Inception Score examine overall content distribution differences between AI-generated and natural images across the entire dataset.
 These metrics capture broad differences in content quality.
-For alignment, we utilize the CLIP Score (Radford et al.
-2021b) to measure the correspondence between text descriptions and the generated images.
+For alignment, we utilize the CLIP Score (Radford et al., 2021a) to measure the correspondence between text descriptions and the generated images.
 This metric evaluates how well the generated content aligns with the given prompts by embedding both text and image into a shared space and calculating their similarity.
-In terms of aesthetic evaluation, we employ NIMA and LAION-AES metrics to assess the visual appeal and artistic quality of the images.
+In terms of aesthetic evaluation, we employ NIMA, LAION-AES and QAlign metrics to assess the visual appeal and artistic quality of the images.
 These metrics predict human perception of image aesthetics based on deep learning models and various visual features.
@@
 The chosen metrics, as detailed in Table 7, collectively provide a comprehensive toolkit for evaluating image quality, alignment, and aesthetics, each focusing on different aspects of visual perception and fidelity.
-Naive Image Quality
+C.2 Naive Image Quality
 We evaluated the naive image quality of AI-generated images versus natural images using a range of metrics.
@@
 Pixel-Level Quality Metrics:
-  • SSIM (Wang et al.
-2004) (Structural Similarity Index) compares the structural information between a reference and a test image, considering luminance, contrast, and structure for a comprehensive measure of similarity.
+  • SSIM (Wang et al., 2004) (Structural Similarity Index) compares the structural information between a reference and a test image, considering luminance, contrast, and structure for a comprehensive measure of similarity.
   • PSNR (Peak Signal-to-Noise Ratio) measures the ratio between the maximum possible power of an image and the power of corrupting noise, with higher values indicating better image quality.
-  • VIF (Sheikh and Bovik 2006) (Visual Information Fidelity) quantifies the amount of visual information preserved in the test image relative to the reference, based on natural scene statistics and the human visual system.
-  • VSI (Zhang, Shen, and Li 2014) (Visual Saliency-based Index) evaluates perceptual similarity by focusing on how salient (prominent) features in the images compare.
-  • FSIM (Zhang et al. 2011) (Feature Similarity Index) measures similarity using phase congruency and gradient magnitude, emphasizing feature similarity, which is crucial for human perception.
-  • LPIPS (Zhang et al.
-2018) (Learned Perceptual Image Patch Similarity) uses deep network features to assess perceptual similarity, capturing human visual perception more effectively by comparing feature activations in a deep neural network.
-  • DISTS (Ding et al. 2022) (Deep Image Structure and Texture Similarity) combines structural and textural similarity metrics using deep learning features to evaluate overall image similarity.
-  • MAD (Larson and Chandler 2010) (Mean Absolute Deviation) provides a straightforward assessment of pixel-level differences, with lower values indicating greater similarity between images.
+  • VIF (Sheikh & Bovik, 2006) (Visual Information Fidelity) quantifies the amount of visual information preserved in the test image relative to the reference, based on natural scene statistics and the human visual system.
+  • VSI (Zhang et al., 2014) (Visual Saliency-based Index) evaluates perceptual similarity by focusing on how salient (prominent) features in the images compare.
+  • FSIM (Zhang et al., 2011) (Feature Similarity Index) measures similarity using phase congruency and gradient magnitude, emphasizing feature similarity, which is crucial for human perception.
+  • LPIPS (Zhang et al., 2018) (Learned Perceptual Image Patch Similarity) uses deep network features to assess perceptual similarity, capturing human visual perception more effectively by comparing feature activations in a deep neural network.
+  • DISTS (Ding et al., 2022) (Deep Image Structure and Texture Similarity) combines structural and textural similarity metrics using deep learning features to evaluate overall image similarity.
+  • MAD (Larson & Chandler, 2010) (Mean Absolute Deviation) provides a straightforward assessment of pixel-level differences, with lower values indicating greater similarity between images.
 Frame-Level Quality Metrics:
-  • PIQE (N. et al. 2015) (Perception-based Image Quality Evaluator) provides a no-reference quality assessment by analyzing image blocks and estimating distortion levels.
-  • IL-NIQE (Zhang, Zhang, and Bovik 2015) (Integrated Local Natural Image Quality Evaluator) evaluates image quality based on natural scene statistics and local features, offering a no-reference assessment.
-  • MUSIQ (Ke et al. 2021) (Multi-Scale Image Quality) assesses image quality at multiple scales, considering both local and global features for a comprehensive no-reference evaluation.
-  • DBCNN (Zhang et al. 2020) (Deep Bilinear Convolutional Neural Network) predicts image quality using deep learning, based on bilinear pooling of convolutional features.
-  • LIQE (Ma et al. 2018) (Learning-based Image Quality Evaluator) combines traditional quality metrics with learned features for robust no-reference quality assessment.
-  • CNNIQA (Kang et al. 2014) (Convolutional Neural Network Image Quality Assessment) leverages deep convolutional networks for no-reference image quality assessment using learned features.
-  • CLIPIQA (Radford et al. 2021a) (CLIP Image Quality Assessment) uses CLIP model embeddings to assess image quality by comparing the alignment between image and text descriptions.
-  • TReS (Radford et al. 2021a) (Textural and Edge-based Similarity) evaluates image quality by focusing on textural and edge-based features, providing a no-reference assessment.
-  • HyperIQA (Su et al. 2020) (Hyper Network-based Image Quality Assessment) uses a hypernetwork to predict image quality, leveraging deep learning for no-reference quality assessment.
-  • UNIQUE (Zhang et al. 2021) (Universal Quality Index with Deep Features) evaluates unique aspects of image quality using deep learning features, offering a no-reference assessment.
-  • BRISQUE (Mittal, Moorthy, and Bovik 2012) (Blind/Referenceless Image Spatial Quality Evaluator) assesses image quality based on natural scene statistics, providing a no-reference measure.
-  • NIQE (Mittal, Soundararajan, and Bovik 2013) (Natural Image Quality Evaluator) evaluates how natural an image appears by comparing its features with those derived from a dataset of natural images, without needing a reference image.
-  • NRQM (Ma et al.
-2018) (Naturalness Image Quality Metric) assesses how closely an image aligns with the statistical properties of natural scenes based on natural scene statistics and perceptual models.
+  • PIQE (N. et al., 2015) (Perception-based Image Quality Evaluator) provides a no-reference quality assessment by analyzing image blocks and estimating distortion levels.
+  • IL-NIQE (Zhang et al., 2015) (Integrated Local Natural Image Quality Evaluator) evaluates image quality based on natural scene statistics and local features, offering a no-reference assessment.
+  • MUSIQ (Ke et al., 2021) (Multi-Scale Image Quality) assesses image quality at multiple scales, considering both local and global features for a comprehensive no-reference evaluation.
+  • DBCNN (Zhang et al., 2020) (Deep Bilinear Convolutional Neural Network) predicts image quality using deep learning, based on bilinear pooling of convolutional features.
+  • LIQE (Ma et al., 2018) (Learning-based Image Quality Evaluator) combines traditional quality metrics with learned features for robust no-reference quality assessment.
+  • CNNIQA (Kang et al., 2014) (Convolutional Neural Network Image Quality Assessment) leverages deep convolutional networks for no-reference image quality assessment using learned features.
+  • CLIPIQA (Radford et al., 2021b) (CLIP Image Quality Assessment) uses CLIP model embeddings to assess image quality by comparing the alignment between image and text descriptions.
+  • TReS (Radford et al., 2021b) (Textural and Edge-based Similarity) evaluates image quality by focusing on textural and edge-based features, providing a no-reference assessment.
+  • HyperIQA (Su et al., 2020) (Hyper Network-based Image Quality Assessment) uses a hypernetwork to predict image quality, leveraging deep learning for no-reference quality assessment.
+  • UNIQUE (Zhang et al., 2021) (Universal Quality Index with Deep Features) evaluates unique aspects of image quality using deep learning features, offering a no-reference assessment.
+  • BRISQUE (Mittal et al., 2012) (Blind/Referenceless Image Spatial Quality Evaluator) assesses image quality based on natural scene statistics, providing a no-reference measure.
+  • NIQE (Mittal et al., 2013) (Natural Image Quality Evaluator) evaluates how natural an image appears by comparing its features with those derived from a dataset of natural images, without needing a reference image.
+  • NRQM (Ma et al., 2018) (Naturalness Image Quality Metric) assesses how closely an image aligns with the statistical properties of natural scenes based on natural scene statistics and perceptual models.
+  • QAlign (Wu et al., 2024b) evaluates image quality by assessing its alignment with semantic and perceptual features indicative of high visual fidelity.
 Visual Content Distribution Metrics:
-  • FID (Heusel et al.
-2017) (Fréchet Inception Distance) measures the similarity between two sets of images by comparing the means and covariances of their feature representations, providing a comprehensive assessment of both quality and diversity.
-  • Inception Score (Salimans et al.
-2016) evaluates the quality of generated images by considering the confidence of the classifier and the diversity of the generated samples, with higher scores indicating better quality and diversity.
-We present the additional results of the Naive Image Quality assessment in Figure 9.
-Alignment
-To assess alignment, we employ the CLIP Score (Contrastive Language-Image Pre-training Score), which measures the correspondence between text descriptions and generated images.
+  • FID (Heusel et al., 2017) (Fréchet Inception Distance) measures the similarity between two sets of images by comparing the means and covariances of their feature representations, providing a comprehensive assessment of both quality and diversity.
+  • CLIP-FID (Kynkäänniemi et al., 2023) (CLIP Fréchet Inception Distance) is an extension of FID that measures the similarity between two sets of images by utilizing CLIP’s multi-modal feature space.
+Instead of relying solely on Inception-based features, CLIP-FID leverages the semantic-rich embeddings from CLIP, which are aligned across image and text modalities.
+This provides a more nuanced assessment of image quality and semantic alignment, particularly for generative models designed to capture complex textual descriptions.
+  • Inception Score (Salimans et al., 2016) evaluates the quality of generated images by considering the confidence of the classifier and the diversity of the generated samples, with higher scores indicating better quality and diversity.
+C.3 Alignment
+To assess alignment, we employ the CLIP Score (Contrastive Language-Image Pre-training Score, CS), which measures the correspondence between text descriptions and generated images.
 This metric evaluates how well the visual content aligns with the textual prompts by embedding both into a shared space and calculating their cosine similarity.
-Aesthetic
-  • NIMA (Esfandarani and Milanfar 2018) (Neural Image Assessment) is a deep learning-based model that predicts the aesthetic quality of images.
+C.4 Aesthetic
+  • NIMA (Esfandarani & Milanfar, 2018) (Neural Image Assessment) is a deep learning-based model that predicts the aesthetic quality of images.
 It generates a score that reflects human perception of visual appeal.
-  • LAION-AES (Schuhmann et al. 2022) (LAION Aesthetic Score) evaluates images’ aesthetic quality using various visual features, providing a numerical value to indicate their aesthetic appeal.
+  • LAION-AES (Schuhmann et al., 2022a) (LAION Aesthetic Score) evaluates images’ aesthetic quality using various visual features, providing a numerical value to indicate their aesthetic appeal.
+  • QAlign (Wu et al., 2024b) assesses image aesthetics by analyzing its adherence to human visual preferences and artistic principles, leveraging semantic coherence and appeal.
 Together, these metrics offer a robust toolkit for evaluating images, focusing on alignment with prompts and aesthetic quality, thereby covering key aspects of visual perception and fidelity.
-Aspects Sub_aspects Netrucs Reference
+Table 7: The detail of evaluation metrics.
+Aspects Sub aspects Metrics Reference
 Naive Quality Pixel-level SSIM TRUE
@@
 NRQM FALSE
+QAlign (quality) FALSE
 Content Distribution FID TRUE
+CLIP-FID TRUE
 Inception Score TRUE
@@
 LAION-AES FALSE
-Table 9: The detail of evaluation metrics.
-D Human Perceptual Evaluation
-Interface for Human Evaluation
-The human evaluation of our DNAI dataset was conducted using a custom-designed interface (Figure 11).
+QAlign (aesthetic) FALSE
+Figure 12: We evaluate nine image-generating models from 14 aspects. The numerical values in the radar chart represent the mean difference rate of each model and each metric.
+Appendix D Evaluation Results
+D.1 Naive Quality
+We present the additional results of the Naive Image Quality assessment in Figure 12.
+D.2 Applicability
+We put the applicability results of Image Recognition and Semantic Segmentation in Tables 8 and  9.
+Table 8: Image Recognition Discrepancy Rate ($DR_{re}$).
+GALIP DFGAN SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
+T2I 68.60 85.46 66.14 65.50 68.12 64.88 64.88 65.36 69.26
+I2I - - 94.29 94.07 40.90 83.30 27.96 47.98 -
+TI2I - - 61.62 59.71 60.05 55.54 27.46 - -
+Table 9: IoU of Object Segmentation.
+GALIP DFGAN SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
+T2I 0.23 0.24 0.26 0.25 0.26 0.25 0.26 0.26 0.28
+I2I - - 0.38 0.38 0.69 0.50 0.82 0.42 -
+TI2I - - 0.43 0.43 0.29 0.53 0.82 - -
+D.3 Human Perceptual Evaluation
+D.3.1 Interface for Human Evaluation
+The human evaluation of our D-ANI dataset was conducted using a custom-designed interface (Figure 14).
 This interface allowed participants to evaluate AI-generated images in comparison to reference content across three key aspects: Quality, Alignment, and Aesthetics.
 The reference content included prompts used to generate the images, categorized into three types: 1) text, 2) image, and 3) text and image.
-The AI-generated images evaluated were produced by four different models: SD_V2.1, SD_XL, Versatile Diffusion, and DALL·E 2.
-Evaluation Process
+The AI-generated images evaluated were produced by four different models: SD21, SDxl, Versatile Diffusion, and DALL·E 2.
+D.3.2 Evaluation Process
 A total of 58 volunteers participated in the evaluation process.
@@
 Only data from participants who showed consistent scoring were considered valid, resulting in 47 effective samples.
-Participant Demographics
+D.3.3 Participant Demographics
 The volunteers who participated in the evaluation ranged in age from 19 to 46 years old, with a gender ratio of 6:4 (male to female).
@@
 Upon completing their evaluations, participants were also invited to provide feedback on the evaluation system and their experience with assessing the differences between natural and AI-generated images.
-Evaluation Metrics
+D.3.4 Evaluation Metrics
 Participants were instructed to evaluate each image based on three aspects:
@@
 These examples illustrated what constituted good or poor performance in each aspect without giving specific scores, thereby preventing any potential bias in participant evaluations.
-Summary
-The comprehensive human evaluation process ensured that the collected data was reliable and reflected genuine human perception.
-The demographic diversity and the rigorous validation of participant responses provided a robust foundation for assessing the performance of various AIGC models in generating high-quality, aligned, and aesthetically pleasing images.
-This approach not only highlighted the strengths and weaknesses of the evaluated models but also offered valuable insights into the perceptual differences between AI-generated and natural images.
 Figure 13: Examples for unsafe references.
@@
 Figure 15: Before the evaluation process, the participant will be given some examples to learn which image should be good or bad in a specific aspect.
-E More Evaluations
+Figure 16: The differential rate of data from different categories models across naive image quality, alignment, and aesthetic.
+Figure 17: The FID, CLIP-FID and Inception Score (IS) of AI-generated images from different source models and categories from content distribution level.
+D.3.5 Summary
+The comprehensive human evaluation process ensured that the collected data was reliable and reflected genuine human perception.
+The demographic diversity and the rigorous validation of participant responses provided a robust foundation for assessing the performance of various AIGC models in generating high-quality, aligned, and aesthetically pleasing images.
+This approach not only highlighted the strengths and weaknesses of the evaluated models but also offered valuable insights into the perceptual differences between AI-generated and natural images.
+D.4 Discrepancies Across Categories
 This section provides a detailed evaluation of AI-generated images across different categories, focusing on naive image quality, alignment, aesthetics, pixel-level similarity, and content distribution.
 These analyses aim to uncover discrepancies between AI-generated and natural images and identify category-specific challenges.
-From Figure 13, we observe notable variability in the differential rates across categories such as ”indoor” and ”sports” among different generative models.
-Images in the ”indoor” category exhibit significantly higher discrepancies in alignment (CLIP Score) and aesthetics (LAION-AES) compared to ”sports”, indicating that AI-generated sports images are more aligned with natural images.
+From Figure 16, we observe notable variability in the differential rates across categories such as ‘indoor’ and ‘sports’ among different generative models.
+Images in the ‘indoor’ category exhibit significantly higher discrepancies in alignment (CLIP Score) and aesthetics (LAION-AES) compared to ‘sports’, indicating that AI-generated sports images are more aligned with natural images.
 In contrast, generating natural-looking indoor images remains a challenge due to their complex structures and diverse visual features.
-Figure 14 presents trends in SSIM and FID metrics across categories and models, where SSIM evaluates pixel-level similarity, and FID measures distribution-level content alignment.
-The ”accessory” category consistently shows higher FID values, reflecting poorer content distribution, while the ”animal” category achieves higher SSIM scores, demonstrating better pixel-level similarity.
-These findings suggest that ”accessory” and ”indoor” images pose greater challenges for generative models in maintaining semantic consistency and achieving balanced content generation compared to other categories.
+Figure 17 presents trends in SSIM, FID, CLIP-FID, and Inception Score.
+The ‘accessory’ category consistently shows higher FID values, reflecting poorer content distribution, while the ‘animal’ category achieves higher SSIM scores, demonstrating better pixel-level similarity.
+These findings suggest that ‘accessory’ and ‘indoor’ images pose greater challenges for generative models in maintaining semantic consistency and achieving balanced content generation compared to other categories.
+Additionally, the results in Figures 17 and Figures 16 show that compared to other generative models, images generated by D·E 3 display greater differences from natural reference images across various metrics.
+This might be due to D·E 3’s tendency to produce high-quality images with a surreal style, whereas other models primarily focus on generating images with a more realistic style (see Appendix Sec.
+B, Figure 9 for visual comparisons).
 Overall, these results indicate that AI-generated images still lag behind natural images in achieving comparable levels of naturalness and image quality.
@@
 This imbalance likely limits their ability to generalize effectively across diverse content types, underscoring the need for more balanced datasets and targeted advancements in model design.
-Figure 13: The differential rate of data from different categories models across naive image quality, alignment, and aesthetic.
-Figure 14: The FID and SSIM of AI-generated images from different source models and categories. The FID and SSIM are used to evaluate image quality at the pixel level and content distribution level.
-F Discussion
-Safety Mechanism Activation in SD 1.4, 1.5 DALL·E 2
-T2I I2I TI2I T2I(%) I2I(%) TI2I(%)
-SD v1.4 104 263 141 4.16 10.51 5.64
-SD v1.5 133 270 145 5.32 10.79 5.80
-DALL·E 2 20 - - 4 - -
-Table 10: Unsafe Prompt.
+Appendix E Discussion
+Table 10: Prompts trigger the safe-checker of different models.
+T2I I2I TI2I T2I(‰) I2I(‰) TI2I(‰)
+SD14 104 263 141 4.16 10.51 5.64
+SD15 133 270 145 5.32 10.79 5.80
+D·E 2 48 - - 9.60 - -
+D·E 3 45 - - 9.00 - -
+Unexpected Safety Mechanism Activation.
 In the rapidly evolving field of AI-generated content (AIGC), the balance between model robustness and safety mechanisms remains a critical area of study.
-A noteworthy issue has emerged with popular image synthesis models, such as Stable Diffusion v1.4, v1.5, and DALL·E 2, which occasionally fail to generate images in response to seemingly benign prompts.
+A noteworthy issue has emerged with popular image synthesis models, such as Stable Diffusion v1.4, v1.5, DALL·E 2 and DALL·E 3, which occasionally fail to generate images in response to seemingly benign prompts.
 Some examples of these prompts are shown in Figure 13.

### v3

@@
-D-Judge: How Far Are We? Accessing the Discrepancies Between AI-synthesized Images and Natural Images through Multimodal Guidance
-Abstract
-In Artificial Intelligence Generated Content (AIGC), distinguishing AI-synthesized images from natural ones remains a key challenge.
-Despite advancements in generative models, significant discrepancies persist.
-To systematically investigate and quantify these discrepancies, we introduce an AI-Natural Image Discrepancy accessing benchmark (D-Judge) aimed at addressing the critical question: how far are AI-generated images (AIGIs) from truly realistic images?
-We construct D-ANI, a dataset with 5,000 natural images and over 440,000 AIGIs generated by nine models using Text-to-Image (T2I), Image-to-Image (I2I), and Text and Image-to-Image (TI2I) prompts.
-Our framework evaluates the discrepancy across five dimensions: naive image quality, semantic alignment, aesthetic appeal, downstream applicability, and human validation.
-Results reveal notable gaps, emphasizing the importance of aligning metrics with human judgment.
-Source code and datasets are available at https://shorturl.at/l83W2.
-Artificial Intelligence Generated Content, AI-synthesized Images, Multimodal, Discrepancy, Evaluation Framework
-1 Introduction
+D-Judge: How Far Are We? Accessing the Discrepancies Between AI-synthesized and Natural Images through Multimodal Guidance
+Abstract.
+In the rapidly evolving field of Artificial Intelligence Generated Content (AIGC), a central challenge is distinguishing AI-synthesized images from natural ones.
+Despite the impressive capabilities of advanced generative models in producing visually compelling images, significant discrepancies remain when compared to natural images.
+To systematically investigate and quantify these differences, we construct a large-scale multimodal dataset, D-ANI, comprising 5,000 natural images and over 440,000 AIGI samples generated by nine representative models using both unimodal and multimodal prompts, including Text-to-Image (T2I), Image-to-Image (I2I), and Text and Image-to-Image (TI2I).
+We then introduce an AI–Natural Image Discrepancy assessment benchmark (D-Judge) to address the critical question: how far are AI-generated images (AIGIs) from truly realistic images?
+Our fine-grained evaluation framework assesses the D-ANI dataset across five dimensions: naive visual quality, semantic alignment, aesthetic appeal, downstream task applicability, and coordinated human validation.
+Extensive experiments reveal substantial discrepancies across these dimensions, highlighting the importance of aligning quantitative metrics with human judgment to achieve a comprehensive understanding of AI-generated image quality.
+The code [^1] https://github.com/ryliu68/DJudge and data [^2] https://huggingface.co/datasets/Renyang/DANI are publicly available.
+Discrepancy Evaluation Benchmark, Image Naturalness Assessment, Multimodal Guided Generation, AI-generated images, AIGC
+copyright: acmlicensedjournalyear: 2025copyright: ccdoi: 10.1145/3746027.3754837conference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Computer visionccs: Computing methodologies Neural networksccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learning
+1. Introduction
 With the rapid advancement of deep learning techniques, the proliferation of Artificial Intelligence Generated Content (AIGC) has garnered significant attention across various domains, such as e-commerce, gaming, medicine, animation, and autonomous driving (Li et al., 2024; Qian et al., 2024).
-AI-generated Images (AIGI) have been one of the mainstream forms of AIGC, and a variety of AI image generative models have been proposed to make the generated synthetic images as realistic as natural images, ranging from the earlier generative adversarial networks (GANs) (Tao et al., 2023, 2022), advanced diffusion models (DMs) (Xu et al., 2023b; Wei et al., 2023) to the large multimodal generative model like DALL·E (Ramesh et al., 2022).
-Figure 1: Toy example demonstrating the key aspects of D-Judge.
-Despite the rapid development of AI image generative models, AI-generated images still fall short of real-world application standards due to discrepancies with realistic natural images (Li et al., 2023a).
-To address this, various AI-generated image quality assessment methods have emerged (Wang et al., 2023; Xu et al., 2023a; Hu et al., 2023).
+AI-generated Images (AIGI) have been one of the mainstream forms of AIGC, and a variety of AI image generative models have been proposed to make the generated synthetic images as realistic as natural images, ranging from the earlier generative adversarial networks (GANs) (Tao et al., 2023, 2022), advanced diffusion models (DMs) (Xu et al., 2023; Wei et al., 2023) to the large multimodal generative model like DALL·E (Ramesh et al., 2022).
+Despite the rapid development of AI image generative models, AI-generated images still fall short of real-world application standards due to discrepancies with realistic natural images (Li et al., 2024).
+To address this, various AI-generated image quality assessment methods have emerged (Yuan et al., 2023; Xu et al., 2023; Hu et al., 2023).
 For instance, Wang et al.
-(Wang et al., 2023) introduced the AIGCIQA2023 database and a subjective evaluation framework based on quality, authenticity, and text-image correspondence.
+(Yuan et al., 2023) introduced the AIGCIQA2023 database and a subjective evaluation framework based on quality, authenticity, and text-image correspondence.
 Similarly, PKU-I2IQA (Yuan et al., 2023) developed a perception-based database for image-to-image generation using no-reference and full-reference methods.
-Other studies, such as Pick-a-Pic (Kirstain et al., 2023), HPS v2 (Wu et al., 2023), TIFA (Hu et al., 2023), and ImageReward (Xu et al., 2023a), focused on training unified score models for automatic AIGC image quality assessment, moving beyond traditional metrics like Inception Score (Salimans et al., 2016) and FID (Heusel et al., 2017).
+Other studies, such as Pick-a-Pic (Kirstain et al., 2023), HPS v2 (Wu et al., 2023), TIFA (Hu et al., 2023), and ImageReward (Xu et al., 2023), focused on training unified score models for automatic AIGC image quality assessment, moving beyond traditional metrics like Inception Score (Salimans et al., 2016) and FID (Heusel et al., 2017).
 Despite advancements in AIGC research, several critical issues remain unresolved.
@@
 Second, previous research predominantly considers unimodal prompts, such as T2I or I2I, while neglecting multimodal prompts like TI2I, and the datasets used are often small, typically comprising only a few thousand images, which limits their capacity for comprehensive analysis and evaluation.
+Figure 1. Toy example demonstrating the key aspects of D-Judge.
 To address the above challenges, we proposed an AI-Natural Image Discrepancy Evaluation Benchmark (D-Judge) to investigate and interpret what discrepancies still remained between AI-generated images and natural images, finally answering the important question “How far are AI generative models with respect to the visual forms of images?”.
@@
 The data size of our DNAI dataset scales to 100x that of prior datasets.
-Fine-grained Evaluation Framework, we propose a fine-grained evaluation framework to conduct the systematic and comprehensive assessment and evaluation of DNAI, covering 5 diverse and important aspects, including naive visual feature quality, semantic alignment among multimodal generation, aesthetic appeal, downstream applicability, and the coordinated human validation.
+Fine-grained Evaluation Framework, we propose a fine-grained evaluation framework to conduct the systematic and comprehensive assessment and evaluation of DNAI, covering 5 diverse and important aspects, including naive visual feature quality, semantic alignment among multimodal generation, aesthetic appeal, downstream applicability, and coordinated human validation.
 The toy example is illustrated in Figure 1.
@@
 It validates the necessity of incorporating human evaluation for coordination.
-2 Related Work
+2. Related Work
 Natural Image Evaluation:
@@
 et al., 2015).
-These methods have focused on various visual features and properties, such as perceptual appearance (Zhang et al., 2018), naturalness (Ma et al., 2018), and aesthetics (Esfandarani & Milanfar, 2018).
+These methods have focused on various visual features and properties, such as perceptual appearance (Zhang et al., 2018), naturalness (Ma et al., 2018), and aesthetics (Esfandarani and Milanfar, 2018).
 For instance, BRISQUE evaluates natural image quality by analyzing spatial domain features, while PIQE assesses perceptual quality through block-based image segmentation, both serving as effective no-reference metrics.
-Other notable measures include FID and Inception Scores, which assess the quality and diversity of generated images, and SSIM (Wang et al., 2004) and PSNR, which quantify structural similarity and pixel-level accuracy.
+Other notable measures include FID and Inception Scores, which assess the quality and diversity of generated images, and SSIM (Wang et al., 2004) and PSNR, which quantify structural similarity and structure-level accuracy.
 Beyond quantitative measures, human evaluations have also been leveraged to estimate natural image quality.
-For example, NIMA (Esfandarani & Milanfar, 2018), proposed by Talebi and Milanfar, uses a deep CNN trained on human-rated images to predict aesthetic quality.
+For example, NIMA (Esfandarani and Milanfar, 2018), proposed by Talebi and Milanfar, uses a deep CNN trained on human-rated images to predict aesthetic quality.
 Similarly, Wong et al.
@@
 Several methods and benchmarks have been introduced to evaluate AI-generated images (Zhang et al., 2023; Hu et al., 2023).
-These methods primarily focus on three aspects: perceptual quality, text-image correspondence, and aesthetics (Xu et al., 2023a; Wang et al., 2023), often training unified models to assess the overall quality of AI-generated images.
-For example, ImageReward (Xu et al., 2023a) presents a framework for aligning image generation with human preferences, AGIQA-3k (Li et al., 2023a) offers a comprehensive benchmark for assessing AI-generated image quality, and QBench (Wu et al., 2024a) evaluates content quality across multiple dimensions.
+These methods primarily focus on three aspects: perceptual quality, text-image correspondence, and aesthetics (Xu et al., 2023; Yuan et al., 2023; Li et al., 2024, 2025; Chen et al., 2025), often training unified models to assess the overall quality of AI-generated images.
+For example, ImageReward (Xu et al., 2023) presents a framework for aligning image generation with human preferences, AGIQA-3k (Li et al., 2024) offers a comprehensive benchmark for assessing AI-generated image quality, and QBench (Wu et al., 2024) evaluates content quality across multiple dimensions.
 Despite advancements, little research has delved into the fine-grained differences between AI-generated and natural images, and existing datasets are often too small for comprehensive evaluation.
 To address these gaps, we propose D-Judge as AI-Natural image difference Assessment Benchmark, featuring a large Distinguishing AI-Natural Image Dataset and a systematic, fine-grained evaluation framework to thoroughly explore these differences.
-3 D-Judge
-Figure 2: The D-ANI dataset of D-Judge, guided by three types of prompts. Each generated image is paired with both the corresponding text and image from the naturally collected dataset (e.g., COCO).
-Figure 3: The framework of D-Judge evaluates the Discrepancy Rate ($DR$) between AI-generated images and natural reference images across five dimensions: ➊ Naive Quality, ➋ Alignment, ➌ Aesthetic, and ➍ Applicability, complemented by ➎ human validation.
-The images with green borders represent reference natural images, while those with yellow borders are AI-generated counterparts.
-This comprehensive evaluation integrates pixel-level, frame-level, and distribution-level quality checks, alongside advanced tools like OpenAI CLIP, OpenAI Laion Aesthetic Evaluator, Qalign, ViT, U2Net, and BLIP2 to robustly assess the gap between the two image types.
-Where $CS$ means CLIP Score for the text and given images, and $AES$ means Aesthetic score.
+3. D-Judge
 To tackle the challenges of current AI-generated image evaluation, we construct an AI-Natural Image Discrepancy evaluation benchmark (D-Judge) to assess the potential discrepancies between AI-generated images and natural images and answer how far AIGIs are from Natural Images.
@@
 In the following parts, we introduce the D-ANI Dataset and the evaluation framework in detail.
-3.1 Distinguishing AI-Natural Image Dataset
-We construct the Distinguished Natural and AI-generated Image Dataset based on the classical natural image dataset MS COCO dataset.
-We select 5,000 images covering diverse captions from the MS COCO dataset and collect a total of 25,000 text-image pairs by pairing each image with five different pieces of text descriptions.
-The 25,000 text-image pairs are seen as the referenced natural image set.
-For each text-image sample in the referenced natural image set, we collect AI-generated images from 9 representative generative models, guided by three types of prompts, including text-only prompts, image-only prompts, and text-image prompts.
-The 9 representative generative models are from diverse generative models ranging from the earlier generative adversarial networks (GANs) (e.g.
-GALIP (Tao et al., 2023), DF-GAN (Tao et al., 2022)) to the recent diffusion models (DMs) like Stable Diffusion v1.4 (SD14), v1.5 (SD15), v2.1 (SD21), XL (SDxl) (Rombach et al., 2022), Versatile-Diffusion (VD) (Xu et al., 2023b), OpenAI DALL·E 2 (D·E 2) (Ramesh et al., 2022) and DALL·E 3 (D·E 3) (OpenAI, 2024).
-The referenced natural images and the generated images comprise our distinguishing AI-Natural image dataset.
-More details about data construction and collection are presented in Appendix Sec.
-B.
-Table 1 shows the statistics and properties of our D-ANI Dataset compared with the existing evaluation dataset.
-Our D-ANI Dataset has a total of 440,000 AI-generated images from 9 representative generated models.
-For details, please refer to Appendix Sec.
-B.
-Compared with state-of-the-art evaluation datasets, our D-ANI Dataset has convincing contributions:
-  • D-ANI Dataset is the most extensive dataset for AI-generated image evaluation, and the data size scales to 100x.
-  • D-ANI Dataset is guided by both unimodal prompts and multi-modal prompts (text-image). Prior datasets barely considered text prompts.
-  • Comprehensive and Diverse generative models are exploited, ranging from the earlier GAN models, the recent popular diffusion models, and the influential large commercial generative model DALL·E 2 and DALL·E 3 issued by OpenAI.
-Our large and comprehensive dataset works as the foundation of the assessment benchmark and enables thorough evaluation and analysis of the differences between AI-generated images and realistic natural images.
-Table 1: Comparison of D-ANI dataset and Counterparts.
+3.1. D-ANI Dataset
+Figure 2. The D-ANI dataset of D-Judge, guided by three types of prompts. Each generated image is paired with both the corresponding text and image from the naturally collected dataset (e.g., COCO).
+3.1.1. Natural Images
+We constructed the Distinguished Natural and AI-generated Image dataset (D-ANI) based on the classical MS COCO natural image dataset (Lin et al., 2014).
+Specifically, we selected 5,000 carefully curated images from the MS COCO validation set, each accompanied by at least five unique captions, resulting in a total of 25,000 distinct text-image pairs.
+This rich textual diversity enhances the dataset’s applicability across various image-to-text and text-to-image tasks, making it an invaluable resource for evaluating and training image generation and understanding models.
+The 25,000 text-image pairs form the referenced natural image set, guiding generative models in producing new images in our study.
+3.1.2. Generated Images
+We collected AI-generated images from nine representative generative models using three distinct types of prompts: text-only (Text-to-Image, T2I), image-only (Image-to-Image, I2I), and combined text-image (Text-and-Image-to-Image, TI2I).
+These models range from earlier generative adversarial networks (GANs) like GALIP (Tao et al., 2023) and DF-GAN (Tao et al., 2022), to recent diffusion models (DMs) such as Stable Diffusion (versions v1.4, v1.5, v2.1, and XL) (Rombach et al., 2022), Versatile Diffusion (VD) (Xu et al., 2023), and OpenAI’s commercial models DALL·E 2 (D·E 2) (Ramesh et al., 2022) and DALL·E 3 (D·E 3) (OpenAI, 2024), which generate images via OpenAI’s API, while other models generate images locally.
+Detailed information w.r.t these models can be found in Appendix B.
+Each model generated 25,000 images per guidance mode, except for DALL·E 2 and DALL·E 3, which produced 5,000 images per mode, leading to a total of 440,000 images.
+Due to intrinsic guidance capabilities, DF-GAN, GALIP, and DALL·E 3 were exclusively used in T2I mode, whereas DALL·E 2 operated in both T2I and I2I modes.
+The details are shown in Table 1.
+The image resolution varies across these models: GALIP produces images at 224x224 pixels, DF-GAN at 256x256 pixels, models such as various versions of Stable Diffusion and DALL·E 2 at 512x512 pixels, and DALL·E 3 provides the highest resolution of 1024x1024 pixels.
+These generated images, combined with the referenced natural images, constitute our distinguishing AI-Natural image dataset.
+Table 1. The details of the D-ANI dataset.
+Types DFGAN GALIP SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
+T2I 25,000 25,000 25,000 25,000 25,000 25,000 25,000 5,000 5,000
+I2I - - 25,000 25,000 25,000 25,000 25,000 5,000 -
+TI2I - - 25,000 25,000 25,000 25,000 25,000 - -
+Table 2. Comparison of D-ANI dataset and Counterparts.
 Dataset Model Text Image Text vs. Image AIGI
-AGIQA-1K 2 1,080 - - 1,080
-AGIQA-3K 6 300 - - 2,982
-AIGCIQA2023 6 100 - - 2,400
-PKU-I2IQA 2 200 200 - 1,600
-D-ANI (Ours) 9 25,000 5,000 25,000 445,000
-3.2 D-Judge Framework
+AGIQA-1K (Zhang et al., 2023) 2 1,080 - - 1,080
+AGIQA-3K (Li et al., 2024) 6 300 - - 2,982
+AGIQA-20K (Li et al., 2024) 15 200,000 - - 200,000
+AIGCIQA2023 (Wang et al., 2023) 6 100 - - 2,400
+ImageReward (Xu et al., 2023) 3 8,878 - - 136,892
+HPDv2 (Wu et al., 2023) 9 108,000 - - 381,304
+PKU-I2IQA (Yuan et al., 2023) 2 200 200 - 1,600
+D-ANI (Ours) 9 25,000 5,000 25,000 440,000
+Table 2 presents the key statistics and characteristics of our proposed D-ANI dataset in comparison with existing AIGC evaluation datasets.
+Specifically, the D-ANI dataset consists of 440,000 AI-generated images produced by nine representative generative models, alongside 25,000 human-annotated image-text pairs from the MS COCO validation set.
+This results in a total of 445,000 high-quality text-image pairs, supporting both generation and evaluation tasks.
+In contrast to state-of-the-art benchmarks, the D-ANI dataset offers several notable advantages:
+  • It represents the most comprehensive dataset to date for evaluating AI-generated images, expanding the data volume by approximately 100 times compared to existing benchmarks.
+  • It uniquely integrates prompts from both unimodal (text-only or image-only) and multi-modal (text-image combined) inputs, effectively addressing the prevalent limitation of missing multi-modal context in prior datasets.
+  • It encompasses a broad and diverse range of generative models, spanning from early GAN-based architectures to the latest state-of-the-art diffusion techniques, including high-impact commercial systems such as OpenAI’s DALL·E 2 and DALL·E 3.
+The extensive scale and diversity of the D-ANI dataset provide a robust foundation for systematically analyzing the differences between natural and AI-generated images under various conditions and prompt modalities.
+By encompassing a wide spectrum of image styles, resolutions, semantic content, and textual descriptions, the dataset facilitates rigorous and holistic evaluations.
+This, in turn, supports the development and benchmarking of more precise, robust, and trustworthy AI image generation systems for real-world deployment.
+Figure 3.
+The framework of D-Judge evaluates the Discrepancy Rate ($DR$) between AI-generated images and natural reference images across five dimensions: ➊ Naive Quality, ➋ Alignment, ➌ Aesthetic, and ➍ Applicability, complemented by ➎ human validation.
+The images with green borders represent reference natural images, while those with yellow borders are AI-generated counterparts.
+This comprehensive evaluation integrates structure-level, frame-level, and distribution-level quality checks, alongside advanced tools like OpenAI CLIP, OpenAI Laion Aesthetic Evaluator, QAlign, ViT, U2Net, and BLIP2 to robustly assess the gap between the two image types.
+Where $CS$ means CLIP Score for the text and given images, and $AES$ means Aesthetic score.
+3.2. D-Judge Framework
 On top of the D-ANI dataset, we construct the fine-grained D-Judge to systematically measure and interpret the differences between AI-generated images and realistic natural images from five key different aspects as following aspects.
@@
 We leverage traditional image quality assessment methods for natural images to measure the AI-Natural image differences.
-In order to perform fine-grained analysis and exploration, we estimate the AI-Natural image differences from three levels by considering the inherent properties of image-style content, ranging from the pixel-level visual features, frame-level visual features to the holistic content distribution.
-  • First, we use the full-referenced pixel-level similarity measures like SSIM, LPIPS, and DISTS to compute the visual similarities between the paired AI-generated image and natural image and use these calculated values to quantify the AI-Natural image differences from the low-level visual view.
+In order to perform fine-grained analysis and exploration, we estimate the AI-Natural image differences from three levels by considering the inherent properties of image-style content, ranging from the structure-level visual features, frame-level visual features to the holistic content distribution.
+  • First, we use the full-referenced structure-level similarity measures like SSIM, LPIPS, and DISTS to compute the visual similarities between the paired AI-generated image and natural image and use these calculated values to quantify the AI-Natural image differences from the low-level visual view.
   • Second, we use non-referenced frame-level visual features to qualify the perceptional quality of images and compute the frame-level visual quality difference rates between AI-generated images and natural images from the high-level visual view.
 The exploited frame-level visual features include PIQE (N.
-et al., 2015), IL-NIQE (Zhang et al., 2015), MUSIQ (Ke et al., 2021), DBCNN (Zhang et al., 2020), LIQE (Ma et al., 2018), Inception Score (Salimans et al., 2016), CLIPIQA (Radford et al., 2021b), TReS (Gu et al., 2015), HyperIQA (Su et al., 2020), UNIQUE (Zhang et al., 2021), BRISQUE , NIQE (Mittal et al., 2013), NRQM (Ma et al., 2018), etc.
+et al., 2015), IL-NIQE (Zhang et al., 2015), MUSIQ (Ke et al., 2021), DBCNN (Zhang et al., 2020), CLIPIQA (Radford et al., 2021), CNNIQA (Kang et al., 2014), TReS (Gu et al., 2015), HyperIQA (Su et al., 2020), LIQE (Ma et al., 2018), UNIQUE (Zhang et al., 2021), BRISQUE (Mittal et al., 2012) , NIQE (Mittal et al., 2013), NRQM (Ma et al., 2018), QAlign (Wu et al., 2024), Inception Score (IS) (Salimans et al., 2016), etc.
 The Discrepancy Rate ($DR$) of these metrics provides insights into frame-level visual quality differences.
-  • Third, we investigate the full-reference visual content distribution to capture the holistic differences between AI-generated images and natural images by computing the FID, CLIP-FID, and Inception Score of the AI-generated image set and the referenced natural image set.
+  • Third, we investigate the full-reference visual content distribution to capture the holistic differences between AI-generated images and natural images by computing the FID, CLIP-FID, and IS of the AI-generated image set and the referenced natural image set.
 The results of these set-level measures can reflect the difference between AIGIs and natural ones from an integral perspective.
 Based on the different levels of image quality measures, our framework can assess and measure the fine-grained AI-Natural images differences with Full-reference and Non-reference angles, covering both low-level and high-level.
-More details about the utilized traditional image quality assessment measures are described in Appendix Sec.
-C.
+More details about the utilized traditional image quality assessment measures are described in Appendix C.
 2. Semantic Alignment Discrepancy:
@@
 Therefore, semantic alignment can be an important quality indicator of the AIGIs.
-We use the widely used CLIP model (Radford et al., 2021a) to calculate the difference rates of Clip Scores, $DR_{CS}$ to represent alignment discrepancies between AIGIs and natural images, and provide a fine-grained analysis with respect to both unimodal and multimodal guidance.
+We use the widely used CLIP model to calculate the difference rates of CLIP Scores (Hessel et al., 2021), $DR_{CS}$ to represent alignment discrepancies between AIGIs and natural images, and provide a fine-grained analysis with respect to both unimodal and multimodal guidance.
 3. Aesthetic Appeal Discrepancy:
 Aesthetic Appeal is to estimate the visual appeal and artistic quality of images, reflecting the visual attractiveness and artistic quality of images.
-We utilize the classical aesthetic measures NIMA (Esfandarani & Milanfar, 2018), LAION-AES (Schuhmann et al., 2022a), and QAlign (Wu et al., 2024b) as the quantitative metrics and compare the difference rates of aesthetic scores $DR_{aes}$ for the paired AI-generated and natural images.
+We utilize the classical aesthetic measures NIMA (Esfandarani and Milanfar, 2018), LAION-AES (Schuhmann et al., 2022), and QAlign (Wu et al., 2024) as the quantitative metrics and compare the difference rates of aesthetic scores $DR_{aes}$ for the paired AI-generated and natural images.
 4. Downstream Applicability Discrepancy:
@@
 For VQA, we design three types of questions: ‘object’ (identifying objects in the image), ‘alignment’ (determining if the image aligns with a textual description, returning ‘yes’ or ‘no’), and ‘similarity probability’ (calculating the alignment probability between text and image).
-Using BLIP2 (Li et al., 2023b), we evaluate the Difference Rate ($DR$) between AI-generated images and reference natural images for each question type.
+Using BLIP2 (Li et al., 2023), we evaluate the Difference Rate ($DR$) between AI-generated images and reference natural images for each question type.
 5. Human Validation:
@@
 Human participants are provided with example images before the evaluation to understand high and low scores without being given specific numerical values, ensuring unbiased and informed ratings.
-Detailed information about human assessment procedures is provided in Appendix Sec.
-D.3.
+Detailed information about human assessment procedures is provided in Sec.
+4.6.
 Through the integration of diverse and comprehensive evaluated dimensions, our D-Judge offers a systematic assessment solution to investigate and interpret the differences that still remain between AI-generated images and natural images.
-Figure 4: We evaluate nine image-generating models from 6 aspects. The numerical values in the radar chart represent the mean difference rate of each model.
-4 Evaluation
+Figure 4. We evaluate nine image-generating models from 6 aspects. The numerical values in the radar chart represent the mean difference rate of each model.
+4. Evaluation
 With respect to the 5 different aspects, we conducted benchmark experiments and performed experimental evaluation alongside the following research questions.
 RQ1: What are the fine-grained naive image quality discrepancies between AIGIs and natural images?
-  • RQ1.a: What discrepancies between I2I-guided AIGIs and their natural counterparts can be interpreted via the pixel-level quality measures?
+  • RQ1.a: What discrepancies between I2I-guided AIGIs and their natural counterparts can be interpreted via the structure-level quality measures?
   • RQ1.b: What discrepancies between multimodal-guided AIGIs and their natural counterparts can be revealed via the frame-level metrics?
@@
 RQ5: Are human assessment results consistent with quantitative measures? What are the discrepancies revealed from human evaluation?
-4.1 Experimental Setting
-We use the described quantitative evaluation metrics, which can be divided into Full-reference and No-reference, to present the $DR$ for involved assessment dimensions (more details of all metrics are presented in Appendix Sec.
+4.1. Experimental Setting
+We use the described quantitative evaluation metrics, which can be divided into Full-reference and No-reference, to present the $DR$ for involved assessment dimensions (more details of all metrics are presented in Appendix.
 C.
 Full-reference metrics:
-For full-reference metrics, which are used for pixel-level image quality, visual content distribution, and downstream object segmentation task, i.e., IoU, we directly report the calculated value to reflect the difference cause it already can present the discrepancy.
+For full-reference metrics, which are used for structure-level image quality, visual content distribution, and downstream object segmentation task, i.e., IoU, we directly report the calculated value to reflect the difference cause it already can present the discrepancy.
 No-reference metrics
 We calculate Difference Rate (DR) for each quantitative metric in non-reference scenarios, by subtracting the value of each AI-generated image from its corresponding natural reference image and then averaging these difference rates, defined as follows:
-$\text{{DR}}=\frac{1}{n}\sum_{i=1}^{n}\frac{\lvert((M(\mathbf{X}_{i})-m_{\text{min}})-(M(\mathbf{N}_{i})-m_{\text{min}}))\rvert}{m_{\text{max}}-m_{\text{min}}}$ (1)
+(1) $\text{{DR}}=\frac{1}{n}\sum_{i=1}^{n}\frac{\lvert((M(\mathbf{X}_{i})-m_{\text{min}})-(M(\mathbf{N}_{i})-m_{\text{min}}))\rvert}{m_{\text{max}}-m_{\text{min}}},$
 where $n$ is the number of images, $M$ represents the metric, $\mathbf{X}$ denotes the generated images, and $\mathbf{N}$ refers to the referenced natural images.
 Additionally, $m_{\text{min}}$ and $m_{\text{max}}$ represent the minimum and maximum values across all images, respectively, especially when the value range of the corresponding metric is undefined.
-4.2 RQ1: Naive Quality Results
-To evaluate the image quality of AI-generated images versus natural images, we conducted a comprehensive analysis using a variety of metrics that assess the discrepancies between AIGIs and natural images across pixel-level similarity, frame-level quality and visual content distribution.
+4.2. RQ1: Naive Quality Results
+To evaluate the image quality of AI-generated images versus natural images, we conducted a comprehensive analysis using a variety of metrics that assess the discrepancies between AIGIs and natural images across structure-level similarity, frame-level quality and visual content distribution.
 Our findings reveal substantial discrepancies between AIGIs and natural images across these dimensions.
-4.2.1 RQ1.a Pixel-level
+4.2.1. RQ1.a Structure-level
 We evaluated the visual similarity between AI-generated and natural images using structural metrics like SSIM, LPIPS, DISTS, PSNR, etc.
@@
 Additionally, PSNR values between 12.29 and 23.11 suggest that AI-generated images have higher noise levels, reducing their overall visual similarity.
-Table 2: Full-reference Evaluation Results (I2I).
-SD14 SD15 VD SD21 SDxl D·E 2
+Table 3. Full-reference Evaluation Results (I2I).
+Metrics SD14 SD15 VD SD21 SDxl D·E 2
 SSIM ↑ 0.35 0.35 0.46 0.43 0.70 0.37
@@
 MAD ↓ 212.91 213.09 197.92 203.89 144.92 217.12
-4.2.2 RQ1.b Frame-level
+4.2.2. RQ1.b Frame-level
 We assessed the frame-level discrepancies of AI-generated and natural images using a suite of metrics, including DBCNN, HyperIQA, LIQE, PIQE, NRQM and QAlign.
@@
 The NRQM difference rates range from 5% to 32%, with images generated by D·E 3 exhibiting particularly high rates, underscoring the challenges AI-generated images encounter in replicating the inherent naturalness of real-world scenes.
-More detailed results with more metrics are presented in Appendix Sec.
+More detailed results with more metrics are presented in Appendix.
 D.1 Figure 13.
-Figure 5: The FID (left, ↓) and Inception Score (right, ↑) of AI-generated images from different source models and guidance.
-4.2.3 RQ1.c Visual Content Distribution
+Figure 5. The FID (left, ↓) and Inception Score (right, ↑) of AI-generated images from different source models and guidance.
+4.2.3. RQ1.c Visual Content Distribution
 We present the FID and Inception Score results, presented in Figure 5, suggesting that AI-generated images often exhibit relatively higher FID scores (up to 50) and lower Inception Scores (low to 14).
@@
 Notably, these models tend to generate better images when a reference image is provided (TI2I), as it has more detailed guidance.
-In conclusion, our evaluation of image quality across pixel-level, frame-level, and content distribution-level reveals significant disparities between AI-generated and natural images.
+In conclusion, our evaluation of image quality across structure-level, frame-level, and content distribution-level reveals significant disparities between AI-generated and natural images.
 These differences highlight the current limitations of AIGC technologies in matching the visual quality and realism of natural images.
-4.3 RQ2: Alignment Results
-We calculate the difference rate of CLIP Scores ($DR_{CS}$) to assess how well the AIGIs align with natural ones in terms of semantic alignment across different guidance types: Text-to-Image (T2I), Image-to-Image (I2I), and Text-and-Image-to-Image (TI2I).
-And we find that AIGIs often struggle to maintain semantic alignment as natural ones, especially the image-only reference provided.
-The results, shown in Table 3, indicate that the $DR_{CS}$ range from 4.02% to 36.69%.
-We also find that even if the AIGIs can get a higher CS sometimes, especially with text and image-guided, they are looking stranger to humans.
-This suggests that current AIGC models often struggle to maintain high fidelity to the provided references, leading to variations that may impact the intended alignment.
-Our findings underscore the challenges that these models face in consistently generating content to accurately reflect the specified input conditions, highlighting both their strengths and limitations in real-world applications.
-Table 3: The results of $DR_{CS}$ (%).
+4.3. RQ2: Alignment Results
+We calculate the difference rate of CLIP Scores ($DR_{CS}$) to assess how well AI-generated images (AIGIs) align with natural images in terms of semantic consistency across three types of guidance: Text-to-Image (T2I), Image-to-Image (I2I), and Text-and-Image-to-Image (TI2I).
+Our results show that AIGIs often struggle to achieve the same level of semantic alignment as natural images, particularly when guided solely by image input.
+As presented in Table 4, the $DR_{CS}$ ranges from 4.02% to 36.69%.
+This suggests that current AIGIs often face difficulties in preserving high-fidelity alignment with the provided inputs, leading to outputs that deviate from natural counterparts.
+These findings highlight the ongoing challenges that AIGIs encounter in faithfully and consistently capturing the intended semantics, underscoring both their capabilities and limitations in real-world scenarios.
+Table 4. The results of $DR_{CS}$ (%).
 GALIP DFGAN SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
@@
 TI2I - - 9.90 9.73 8.22 9.20 3.57 - -
-4.4 RQ3: Aesthetic Appeal
-We assessed the aesthetic appeal of AI-generated and natural images using NIMA and LAION-AES metrics and found that AI-generated images generally fall short of natural images in terms of aesthetic quality.
-Figure 6 presents the aesthetic difference rate ($DR_{aes}$) results, which reveal that AI-generated images have noteworthy discrepancies compared to natural ones, with $DR_{aes}$ ranging from 3.51% to 27.66%.
-These findings indicate that while AI models can produce visually appealing images, AIGIs’ often lack the nuanced artistic quality and emotional impact that are characteristic of natural images.
-Figure 6: The $DR_{aes}$ between AIGIs and natural ones.
-4.5 RQ4: Applicability Results
-When evaluating the applicability of AI-generated images, we focused on three crucial downstream tasks: image recognition, semantic segmentation and VQA.
-Our findings indicate significant discrepancies between the performance of AIGC images and natural images in these tasks.
-Image Recognition:
-We measured the classification Difference Rate (DR) (%) to identify the differences in predicted labels between AIGC images and their natural counterparts.
-The results in Table 8 showed a substantial gap, with the DR ranging from 27.46% to 94.29%, indicating that the generated images can’t accurately present the object as expected.
-Semantic Segmentation:
-We evaluated the performance using the Intersection over Union (IoU) metric and show the discrepancy results in Table 9.
-The $IoU$ values (higher is better) for AIGC images varied between 0.33 and 0.82, indicating a wide range of segmentation between AIGC images and their natural counterparts.
-Visual Question Answering:
-Further, we apply these AI-generated image to the VQA task by BLIP-2 (Li et al., 2023b) involved three different types of questions: Object, Alignment (Align), and similarity probability (Prob.).
-Table 4 summarizes the discrepancies (%) between natural and AI-generated images across various questions.
-For Object, the highest discrepancies occurred in the I2I setting, with SD14 and SD15 reaching 96.41% and 96.95%, respectively, while T2I discrepancies ranged from 64.07% (DALL·E 3) to 88.43% (DFGAN).
-Align questions showed lower discrepancies, with T2I values ranging from 8.55% (DFGAN) to 21.18% (DALL·E 3) and I2I values from 7.87% (VD) to 10.07% (SD15).
-For similarity probability (Prob.), T2I discrepancies varied between 37.31% (DALL·E 2) and 62.69% (DFGAN), while I2I results ranged from 32.81% (DALL·E) to 61.56% (SD15).
-Overall, our findings highlight significant limitations of AI-generated images in tasks such as image recognition, semantic segmentation, and visual question answering (VQA).
-While AI-generated images perform comparably to natural images in simpler tasks, such as binary text-image alignment (e.g., responding with “yes” or “no”), they fall short in more complex scenarios, including accurately estimating text-image alignment probabilities, identifying specific objects, and segmenting object areas.
-These shortcomings limit their broader applicability in diverse downstream tasks compared to natural images.
-Our results underscore the need for continued research and development to improve the reliability and accuracy of AIGC models, enabling their effective deployment in real-world applications.
-Table 4: The Difference Rate (DR) of VQA tasks.
+4.4. RQ3: Aesthetic Appeal
+We assessed the aesthetic appeal of AIGIs and natural images using two widely adopted metrics: NIMA and LAION-AES.
+Our results indicate that AI-generated images generally fall short of natural images in terms of aesthetic quality.
+As shown in Figure 6, the aesthetic difference rate ($DR_{aes}$) ranges from 3.51% to 27.66%, highlighting noticeable perceptual discrepancies between AIGIs and their natural counterparts.
+These findings suggest that although current AI models are capable of generating visually appealing content, AIGIs often lack the nuanced artistic detail and emotional resonance that are characteristic of natural images.
+This aesthetic gap underscores the limitations of existing generative systems in replicating the subtle visual cues and expressive depth found in human-created imagery.
+Figure 6. The $DR_{aes}$ between AIGIs and natural ones.
+Table 5. Image Recognition Discrepancy Rate ($DR_{re}$).
+GALIP DFGAN SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
+T2I 68.60 85.46 66.14 65.50 68.12 64.88 64.88 65.36 69.26
+I2I - - 94.29 94.07 40.90 83.30 27.96 47.98 -
+TI2I - - 61.62 59.71 60.05 55.54 27.46 - -
+Table 6. IoU of Object Segmentation.
+GALIP DFGAN SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
+T2I 0.23 0.24 0.26 0.25 0.26 0.25 0.26 0.26 0.28
+I2I - - 0.38 0.38 0.69 0.50 0.82 0.42 -
+TI2I - - 0.43 0.43 0.29 0.53 0.82 - -
+4.5. RQ4: Applicability Results
+We evaluated the applicability of AIGIs across three representative downstream tasks: image recognition, semantic segmentation, and visual question answering (VQA).
+Our results reveal substantial performance discrepancies between AI-generated and natural images across all tasks, raising concerns about the practical utility of current AIGIs in real-world applications.
+4.5.1. Image Recognition
+We quantified recognition inconsistencies using the Difference Rate ($DR_{re}$, %) of predicted labels.
+As shown in Table 5, $DR_{re}$ values ranged from 27.46% to 94.29%, indicating that AI-generated images often fail to accurately represent the intended visual concepts, leading to incorrect or unstable recognition outcomes.
+4.5.2. Semantic Segmentation
+Semantic segmentation performance was evaluated using Intersection over Union (IoU), computed based on segmentation results produced by U2Net (Qin et al., 2020).
+Table 6 reports IoU values ranging from 0.33 to 0.82, with a mean IoU of 0.40.
+These results suggest that generated images frequently suffer from imprecise object boundaries and incorrect spatial configurations, limiting their usability for structure-level understanding tasks.
+4.5.3. Visual Question Answering
+We further assessed AI-generated images in the context of visual question answering using BLIP-2 (Li et al., 2023), focusing on three types of questions: Object Recognition, Semantic Alignment (Align), and Similarity Probability (Prob.) (The prompt template used here please refer to Appendix D.2).
+As summarized in Table 7, object-based questions exhibited the highest discrepancy rates, particularly under the I2I setting, reaching up to 96.41% (SD14) and 96.95% (SD15).
+Binary alignment questions (only answer ”yes” or ”no”) yielded lower discrepancies, with an average $DR$ of 11.55%, while similarity-based probabilistic questions underscoring substantial discrepancies compared to natural images, with an average $DR$ of 50.41%.
+Summary.
+Although AI-generated images perform relatively well as natural ones on simpler binary tasks, their effectiveness significantly deteriorates in more complex settings that require fine-grained visual understanding, such as object recognition, spatial reasoning, and nuanced semantic interpretation.
+These limitations currently hinder the broader adoption of generative images in downstream vision tasks.
+Our findings highlight an urgent need for further research into improving the structural fidelity and semantic coherence of generated content to enable trustworthy and task-reliable deployment.
+Table 7. The Difference Rate (DR,%) of VQA tasks.
 GALIP DFGAN SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
@@
 TI2I - - 50.96 50.34 45.93 49.97 41.48 - -
-4.6 Human Validation
-To rigorously assess the differences between AI-Natural images and to evaluate the effectiveness of existing metrics in measuring AIGC image quality, we conducted human evaluations focusing on three key aspects: Image Quality, Alignment, and Aesthetic Appeal.
-  • Image Quality: Participants assessed the overall quality of the images, considering factors such as clarity, detail, and the presence of artifacts.
-  • Alignment: This aspect measured how well the generated images matched the given prompts.
-  • Aesthetic: The aesthetic appeal of the images was rated based on visual attractiveness, composition, and overall artistic quality.
-We excluded results for Applicability since this dimension inherently involves human-generated ground truth labels.
-Participants rated images on a scale from 1 (poor) to 5 (excellent).
-Our human evaluation results are shown in Table 5 and Figure 7 reveal significant discrepancies between AI-generated images and natural images across all three assessed aspects: quality, alignment, and aesthetic appeal.
-AIGC images consistently received lower scores across all three dimensions; AI-generated images generally received lower scores, often between 2 and 3, compared to natural images.
-This emphasizes the need for further advancements to narrow these gaps.
-These findings underscore the importance of continued research and development in AI image generation to bring AIGC images closer to the standards of natural images.
-Moreover, the result in Figure 7 from human validation indicates that traditional quantitative metrics often fail to align with human preferences, suggesting that they are not fully compatible with assessing AIGC images.
-This highlights the necessity for developing new measures tailored specifically to the unique characteristics of AI-generated content.
-Table 5: Human score results across three aspects.
-SD21 SDxl VD D·E 2
+4.6. Human Validation
+To rigorously examine the gap between AI-generated images and human expectations, as well as to evaluate the reliability of existing metrics in assessing AI-generated images, we conducted a comprehensive human evaluation study.
+Participants were asked to score images across three key dimensions: Image Quality, Alignment, and Aesthetic Appeal.
+  • Image Quality: Evaluates the overall visual fidelity of the image, including clarity, sharpness, and the presence of visual artifacts.
+  • Alignment: Assesses the semantic consistency between the generated image and its corresponding prompt, which could be textual, visual, or a combination of both.
+  • Aesthetic Appeal: Measures the artistic and visual attractiveness of the image, considering aspects such as composition, color balance, and overall presentation.
+Participants were instructed to rate each image on a 5-point scale (1 = poor, 5 = excellent) for each of the three evaluation aspects using a radio-button interface.
+4.6.1. Interface for Human Evaluation
+The evaluation was implemented through a custom-designed interface (please refer to Appendix.
+D.3, Figure 15).
+This interface enabled participants to view AI-generated images alongside their reference prompts and assign scores accordingly.
+The prompts used in the evaluation were drawn from the D-ANI dataset and were provided in three modalities: (1) text-only, (2) image-only, and (3) text–image combinations.
+The evaluated images were generated by four state-of-the-art models: SD21, SDxl, Versatile Diffusion, and DALL·E 2.
+To support score calibration and mitigate evaluation bias, we incorporated qualitative reference examples into the interface (please refer to Appendix.
+D.3, Figure 16).
+These examples illustrated typical characteristics of both strong and weak performance across the three evaluation aspects, without disclosing specific numeric scores, thereby reducing potential anchoring effects.
+4.6.2. Evaluation Process
+A total of 58 volunteers participated in the evaluation.
+Each participant rated a curated set of images based on the provided prompts and scoring criteria.
+To assess the consistency and reliability of the responses, we included duplicated image entries within the evaluation set.
+By comparing the scores assigned to these repeated images, we identified inconsistent raters and excluded their data from further analysis.
+After this filtering process, 47 participants’ responses were retained as valid samples for final analysis.
+4.6.3. Participant Demographics
+The participants ranged in age from 19 to 46 years, with a gender distribution of 60% male and 40% female.
+The majority held at least an undergraduate degree, ensuring a well-educated evaluation cohort.
+In addition to scoring data, we collected basic demographic information and invited participants to provide open-ended feedback on their evaluation experience, particularly regarding the perceived differences between AI-generated and natural images.
+4.6.4. Results and Observations
+Quantitative results from the human study are summarized in Table 8, and visualized in Figure 7.
+The findings reveal a pronounced gap between AI-generated images and human expectations across all three evaluation dimensions.
+Specifically, AIGC images consistently received modest scores—typically between 2 and 3—indicating perceptible shortcomings in terms of visual quality, semantic fidelity, and aesthetic appeal.
+These results highlight the need for continued advancements in image generation technologies to better align with human perceptual standards.
+Moreover, the analysis reveals that conventional quantitative metrics often fail to reflect human perceptual judgments (see Figure 7), suggesting that such metrics may be insufficient for evaluating AIGC images.
+This underscores the importance of developing new, human-aligned evaluation measures tailored to the unique characteristics of AI-generated visual content.
+Table 8. Human scores across three aspects.
+Aspect SD21 SDxl VD D·E 2
 T2I I2I TI2I T2I I2I TI2I T2I I2I TI2I T2I I2I
@@
 Aesthetic 3.38 3.55 2.70 3.21 2.35 4.29 3.96 3.32 3.19 4.28 3.23
-Figure 7: The difference rate provided by human & quantitative statistics, where ‘H’ represents human and ‘Q’ represents quantitative.
-4.7 Discrepancies Across Categories
-To analyze the differences between AI-generated and natural images across various categories, we evaluated the discrepancy of AIGI from eight categories (‘person’, ‘animal’, ‘indoor’, ‘outdoor’, ‘vehicle’, ‘food’, ‘sports’, ‘accessory’) across three key aspects: ➊ Naive Quality, including pixel-level similarity (SSIM), frame-level quality (CLIPIQA), and content distribution (FID, CLIP-FID, and Inception Score); ➋ Alignment, measured by CLIP Score; and ➌ Aesthetic Quality, assessed by LAION-AES.
-The results in Figure 8 reveal substantial variability across different categories, even for the AIGI from the same model (please refer to Appendix Sec.
-D.2, Figures 16 and 17 for integrated details).
-This could be attributed to the significant imbalance in the distribution of training data across these categories (Schuhmann et al., 2022b; Aghajanyan et al., 2023).
-Figure 8: The difference rate on food and sport subcategories across quality, alignment, and aesthetic.
-5 Conclusion and Discussion
-In this paper, we develop a systematic and comprehensive D-Judge to evaluate the discrepancies between AI-Natural images using the constructed D-ANI dataset.
-Our findings reveal significant disparities in naive image quality, semantic alignment, aesthetics, and downstream task applicability, especially in the low-level and content distribution-level image quality and fine-grained downstream tasks, like object recognition and object VAQ.
-While AIGC models have achieved impressive innovations, AIGIs still lag behind natural images in these critical areas.
-Besides, human evaluations also reveal lower scores for AIGC in terms of visual quality, semantic alignment, and aesthetics, emphasizing the gaps between AIGIs and natural images.
-Finally, the poor utility of AIGIs in downstream tasks indicates practical application challenges for AIGC.
-This study also has limitations, including reliance on specific quantitative metrics and models available at the time.
-Future studies should incorporate more generative models, larger datasets, and sophisticated metrics to capture further discrepancies.
-Addressing these limitations and improving models will help reduce the perceptual gaps between AIGC images and natural images, facilitating better integration into real-world applications.
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Figure 7. The difference rate provided by human & quantitative statistics, where ‘H’ represents human and ‘Q’ represents quantitative.
+4.7. Discrepancies Across Categories
+To systematically investigate the differences between AI-generated images (AIGI) and natural images across various content categories, we evaluate discrepancies in eight representative categories: person, animal, indoor, outdoor, vehicle, food, sports, and accessory.
+The evaluation is conducted along three critical dimensions: ➊ Naive Quality, assessed via structure-level similarity (SSIM), frame-level quality (CLIPIQA), and content distribution metrics (FID, CLIP-FID, and Inception Score); ➋ Alignment, measured using CLIP Score; and ➌ Aesthetic Quality, evaluated through LAION-AES.
+Representative results are illustrated in Figures 8 and 9.
+Specifically, Figure 8 presents the difference rate ($DR$) across CLIPIQA, CLIP Score, and LAION_AES on the food and sports subcategories, reflecting variations in visual quality, semantic alignment, and aesthetic perception.
+Figure 9 further provides FID and Inception Score comparisons that reveal discrepancies in content distribution for the person and accessory categories.
+These results demonstrate that the extent of divergence between AI-generated and natural images varies significantly across categories, even when produced by the same generative model.
+A potential cause of this variability is the inherent imbalance in training data distributions across different semantic categories (Schuhmann et al., 2022; Aghajanyan et al., 2023).
+Additional results and visualizations are provided in Appendix D.4, Figures 17 and 18, offering a comprehensive view of these discrepancies.
+Notably, Figures 18 suggest that images generated by D·E 3 display the most pronounced deviations from natural references across multiple metrics.
+This could be attributed to D·E 3’s inherent tendency to generate visually appealing but stylized or surreal imagery, which differs from the more realism-oriented outputs of other models (see  B, Figure 10 for visual comparisons).
+Figure 8. The difference rate on food and sport subcategories across quality, alignment, and aesthetic.
+Figure 9. The FID and Inception Score (IS) of AI-generated images from different source models and categories from content distribution level.
+5. Conclusion
+In this paper, we present a comprehensive D-Judge based on the constructed D-ANI dataset to evaluate the discrepancies between AI-generated and natural images.
+The results reveal substantial gaps across key dimensions such as image quality, semantic alignment, and downstream applicability.
+Despite recent progress in AIGC models, human evaluations consistently show that AI-generated images underperform in visual fidelity, semantic relevance, and aesthetic quality.
+Our benchmark also yields actionable guidance for improving model design, such as enhancing semantic consistency and prompt-image alignment, particularly in I2I tasks.
+Promising directions include reference-aware attention, perceptual alignment losses, and joint image-text encoding.
+Limitation and Future Work.
+This study has certain limitations, including its reliance on quantitative metrics and a limited set of generative models.
+Future work should incorporate more diverse and balanced datasets, a broader range of contemporary models, and task-aware evaluation protocols to uncover deeper discrepancies.
+Beyond evaluation, D-Judge can also serve as a training signal—for instance, as a reward module in reinforcement learning or preference-based fine-tuning—to jointly optimize semantic alignment and downstream usability.
+We will release our dataset and code to support such use cases.
+Addressing these limitations and advancing model architectures will be essential for narrowing the perceptual and functional gaps between AI-generated and natural images, ultimately enabling their trustworthy deployment in real-world scenarios.
+Acknowledgments
+This work is supported by Guangdong Basic and Applied Basic Research Foundation (2023A1515012848) and CCF-DiDi GAIA Collaborative Research Funds.
+This research is supported by A*STAR, CISCO Systems (USA) Pte.Ltd and National University of Singapore under its Cisco-NUS Accelerated Digital Economy Corporate Laboratory (Award I21001E0002).
+We would like to express our gratitude to all the volunteers who participated in our human evaluation experiments.
 Appendix A Appendix Overview
@@
   • Section E: Discussion on the Safety Mechanisms in Existing Diffusion Models.
-Appendix B Details of the D-ANI Dataset
-B.1 Overview
+Appendix B Additional Details of the D-ANI Dataset
 The D-ANI dataset is a carefully curated collection of both natural and AI-generated images, designed to support comprehensive evaluations of image generation models.
 The natural images are sourced from the COCO validation set.
-AI-generated images were produced using eight different models.
+AI-generated images were produced using nine different models.
 These images were generated under three distinct guidance modes: Text-to-Image (T2I), Image-to-Image (I2I), and Text vs.
 Image-to-Image (TI2I).
-Representative samples from the D-ANI dataset are shown in Figures 9,  10 and  11.
-For a detailed breakdown of the D-ANI dataset composition, please refer to Table 6.
-Figure 9: Examples of AI-generated content (AIGC) created using Text-to-Image (T2I) guidance.
-Figure 10: Examples of AI-generated content (AIGC) created using Image-to-Image (I2I) guidance.
-Figure 11: Examples of AI-generated content (AIGC) created using Text-and-Image-to-Image (TI2I) guidance.
-B.2 Natural Images
-The D-ANI dataset features a carefully curated selection of natural images sourced from the COCO validation set.
-This subset comprises 5,000 images, each paired with at least five unique captions, totaling 25,014 captions.
-These rich and varied textual descriptions enhance the dataset’s utility across a wide range of image-to-text and text-to-image tasks, making it an invaluable resource for evaluating and training image generation and understanding models.
-In this study, we use these images and their accompanying (five) captions to guide the generative models in creating new images.
-B.3 Generated Images
-The D-ANI dataset includes AI-generated images created using several advanced generative models across three distinct guidance modes: Text-to-Image (T2I), Image-to-Image (I2I), and Text-and-Image-to-Image (TI2I).
-These modes collectively contributed a significant volume of images to the dataset.
-The models utilized in this process include DF-GAN, GALIP, various versions of Stable Diffusion (v1.4, v1.5, v2.1, XL), Versatile Diffusion, DALL·E 2 and DALL·E 3.
-Each model, except DALL·E 2 and DALL·E 3, generated 25,000 images per guidance mode.
-DALL·E 2 and DALL·E 3 produced 5,000 images for per mode, resulting in a total of 440,000 generated images.
-It is important to note that DF-GAN, GALIP and DALL·E 3 were used only for T2I, while DALL·E 2 was employed for both T2I and I2I modes.
-For detailed information, please refer to Table 6.
+Representative samples from the D-ANI dataset are shown in Figures 10,  11 and  12.
+Figure 10. Examples of AI-generated content (AIGC) created using Text-to-Image (T2I) guidance.
+Figure 11. Examples of AI-generated content (AIGC) created using Image-to-Image (I2I) guidance.
+Figure 12. Examples of AI-generated content (AIGC) created using Text-and-Image-to-Image (TI2I) guidance.
 The involved generative models are as follows:
@@
 While continuing to use the diffusion process, SD15 incorporates better optimization strategies and larger datasets, yielding superior results.
-  • Versatile Diffusion: (Xu et al., 2023b) is a diffusion-based model designed for a variety of image generation tasks, including style transfer, image inpainting, and super-resolution.
+  • Versatile Diffusion: (Xu et al., 2023) is a diffusion-based model designed for a variety of image generation tasks, including style transfer, image inpainting, and super-resolution.
 Its versatility lies in its ability to adapt the diffusion process to the specific requirements of each task, making it suitable for diverse applications.
@@
 From text-to-image generation and high-resolution synthesis to versatile image processing, these models push the boundaries of AI-generated content.
-B.4 Image Resolution
-The resolution of the generated images varies across models.
-GALIP produces images at 224x224 pixels, while DF-GAN generates slightly higher-resolution images at 256x256 pixels.
-Models such as various versions of Stable Diffusion and DALL·E 2 generate images at a resolution of 512x512 pixels.
-Notably, DALL·E 3 delivers the highest resolution among these models, producing images at 1024x1024 pixels.
-B.5 Composition
-The D-ANI dataset offers a robust foundation for evaluating the performance and capabilities of various AIGC models under different conditions and guidance modes.
-By encompassing a broad spectrum of images and resolutions, the dataset ensures comprehensive coverage of the potential use cases and challenges associated with AI-generated content.
-The combination of natural and generated images, alongside diverse textual descriptions, facilitates a thorough assessment of image generation models, supporting the development of more accurate and reliable AI systems.
-Table 6: The details of the D-ANI dataset.
-Types DFGAN GALIP SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
-T2I 25,000 25,000 25,000 25,000 25,000 25,000 25,000 5,000 5,000
-I2I - - 25,000 25,000 25,000 25,000 25,000 5,000 -
-TI2I - - 25,000 25,000 25,000 25,000 25,000 - -
 Appendix C Evaluation Metrics
-C.1 Overview
+C.1. Overview
 In this study, we selected a comprehensive set of metrics to evaluate the image quality, alignment, and aesthetics of both natural and AI-generated images.
 These metrics, summarized in Table 9, are carefully chosen to highlight the differences across multiple dimensions, providing a robust framework for assessing the quality and realism of AIGC in comparison to natural images.
-To evaluate image quality, we divided the assessment into three sub-aspects: pixel-level similarity, frame-level visual metrics, and visual content distribution.
-  • Pixel-level: Metrics such as SSIM, PSNR, LPIPS, DISTS, VIF, VSI, FSIM, and MAD assess structural, perceptual, and textural fidelity.
+To evaluate image quality, we divided the assessment into three sub-aspects: structure-level similarity, frame-level visual metrics, and visual content distribution.
+  • Structure-level: Metrics such as SSIM, PSNR, LPIPS, DISTS, VIF, VSI, FSIM, and MAD assess structural, perceptual, and textural fidelity.
 These metrics measure how closely the generated images match natural ones in terms of structural integrity and perceptual similarity.
@@
 These metrics capture broad differences in content quality.
-For alignment, we utilize the CLIP Score (Radford et al., 2021a) to measure the correspondence between text descriptions and the generated images.
+For alignment, we utilize the CLIP Score (Hessel et al., 2021) to measure the correspondence between text descriptions and the generated images.
 This metric evaluates how well the generated content aligns with the given prompts by embedding both text and image into a shared space and calculating their similarity.
@@
 The chosen metrics, as detailed in Table 9, collectively provide a comprehensive toolkit for evaluating image quality, alignment, and aesthetics, each focusing on different aspects of visual perception and fidelity.
-C.2 Naive Image Quality
+C.2. Naive Image Quality
 We evaluated the naive image quality of AI-generated images versus natural images using a range of metrics.
-These metrics assess both pixel-level and frame-level qualities, as well as visual content distribution, to capture the holistic differences between AI-generated and natural images.
-Pixel-Level Quality Metrics:
+These metrics assess both structure-level and frame-level qualities, as well as visual content distribution, to capture the holistic differences between AI-generated and natural images.
+Structure-Level Quality Metrics:
   • SSIM (Wang et al., 2004) (Structural Similarity Index) compares the structural information between a reference and a test image, considering luminance, contrast, and structure for a comprehensive measure of similarity.
   • PSNR (Peak Signal-to-Noise Ratio) measures the ratio between the maximum possible power of an image and the power of corrupting noise, with higher values indicating better image quality.
-  • VIF (Sheikh & Bovik, 2006) (Visual Information Fidelity) quantifies the amount of visual information preserved in the test image relative to the reference, based on natural scene statistics and the human visual system.
+  • VIF (Sheikh and Bovik, 2006) (Visual Information Fidelity) quantifies the amount of visual information preserved in the test image relative to the reference, based on natural scene statistics and the human visual system.
   • VSI (Zhang et al., 2014) (Visual Saliency-based Index) evaluates perceptual similarity by focusing on how salient (prominent) features in the images compare.
@@
   • DISTS (Ding et al., 2022) (Deep Image Structure and Texture Similarity) combines structural and textural similarity metrics using deep learning features to evaluate overall image similarity.
-  • MAD (Larson & Chandler, 2010) (Mean Absolute Deviation) provides a straightforward assessment of pixel-level differences, with lower values indicating greater similarity between images.
+  • MAD (Larson and Chandler, 2010) (Mean Absolute Deviation) provides a straightforward assessment of structure-level differences, with lower values indicating greater similarity between images.
 Frame-Level Quality Metrics:
@@
   • CNNIQA (Kang et al., 2014) (Convolutional Neural Network Image Quality Assessment) leverages deep convolutional networks for no-reference image quality assessment using learned features.
-  • CLIPIQA (Radford et al., 2021b) (CLIP Image Quality Assessment) uses CLIP model embeddings to assess image quality by comparing the alignment between image and text descriptions.
-  • TReS (Radford et al., 2021b) (Textural and Edge-based Similarity) evaluates image quality by focusing on textural and edge-based features, providing a no-reference assessment.
+  • CLIPIQA (Radford et al., 2021) (CLIP Image Quality Assessment) uses CLIP model embeddings to assess image quality by comparing the alignment between image and text descriptions.
+  • TReS (Radford et al., 2021) (Textural and Edge-based Similarity) evaluates image quality by focusing on textural and edge-based features, providing a no-reference assessment.
   • HyperIQA (Su et al., 2020) (Hyper Network-based Image Quality Assessment) uses a hypernetwork to predict image quality, leveraging deep learning for no-reference quality assessment.
@@
   • NRQM (Ma et al., 2018) (Naturalness Image Quality Metric) assesses how closely an image aligns with the statistical properties of natural scenes based on natural scene statistics and perceptual models.
-  • QAlign (Wu et al., 2024b) evaluates image quality by assessing its alignment with semantic and perceptual features indicative of high visual fidelity.
+  • QAlign (Wu et al., 2024) evaluates image quality by assessing its alignment with semantic and perceptual features indicative of high visual fidelity.
 Visual Content Distribution Metrics:
@@
   • Inception Score (Salimans et al., 2016) evaluates the quality of generated images by considering the confidence of the classifier and the diversity of the generated samples, with higher scores indicating better quality and diversity.
-C.3 Alignment
+C.3. Alignment
 To assess alignment, we employ the CLIP Score (Contrastive Language-Image Pre-training Score, CS), which measures the correspondence between text descriptions and generated images.
 This metric evaluates how well the visual content aligns with the textual prompts by embedding both into a shared space and calculating their cosine similarity.
-C.4 Aesthetic
-  • NIMA (Esfandarani & Milanfar, 2018) (Neural Image Assessment) is a deep learning-based model that predicts the aesthetic quality of images.
+C.4. Aesthetic
+  • NIMA (Esfandarani and Milanfar, 2018) (Neural Image Assessment) is a deep learning-based model that predicts the aesthetic quality of images.
 It generates a score that reflects human perception of visual appeal.
-  • LAION-AES (Schuhmann et al., 2022a) (LAION Aesthetic Score) evaluates images’ aesthetic quality using various visual features, providing a numerical value to indicate their aesthetic appeal.
-  • QAlign (Wu et al., 2024b) assesses image aesthetics by analyzing its adherence to human visual preferences and artistic principles, leveraging semantic coherence and appeal.
+  • LAION-AES (Schuhmann et al., 2022) (LAION Aesthetic Score) evaluates images’ aesthetic quality using various visual features, providing a numerical value to indicate their aesthetic appeal.
+  • QAlign (Wu et al., 2024) assesses image aesthetics by analyzing its adherence to human visual preferences and artistic principles, leveraging semantic coherence and appeal.
 Together, these metrics offer a robust toolkit for evaluating images, focusing on alignment with prompts and aesthetic quality, thereby covering key aspects of visual perception and fidelity.
-Table 7: The detail of evaluation metrics.
+Table 9. The detail of evaluation metrics.
 Aspects Sub aspects Metrics Reference
-Naive Quality Pixel-level SSIM TRUE
+Naive Quality Structure-level SSIM TRUE
 PSNR TRUE
@@
 QAlign (aesthetic) FALSE
-Figure 12: We evaluate nine image-generating models from 14 aspects. The numerical values in the radar chart represent the mean difference rate of each model and each metric.
-Appendix D Evaluation Results
-D.1 Naive Quality
+Figure 13. We evaluate nine image-generating models from 14 aspects. The numerical values in the radar chart represent the mean difference rate of each model and each metric.
+Appendix D Evaluation
+D.1. Naive Quality
 We present the additional results of the Naive Image Quality assessment in Figure 13.
-D.2 Applicability
-We put the applicability results of Image Recognition and Semantic Segmentation in Tables 8 and  9.
-Table 8: Image Recognition Discrepancy Rate ($DR_{re}$).
-GALIP DFGAN SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
-T2I 68.60 85.46 66.14 65.50 68.12 64.88 64.88 65.36 69.26
-I2I - - 94.29 94.07 40.90 83.30 27.96 47.98 -
-TI2I - - 61.62 59.71 60.05 55.54 27.46 - -
-Table 9: IoU of Object Segmentation.
-GALIP DFGAN SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
-T2I 0.23 0.24 0.26 0.25 0.26 0.25 0.26 0.26 0.28
-I2I - - 0.38 0.38 0.69 0.50 0.82 0.42 -
-TI2I - - 0.43 0.43 0.29 0.53 0.82 - -
-D.3 Human Perceptual Evaluation
-D.3.1 Interface for Human Evaluation
-The human evaluation of our D-ANI dataset was conducted using a custom-designed interface (Figure 14).
-This interface allowed participants to evaluate AI-generated images in comparison to reference content across three key aspects: Quality, Alignment, and Aesthetics.
-The reference content included prompts used to generate the images, categorized into three types: 1) text, 2) image, and 3) text and image.
-The AI-generated images evaluated were produced by four different models: SD21, SDxl, Versatile Diffusion, and DALL·E 2.
-D.3.2 Evaluation Process
-A total of 58 volunteers participated in the evaluation process.
-Each volunteer assessed the images based on the provided reference content.
-To ensure the validity of the data collected, we included duplicate images in the evaluation system.
-By comparing the scores given to these identical images, we could determine the consistency of each participant’s evaluations.
-Only data from participants who showed consistent scoring were considered valid, resulting in 47 effective samples.
-D.3.3 Participant Demographics
-The volunteers who participated in the evaluation ranged in age from 19 to 46 years old, with a gender ratio of 6:4 (male to female).
-All participants held at least an undergraduate degree, ensuring a well-educated sample group.
-Alongside scoring data, we collected demographic details such as age and gender.
-Upon completing their evaluations, participants were also invited to provide feedback on the evaluation system and their experience with assessing the differences between natural and AI-generated images.
-D.3.4 Evaluation Metrics
-Participants were instructed to evaluate each image based on three aspects:
-  1. 1.
-    Quality: Overall visual fidelity and clarity of the image.
-  2. 2.
-    Alignment: Accuracy in matching the image to the provided prompt.
-  3. 3.
-    Aesthetic: Visual appeal and artistic quality of the image.
-To facilitate the scoring process, we provided reference examples within the evaluation system (Figure 15).
-These examples illustrated what constituted good or poor performance in each aspect without giving specific scores, thereby preventing any potential bias in participant evaluations.
-Figure 13: Examples for unsafe references.
-Figure 14: Interface for Human Evaluation.
+D.1.1. Aesthetic Appeal
+We present the additional results of QAlign (Wu et al., 2024) regarding to the Aesthetic Appeal assessment in Table 10.
+Table 10. Aesthetic Appeal of QAlign.
+Types DFGAN GALIP SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
+T2I 45.33 39.35 32.44 31.84 35.31 40.91 43.73 26.53 99.34
+I2I - - 23.47 23.09 11.31 20.32 8.17 13.88 -
+TI2I - - 25.8 25.98 23.76 26.62 7.46 - -
+D.2. Applicability Settings
+To facilitate the applicability evaluation in downstream tasks, we construct three standardized Visual Answering Question (VAQ) templates using BLIP-2, each targeting a specific aspect of visual-semantic understanding: object recognition, semantic alignment, and similarity probability.
+As shown in Table 11, each prompt is designed to elicit structured and interpretable responses.
+The object recognition prompt asks the model to identify the main object in the image and return a class label from the COCO taxonomy.
+The alignment prompt evaluates whether the image content semantically matches a given caption, requiring a binary “yes” or “no” answer.
+The similarity probability prompt estimates the degree of alignment between image and caption by requesting a numerical probability.
+These prompts are uniformly applied to both AI-generated and reference images, offering a consistent protocol for probing semantic fidelity across different generation conditions.
+Table 11. The prompt template used for VAQ tasks.
+Task Prompts
+Object Question: what is the object in the image? Only return the class name in {COCO class list}. Answer:
+Alignment Question: the image content aligns very well with {caption}? only return ’yes’ or ’no’. Answer:
+Similarity probability Question: what’s the probability of the image content well align with {caption}? only return a float number. Answer:
+D.3. Human Perceptual Evaluation
+To complement the description in Sec.4.6, we present two figures that detail the setup of our human evaluation protocol.
+Figure15 illustrates the custom-designed interface through which participants evaluated AI-generated images.
+The interface supports side-by-side comparisons and collects ratings across three key dimensions: Quality, Alignment, and Aesthetics.
+For each dimension, participants assigned a score ranging from 1 to 5 using a radio-button selection mechanism, where higher scores indicated better performance.
+Figure 16 shows the set of reference examples embedded within the interface.
+These examples were designed to convey representative cases of both high and low performance for each evaluation dimension.
+They served as non-numeric visual guidelines to help participants develop a consistent evaluation criterion, while avoiding explicit scoring anchors that might introduce bias.
+Figure 14. Examples for unsafe references.
+Figure 15.
+Interface for Human Evaluation.
 Reference content: the prompt guidance used to generate images, including three types: 1) text, 2) image, and 3) text and image.
@@
 Participants can rely on the reference content to evaluate each image on three aspects, i.e., Quality, Alignment, and Aesthetic.
-Figure 15: Before the evaluation process, the participant will be given some examples to learn which image should be good or bad in a specific aspect.
-Figure 16: The differential rate of data from different categories models across naive image quality, alignment, and aesthetic.
-Figure 17: The FID, CLIP-FID and Inception Score (IS) of AI-generated images from different source models and categories from content distribution level.
-D.3.5 Summary
-The comprehensive human evaluation process ensured that the collected data was reliable and reflected genuine human perception.
-The demographic diversity and the rigorous validation of participant responses provided a robust foundation for assessing the performance of various AIGC models in generating high-quality, aligned, and aesthetically pleasing images.
-This approach not only highlighted the strengths and weaknesses of the evaluated models but also offered valuable insights into the perceptual differences between AI-generated and natural images.
-D.4 Discrepancies Across Categories
-This section provides a detailed evaluation of AI-generated images across different categories, focusing on naive image quality, alignment, aesthetics, pixel-level similarity, and content distribution.
+Figure 16. Before the evaluation process, the participant will be given some examples to learn which image should be good or bad in a specific aspect.
+Figure 17. The differential rate of data from different categories models across naive image quality, alignment, and aesthetic.
+Figure 18. The FID, CLIP-FID, and Inception Score (IS) of AI-generated images from different source models and categories from the content distribution level.
+D.4. Discrepancies Across Categories
+This section provides a detailed evaluation of AI-generated images across different categories, focusing on naive image quality, alignment, aesthetics, structure-level similarity, and content distribution.
 These analyses aim to uncover discrepancies between AI-generated and natural images and identify category-specific challenges.
@@
 Figure 18 presents trends in SSIM, FID, CLIP-FID, and Inception Score.
-The ‘accessory’ category consistently shows higher FID values, reflecting poorer content distribution, while the ‘animal’ category achieves higher SSIM scores, demonstrating better pixel-level similarity.
+The ‘accessory’ category consistently shows higher FID values, reflecting poorer content distribution, while the ‘animal’ category achieves higher SSIM scores, demonstrating better structure-level similarity.
 These findings suggest that ‘accessory’ and ‘indoor’ images pose greater challenges for generative models in maintaining semantic consistency and achieving balanced content generation compared to other categories.
-Additionally, the results in Figures 17 and Figures 16 show that compared to other generative models, images generated by D·E 3 display greater differences from natural reference images across various metrics.
-This might be due to D·E 3’s tendency to produce high-quality images with a surreal style, whereas other models primarily focus on generating images with a more realistic style (see Appendix Sec.
-B, Figure 9 for visual comparisons).
 Overall, these results indicate that AI-generated images still lag behind natural images in achieving comparable levels of naturalness and image quality.
@@
 Appendix E Discussion
-Table 10: Prompts trigger the safe-checker of different models.
-T2I I2I TI2I T2I(‰) I2I(‰) TI2I(‰)
+Table 12. Prompts trigger the safe-checker of different models.
+Models T2I I2I TI2I T2I(‰) I2I(‰) TI2I(‰)
 SD14 104 263 141 4.16 10.51 5.64